### v1

# LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data

Abstract.
The semantic capabilities of language models (LMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora. Unfortunately, existing systems lack high-level abstractions to perform semantic queries at scale. We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for semantic queries over datasets (e.g., sorting or aggregating records using natural language criteria). Each operator can be implemented and optimized in multiple ways, opening a rich space for execution plans similar to relational operators. We implement our operators and several optimizations for them in LOTUS, an open source query engine with a Pandas-like API.
We demonstrate LOTUS’ effectiveness across a series of real applications, including fact-checking, extreme multi-label classification, and search. We find that LOTUS’ programming model is highly expressive, capturing state-of-the-art query pipelines with low development overhead across these diverse applications. Specifically, on the FEVER dataset for fact-checking application, LOTUS’ programs can reproduce FacTool, recent state-of-the-art pipeline, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time. In the extreme multi-label classification task on the BioDEX dataset, LOTUS reproduces state-of-the art result quality with its join operator, while providing an efficient algorithm that runs faster than a naive join. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods used by prior works. LOTUS is publicly available at https://github.com/stanford-futuredata/lotus.
1. Introduction
The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI systems that reason over vast knowledge corpora. Many applications require complex reasoning over large amounts of data, including both unstructured and structured data. For example a researcher reviewing recent ArXiv (arx, [n. d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset. Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023). Likewise organizations may wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business needs and productivity (dis, [n. d.]).
Each of these tasks require a form of bulk semantic processing, where the AI system must process large amounts of data in often complex query patterns to perform the reasoning task at hand. Supporting the full generality of these applications with efficient and easily programmable query systems would have transformative impact. This prospect, however, raises two important and challenging questions: first, how should developers express semantic queries, and secondly, how should we design the underlying query system to achieve high efficiency and accuracy.
Unfortunately, existing systems are insufficient for serving applications that require bulk semantic processing.
Many existing LM programming frameworks (lan, [n. d.]; lla, [n. d.])
and research works (Khattab et al., 2023, 2021; Khattab and Zaharia, 2020; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) provide methods, abstractions and optimizations for retrieval-augmented generation (RAG), which first performs a semantic search over the text corpus, then invokes the LM conditioned on the user question and retrieved documents.
While RAG is one common and useful query pattern, it is limited to point lookups into the data corpus and assumes the user query can be answered by one or a small set of retrieved documents. However, bulk semantic processing pipelines may involve more complex patterns not captured by RAG, such as semantic aggregations or transformations over many documents.
Alternatively, several systems (ver, [n. d.]; dat, [n. d.]; noa, 2023a; Liu et al., 2024b) offer and optimize LLM user-defined functions (UDFs) in SQL. This model offers a logically row-wise LM-execution model for batch processing data with arbitrary prompts in composition with arbitrary SQL queries. While useful for batch-processing applications, this model cannot support LM-reasoning patterns across rows and provides a low-level of abstraction with a simple LM() function.
Towards a declarative programming interface for bulk semantic processing, we propose semantic operators, which extend the relational model with AI-based operations that users can compose into powerful, reasoning-based query pipelines over structured and unstructured data. These operators include semantic filters, joins, rankings, aggregations, and projections, which take natural language expressions given by the programmer. To provide an implementation of these operators, we present the LOTUS (LLMs Over Tables of Unstructured and Structured data) system and programming model. LOTUS’ query engine efficiently executes queries with semantic operators using a variety of algorithms and optimizations for each operator, while abstracting away low-level details like model context length limits and choice of algorithms.
| Operator | Description |
| Returns the tuples in a table that pass the provided langex predicate. | |
| Joins a table against a second table by keeping all pairs of tuples that pass the provided langex predicate. | |
| Performs a similarity join where each row of the source table is joined with the most semantically similar tuples from the table , using fields and as the left and right join keys, respectively. | |
| Performs an aggregation over the input tuples according to the langex, which specifies a commutative, associative aggregation function over a list of tuples. | |
| Ranks each tuple and returns the best according to the langex, which specifies a ranking function that sorts a list of tuples. | |
| Performs a projection, returning a new column, according to the provided langex. | |
| Performs a projection according to the langex, returning a column with a list of substrings from the input tuples. | |
| Performs a semantic similarity clustering over column to create groups. | |
| Performs a top- search over column using query . | |
| Creates similarity index for column and saves to path. | |
Figure 1 begins to demonstrate the power of LOTUS’ declarative programming model and optimized query engine. For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we can easily create 3 distinct query pipelines, each written in intuitive LOTUS programs of less than 50 lines of code by composing semantic operators (e.g. filters, maps, search and joins). The modularity and composability of these semantic operators allows us to quickly explore the design space on this task. In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by up to compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), while also providing query efficiency, by default, to maintain modest execution time.
LOTUS’ optimizer also exploits the rich implementation design space of semantic operators to leverage new algorithmic and optimization opportunities. For expensive operators like semantic joins, aggregations, ranking and filters, LOTUS implements novel algorithms that maximize parallel batched-inference opportunities, use model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) with a lightweight scoring function unique to our setting, leverage semantic similarity indices, and perform algorithmic approximations. Figure 1 demonstrates the effectiveness of these methods. Compared to the best-performing un-optimized LOTUS program, Program B shown in the figure, the logically equivalent program implemented using LOTUS’ optimizations attains higher accuracy and runs faster.
We systematically evaluate LOTUS through a series of real applications, including fact-checking, extreme multi-label classification, and search. Our results show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications. Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time. In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, demonstrating the power of LOTUS’ declarative interface. In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods (Qin et al., 2024) used by prior works.
2. The LOTUS Programming Model
We now introduce the LOTUS programming model and show its expressive power in allowing developers to declaratively specify AI-based query pipelines that bulk process large datasets of structured and unstructured data. LOTUS extends the relational model with semantic operators, which we show in Table 1. These operators can be easily composed together with standard relational operators to build powerful programs that are transparently optimized (Section 3). In this section, we describe each semantic operator and our API for them in LOTUS, grounding each in concrete examples. While our current API implementation extends Pandas (pan, [n. d.]), LOTUS’ semantic operators could be used with other existing relational query languages and APIs, such as SQL.
2.1. Datatypes
LOTUS’ data model consists of tables with structured and unstructured text fields, and our current implementation extends Pandas (pan, [n. d.]). Figure 2 shows an example LOTUS program that loads data about ArXiv papers and performs a summarization task, which finds relevant papers, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest. We briefly describe two core components of LOTUS programming model: its data model (Section 2.1.1), and its parameterized natural language expressions (Section 2.1.2) for specifying semantic operations over the data.
2.1.1. Data Model
LOTUS is designed to seamlessly extend the relational model.
Each row in the table, represents a logical entity. Table columns may contain either structured data or unstructured fields with free-form natural-language text. LOTUS’ semantic-relational operators can take both of these data-types as inputs. Figure 3 shows an example table, where each rows represents an ArXiv paper document, with fields for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date. These columns are then passed as parameters to semantic-relational operators, such as sem_search, sem_filter, sem_agg, as shown in Figure 2.
Additionally, LOTUS supports semantic similarity indices over natural-language text columns to provide optimized semantic query processing. These indices create embeddings over each document in the column to capture semantic similarity using embedding distance metrics. Semantic indexes can be created off-line and then loaded using sem_index and load_sem_index. Figure 2 provides an example program, where upon reading the ArXiv papers data from a CSV file, the programmer loads a semantic index over the abstract column. The program then repeatedly uses the semantically-indexed column in subsequent LOTUS operations, involving semantic search, filtering and aggregation.
2.1.2. Parameterized Natural Language Expressions (langex)
A core principle of LOTUS is to provide users with a declarative interface that separates the user-specified, logical query plan from its underlying implementation.
As such, users program with LOTUS’ semantic operators by writing parameterized natural language expressions (langex111Akin to regular expressions (regex), which specify character patterns to match in text, langex are natural language expressions to programmatically specify reasoning-based patterns over structured data and free-form text.),
rather than directly prompting an underlying LM.
Figure 2 shows an example of this, where the programmer provides a langex as the parameter to the sem_filter (line 18) and sem_agg (line 19) operations. Programmers write each langex in natural language text, parameterized by one or more data columns, which are indicated in the double brackets within the formatted string. The function of these expressions varies according to the semantic operator used and may represent a predicate, aggregation, comparator function, or projection in natural language. For instance, as shown in Figure 2, the langex signature of sem_filter provides a predicate that indicates a filter criteria to apply over paper titles and abstracts, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over paper abstracts.
Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user. This makes it simple for the programmer to specify diverse and complex, multi-step query pipelines. All operator-specific prompts are automatically handled by LOTUS’ underlying query engine, and the system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
2.2. Semantic Operators
We now overview each semantic operator and their corresponding LOTUS API. Table 1 provides a concise summary of each operator.
Sem_filter returns the subset of rows that pass the filtering condition, specified by the user’s langex. As Figure 4 shows, the langex signature provides a semantic predicate over one or more table columns and can be answered by a binary ”True” or ”False” answer.
Sem_topk ranks a set of rows according to the user-defined criteria, and returns the rows that best match the ranking criteria. The signature of the langex provides a general ranking criteria according to one or more columns. The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation. As Figure 5 shows, the programmer uses the langex to specify arbitrary reasoning-based ranking criteria, such as ranking paper abstracts by the most outrageous claim made.
Programmers can also optionally specify a group-by parameter to indicate a subset of columns to group over during ranking, as shown in 6. The groupings are defined using standard equality matches over the group-by columns. To use groupings according to semantic similarity, users can also perform a sem_cluster_by and pass the resulting cluster_id column as the group-by column parameter.
Sem_join combines data from two tables, evaluating the user’s predicate to return the set of rows from the left and right table that pass. As Figure 7 shows, users specify the right join table, a langex, and optionally, the join key of the left and right tables. Here the langex contains two or more columns, and provides a predicate over the left and right tables. By default, the operator performs an inner join over the two tables, and can alternatively perform left, right or outer joins, if specified.
Sem_sim_join provides a variant of the semantic join, such that rows are matched according to their semantic similarity, rather than an arbitrary natural-language predicate. Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join. Figure 8 provides an example, where one table contains papers with an indexed column of abstracts, and the other table contains a list of research interests. The user specifies the left and right table join keys, and a parameter . The left join key may or may not be indexed, whereas the the right join key must be a semantically indexed column with its index loaded. The operator performs a left join such that for each row in the left table, the output table will contain matching rows from the right table with the highest similarity scores. The programmer can also optionally specify a return column containing the semantic similarity scores of each joined row.
Sem_agg performs an aggregation over all rows of the table. As Figure 9 shows, the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results. Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents. Similar to sem_topk, users can also specify a group-by parameter to use. We also provide additional flexibility, by allowing the programmer to compose semantic aggregations with sem_partition_by, which provides finer-granularity control over how documents are grouped over in each LM invocation. We describe this further below, and show its use in Section 3 for overriding the default commutativity and associativity assumptions for some tasks.
Sem_partition_by creates a row partitioning over the table, which will be used by sucessive calls to sem_agg to decide which rows to group together within each LM invocation. We find in Section 3 that this can non-trivially affect the result quality of aggregation tasks, like summarization. As Figure 10 shows, sem_partition_by takes a function, which outputs a group-id for each row. LOTUS natively supports a semantic cluster function, which takes the number of clusters to create and a column, but users can also specify arbitrary partition functions. The group-ids output by the partition function indicate to the system which rows should be grouped together, in a best-effort manner, during LM invocations. The system will aggregate over documents within each group, before merging intermediate results across groups.
Sem_map performs a natural language projection over an existing column and outputs a new column in the table. As shown in Figure 11, the user’s langex specifies what to project. This operator provides general functionality and simulates logically row-wise data-flow similar to prior works that use LLM-UDFs in relational languages (Liu et al., 2024a, c). These semantic projections are broadly useful for a variety of tasks, such as row-wise summarization, classification, and entity-extraction.
Sem_extract provides similar functionality to sem_map but provides the answers by returning a list of sub-strings from the source text. This functionality is useful for applications, such as entity extraction or fact-checking, where snippet-finding or verified quotes may be preferable to synthesized LLM-based answers. As Figure 12, the user’s langex signature specifies a projection, similar to sem_map.
Sem_index generates a semantic similarity index over the specified data column, as shown in Figure 13. To generate the semantic index, users first use the settings module to declare a retrieval model, which will be used for generating semantic embeddings and indexing the column data. The sem_index operator takes a column and a local directory path, to which the generated semantic index will be stored. As the figure shows, users can separately index multiple columns of the table.
Load_sem_index re-loads the stored semantic index upon reading the table data from disk, as shown in Figure 14. The user specifies the column corresponding to the semantic similarity index and the directory where the index was saved.
Sem_search performs a top- semantic similarity search over a semantically-indexed column, as Figure 14 shows. The user specifies the table column to search over, a query in natural language, a target of the number of results to return, and optionally indicates whether to return similarity scores as a column in the returned table.
LOTUS also exposes advanced relevance-based re-ranking functionality for search. Users can specify a re-ranker model, as shown in Figure 15, and then set the n_rerank parameter during the semantic search. The semantic search in this case will first find the top- most relevant documents according to the retriever model, and then re-rank the top- found documents to return the top n_rerank ones using the re-ranker model.
Sem_cluster_by assigns a group to each row of the table according to semantic similarity. This operator is akin to a relational group_by, but uses semantic similarity to determine the groups, rather than equality over the specified column. As Figure 16 shows, the user specifies an indexed column to cluster, and the number of clusters to create. The returned table augments the input table with an additional column that specifies the assigned cluster_id of each row. Users can optionally also return the similarity scores between each row and its cluster centroid in an additional column of the returned table.
3. Implementing and Optimizing Semantic Operators
Semantic operators create a rich design space, allowing for a diverse set of algorithmic decisions and optimizations which have significant consequences on system efficiency and accuracy. We identify multiple possible algorithms for these operators, noting their performance implications. LOTUS’ query engine aims to automatically exploit this rich design space to provide an optimized implementation of each individual operator and composite pipelines.
Specifically, LOTUS employs a series of novel algorithms designed to leverage the semantics of each operator in order to maximize parallel batched-inference opportunities, use model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) with a lightweight scoring function unique to our setting, leverage the semantic similarity index, and perform algorithmic approximations for expensive operations (e.g. joins). We see in Section 4 that these decisions can significantly improve efficiency while maintaining high quality results for LOTUS’ query pipelines.
While our initial investigation studies several operator-specific design decisions unique to the semantic bulk processing setting, we envision a rich set of additional optimization opportunities. Several prior works demonstrate performance gains in both logical query plan optimizations (e.g. operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g. code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)), which we find to be promising opportunities for LOTUS’ implementation, left to future work.
LOTUS’ current implementation extends the Pandas API (pan, [n. d.]). We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) for efficient vector similarity search.
3.1. sem_filter
LOTUS’ semantic filter runs batched LLM calls over a set of rows, prompting the model to output a boolean value for each row to indicate whether the record passes the user’s natural language predicate. Figure 18 shows an example of the operator-specific filtering instruction, which is managed transparently by the system.
3.1.1. Optimizations
We leverage model cascades, using an efficient scoring function unique to our setting, to further optimize the filter function. Many prior works have studied model cascades for ML tasks (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022), leveraging a relatively cheap and inaccurate proxy model, along with an accurate but expensive oracle model in order to reduce the inference cost by routing easy queries to the small model, and resorting to the large model where necessary. Typically, a scoring function provides confidence scores over the proxy model’s outputs, and allows the system to decide when to route examples to the larger model. Applying this paradigm to LMs introduces several challenges, including generating a reliable scoring function. Several works study this problem, suggesting to fine-tune a smaller LM to score each question along with the answer produced by the weak LLM (Chen et al., 2023), or invoking multiple sampling paths from the small LM and evaluating self-consistency to build a confidence score (Yue et al., 2024). While these methods provide general-purpose scoring functions, LOTUS’ scoring function leverages the operator semantics to provide a lighter-weight scoring function, which avoids the execution time penalty of scoring with an additional model or re-sampling from the proxy-model.
Specifically, LOTUS’ cascade scoring function leverages the binary-label output of each LM invocation. The system first batch-processes each record using the cheaper LM, then computes confidence scores by exponentiation the log-probabilities for the output LLM tokens corresponding to the True or False answer.
For records with confidence scores below the user-defined threshold, the system then passes these records in batch to the larger LM. As we show, in Section 4, leveraging model cascades in this way for semantic filtering can significantly improves the system’s efficiency while maintaining high quality results for some tasks.
3.2. sem_topk
Performing a semantic top- ranking requires logically reasoning across rows, entailing joint reasoning over often large amounts of data. Implementing an efficient algorithm introduces many important design decisions pertaining to managing LM context length limits, grouping documents within each LM context window to achieve high quality LM-based comparisons over subsets of the data, and choosing an efficient top-k ranking algorithm to combine LM-based comparisons. These decisions have notable performance implications on efficiency and result quality, as we show in Section 4. While prior works have studied LM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n. d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, LOTUS’ implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low query execution time. We leverage several algorithmic design decisions and optimizations to achieve this.
3.2.1. Algorithms
Prior LM-based sorting and ranking works suggest several methods for performing LM-based comparisons. Specifically, prior works primarily study three classes of methods: point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), and pair-wise ranking methods (Shah and Wainwright, 2016; Qin et al., 2024). Point-wise methods output a relevance score for each document independently; while these methods are fully parallelizable across rows, prior work (Desai and Durrett, 2020) shows they provide poor accuracy due to difficulty in calibrating scores across prompts. We note that these methods can be implemented in LOTUS with a sem_map, but are not the focus of our semantic top- operator due to their quality limitations. By contrast, list-wise methods feed a partial list of 10 to 20 documents to the LM and prompt it to output a ranking over these, which can be aggregated using a sliding window approach. Unfortunately, prior works shows these methods are often prone to prediction failures (Sun et al., 2023) and sensitive to document ordering in the prompt (Qin et al., 2024).
Pairwise-prompting methods instead offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label. This method has been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
To aggregate pairwise comparisons, we consider several ranking algorithms, including a quadratic sorting algorithm, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm. The quadratic ranking algorithm obtains a comparison between each pair of input documents, and uses a win rate to determine a ranking over all elements before selecting the best. In contrast, the heap-based algorithm maintains a heap of size , and makes a linear pass over the data, updating the heap when more promising elements are found. Each time a new element is inserted or removed from the heap, the algorithm performs a series of sequential LM comparisons to update the data structure. Lastly the quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot. Because each round is fully parallelizable, we perform these LM-based comparisons efficiently in batch before recursing in the next round.
3.2.2. Discussion
As the core LM-based computation unit over subsets of the data, LOTUS’ current implementation uses pair-wise rankings, following recent prior work (Qin et al., 2024), which demonstrates its effectiveness. We also find this choice useful for implementing further optimizations, such as model cascades, described below. Figure 19 shows an example task-instruction prompt, which we use to implement the LM-based comparison by instructing the model to select between two documents according to the user-defined sorting criteria.
Additionally, for aggregating comparisons, LOTUS’ current implementation leverages the quick-select-based (Hoare, 1961) top-k ranking algorithm, by default. Prior work (Qin et al., 2024) studies several alternative algorithms, including the heap-based ranking implementation and the quadratic sorting algorithm. We study these alternatives in Section 4, and find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for batched calls leading to lower execution time compared to a heap-based implementation. We believe future work may leverage these multiple algorithms to allow the user to declaratively trade-off between performance metrics, like accuracy, execution time, and cost.
3.2.3. Optimizations
LOTUS’ top- algorithm is amenable to several optimizations. First, the top- algorithm’s pair-wise comparisons each invoke the LM to output a binary label, which is amenable to model cascades. To implement this, we use a simple an efficient scoring procedure based on log-probabilities of the generated tokens. We describe this procedure above for sem_filter (Section 3.1.1) and apply an equivalent procedure here. In Section 4, we demonstrate this optimization allows programmers to leverage small, cheap models in conjunction with large, expensive models to obtain high accuracy results at reduced cost.
Additionally, LOTUS can leverage the semantic index to optimize pivot selection for some queries, rather than resorting to random pivot selection. This optimization is useful when there exists correlation between the rankings imposed by the user’s arbitrary sorting criteria and the rankings imposed by semantic similarity scores. In this case, LOTUS can sort the document set based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot. This can reduce the number of LM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss. We believe fruitful future work will automatically estimate correlation between semantic similarity scores and the user-defined ranking criteria to transparently apply this optimization.
3.3. sem_join
Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left and right table. Implemented naively, this can be prohibitively expensive, incurring a quadratic number of LM calls with respect to to the size of the left and right join tables. LOTUS thus implements several algorithms suitable for a variety of settings, highlighting the rich design space, which we plan to further optimize over in future work. Here we describe the design patterns currently implemented, which we find to be useful for real-world tasks that we evaluate in Section 4.
3.3.1. Algorithms
The first join algorithm implements the nested-loop join pattern with efficient LM batch processing to maximize GPU utilization rather than naively looping over each pair of rows and invoking the LM. This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively. As Figure 18 shows, each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate. This quadratic join algorithm is suitable for small join tables.
Alternatively, we implement a map-search-filter join pattern which can be used to approximate the quadratic join algorithm, while incurring fewer LM calls. This algorithm first performs a semantic mapping over the left join key to the domain of the right join key. In the example provided by Figure 20, which joins paper abstracts with the datasets they use, the map step would invoke the LM over each abstract, instructing the model to output the dataset used. This step is un-grounded in the sense that the LM is not given knowledge of right join table, but may optionally leverage user-specified demonstrations. Following the semantic projection, the algorithm then leverages the semantic index and performs a similarity search over the right join table to find candidates likely to pass the predicate. In the example, this search would retrieve a list of datasets from the right table for each paper, using semantic similarity of the datasets to the LM projection output in the first step. The semantic similarity search is parameterized by , which is automatically set based on the user’s specified LM call budget. Lastly, the algorithm performs a filter over the candidate pairs and outputs the joined table of tuples pairs that pass. This join algorithm partially mimics the information flow of prior work (D’Oosterlinck et al., 2024) and abstracts away intermediate steps by instead allowing users to specify an LM call budget for the join. The LM call complexity of this algorithm is .
Lastly, we implement a search-filter join pattern. This algorithm first performs a semantic similarity join using a similarity index on the right join key, with the left join key embedded on the fly. In the example provided by Figure 20, this step would retrieve semantically similar datasets from the right table for each paper abstract in the left table. The batched semantic-similarity search sets the search parameter according to the user’s specified LM call budget, similar to the map-search-filter algorithm. The search results provide candidate tuple pairs, which are then evaluated in batch using the filter operation to gather the final set of rows that pass the user’s language predicate. This algorithm obtains an LM call complexity of .
3.3.2. Discussion
These join patterns offer performance tradeoffs suitable for different settings. The nested-loop algorithm offers an efficient solution when the join tables are sufficiently small, whereas the map-search-filter and search-filter patterns are suitable approximations that can apply efficiently over large tables, due to their linear LM call complexity in table size. The expected result quality of either approximation likely varies depending on the presence of predicate clustering and correlation, as defined by prior work (Patel et al., 2024) between the user’s predicate and the semantic embeddings. The search-filter pattern is likely to produce high quality results under positive query correlation, where entities that are semantically similar are also likely to pass the predicate. On the other hand, the map-search-filter pattern is likely to produce higher quality results when the predicate-embedding correlation is low using the original left join key’s embeddings as queries, but can be increased using the LM-projection over the left join key.
3.4. sem_agg
Performing semantic aggregations are inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows. Thus, the operator’s implementation must efficiently orchestrate the LM over large amounts of data, while managing long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length. LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results and provide the programmer with flexibility.
3.4.1. Algorithms
LOTUS’ implementation builds on and generalizes the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n. d.]; lan, [n. d.]). These implementations primarily leverage one of two aggregation patterns: either a fold pattern, which produces a sequential, linear pass over the data while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains.
3.4.2. Discussion
By default, LOTUS’ aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024). However, LOTUS’ partition function can be used to override the default functionality and generalizes the information flow of prior implementation by allowing users to specify arbitrary groupings over the data. Each unique group specified by the partition function will be aggregated with as few LM invocations as possible according to the model context length, before being merged with other groups. This allows the user to perform a fold pattern, as well as arbitrary user-defined patterns.
3.4.3. Optimizations
Preliminary results show that achieving high-quality semantic-aggregations is non-trivial and sensitive to the document’s input ordering and grouping. We see qualitative evidence of this in a summarization task of 50 ArXiv paper abstracts, which we show in Figure 21. We contrast the summarization results of a naive semantic aggregation (Figure 22) performed over the paper abstracts with the semantic aggregation performed using a partitioning function based on semantic similarity (Figure 23). The semantic aggregation using the partitioner first clusters documents based on semantic similarity, then aggregates each cluster before performing aggregations across clusters. As the figures show the two methods result in significantly different summaries. While the latter demonstrates qualitatively more cohesion and effectively abstracts general themes across papers, the former omits details and tends to list low-level details from individual papers rather than capturing higher-level themes. We leave a quantitative study of this to future work and believe that semantic aggregations create a rich design space for optimization.
3.5. sem_map & sem_extract
Both LOTUS’ semantic map and extract run batched LM calls over a set of rows, prompting the LM with the user’s arbitrary natural-language expression to generate a new column. Both functions can be fully parallelized over rows, and LOTUS implements efficient batched inference with vLLM (Kwon et al., 2023). To implement sem_extract, LOTUS prompts the model to answer the user’s langex with direct quotes, and the system then verifies that these snippets returned by the LM match the reference text.
3.6. Semantic Indexing
LOTUS supports several different algorithms for its semantic index.
To generate the semantic similarity index, LOTUS’ sim_index operator first batch processes the user-specified column to generate semantic embeddings with the configured retriever model, then constructs a vector index over the semantic embeddings. The current implementation uses FAISS’ flat index by default and writes the index locally to disk. User’s can specify the local file path to store the index, and optionally specify alternative vector indices, such as hierarchical navigable small worlds (HNSW) (Malkov and Yashunin, 2018), inverted indices (IVF) (Baranchuk et al., 2018; Ge et al., 2014; Jégou et al., 2011; Johnson et al., 2017), and locality sensitive hash indices (LSH) (Andoni and Indyk, 2008; Andoni and Razenshteyn, 2015; Indyk and Motwani, 1998; Jafari et al., 2020; Li et al., 2020; Liu et al., 2021; Lu and Kudo, 2020; Park et al., 2015; Sundaram et al., 2013; Zheng et al., 2020; Andoni et al., 2015; Gionis et al., 1999; Gong et al., 2020).
In the future, we envision additionally supporting a wider variety of embeddings stores and indices (Patel et al., 2024; Wang et al., 2021; vec, [n. d.]; noa, 2023b; chr, [n. d.]; mos, [n. d.]).
LOTUS’ semantic search, similarity join and cluster operations all leverage the similarity index in their implementation. sem_search first embeds the user’s query string using the configured retriever model, then performs an efficient top-k search using the loaded FAISS index over the search column. sem_sim_join similarly performs a top-k search using the loaded FAISS index over the right key. Here the right-key may be a column of natural-language text, which LOTUS will embed using the retriever model on-the-fly, or an indexed column, for which LOTUS will load the previously-generated embeddings. The right-key embeddings then serve as the queries to perform batched search over the search column, given by the left key. Lastly, LOTUS implements sem_cluster_by using FAISS optimized kmeans library to cluster the user-specified column. Here, the column must be previously indexed, and LOTUS uses the generated embeddings to perform the clustering with the user-specified number of cluster centroids.
4. Evaluation
We now evaluate LOTUS’ programmability and efficiency through three diverse applications: fact-checking (Section 4.1), extreme multi-label classification (Section 4.2), and search and ranking (Section 4.3). For each of these applications, we see that state-of-the-art quality results are achievable with low development overhead, using LOTUS programs with a few lines of code and a few or even one semantic operator. In addition, our results demonstrate interesting implementation and optimization choices introduced by semantic operators. Specifically, we find that:
On the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce FacTool (Chern et al., 2023), a recent state-of-the-art pipeline, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time.
On the BioDEX dataset (D’Oosterlinck et al., 2023) for the extreme multi-label classification task, LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that runs up to faster than a naive join, demonstrating the power of LOTUS’ declarative interface.
On the SciFact dataset (Thakur et al., 2021) and two newly-constructed paper datasets for the search and ranking application, LOTUS’s semantic top- operator achieves higher nDCG@10 than a vanilla retriever and re-ranker, while also running faster than alternative LM-based ranking methods used by prior works (Qin et al., 2024).
Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama 3 models (int, [n. d.]), with a batch size of 64 running on vLLM (Kwon et al., 2023). For our experiments that use OpenAI’s GPT models (ope, [n. d.]), we run with 64-way thread parallelism.
4.1. Application: Fact-Checking
4.1.1. Dataset
We evaluate on FEVER (Thorne et al., 2018), a claim verification dataset. We use the development dataset, which contains about 38,000 total claims, of which we sample 500 for our evaluation. Each claim is labeled with one of three labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”, and the task is to correctly determine the label of each claim, leveraging evidence from a corpus of million Wikipedia articles. We merge the latter two labels in to a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation.
4.1.2. Baselines
FacTool (Chern et al., 2023) is a recent research work that proposes a multi-step pipeline for fact-checking involving, claim extraction, query generation, tool querying, evidence collection, and verification. We use FactTool’s open source codebase (gai, [n. d.]) to measure its performance. FactTool’s pipeline, by default, performs retrieval with a Google Search API (cus, [n. d.]). We evaluate the pipeline with both the default retrieval API, and alternatively test with a ColBERT (Khattab and Zaharia, 2020) index over the document corpus to perform retrieval. We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the implemented LOTUS programs.
| Method | Accuracy | Execution Time, batched (s) | Execution Time, no batching (s) | LoC |
| FacTool | 83.5 | N/A | 1,174.8 | ¿ 750 |
| LOTUS-Factool | 90.0 | 111.48 | 979.50 | ¡ 50 |
| LOTUS-fact-filter | 90.5 | 64.28 | 206.45 | ¡ 50 |
| LOTUS-fact-filter (+ cascades) | 93 | 34.09 | 168.24 | ¡ 50 |
| LOTUS-fact-join | 86.5 | 2,394.22 | 12,923.08 | ¡ 50 |
4.1.3. LOTUS Programs
We compose several intuitive LOTUS programs, each in less than 50 lines of code. For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama 70B as the primary LM, and Llama 8B and TinyLlama for cascade optimizations.
First, we compose a pipeline designed to directly re-implement FacTool’s information flow in LOTUS. Figure 24 shows the pseudocode for the LOTUS-FacTool pipeline. The two tables are shown by wiki_df, which stores the Wikipedia articles, and claim_df, which stores the claims from the FEVER dataset. After loading the semantic index to wiki_df, the pipeline first performs a semantic map over each claim to generate two search queries, which are then used to perform a semantic similarity join over the corpus of Wikipedia articles. The program then concatenates the context retrieved for each claim, and performs a sem_map to output whether the claim is true or false, along with a revised claim if the claim is false. We use the same prompts found in FacTool (gai, [n. d.]), which include 3 demonstrations for generating search queries in the first sem_map, and chain-of-thought prompting in the second sem_map.
The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the query pipeline, with LOTUS’ semantic filter operation. We use 3 demonstrations for the filter. Figure 25 shows the pseudocode for this program. We evaluate this program with and without model cascades for the semantic filter operation.
Lastly, we compose an alternative pipeline, LOTUS-fact-join. As the pseudocode shows in Figure 26, the pipeline first performs a semantic map over each claim to obtain a set of sub-claims. From this, we create the claimed_facts_df, which separates each sub-claim into a different row. Next, the pipeline uses these sub-claims to perform a semantic similarity join over the Wikipedia corpus, then performs a semantic map over each retrieved article to generate the important facts in each one. Lastly, the program performs a semantic join between the sub-claims and the facts described in the retrieved passages. If the returned table contains a supporting fact for each sub-claim, the claim is labeled as ”Supported”. For the sem_map and join operations, we use 3 demonstrations each.
4.1.4. Results
Table LABEL:tab:factchecking demonstrates the powerful abstraction that LOTUS provides, allowing programmers to quickly write and test programs that compose simple operators to obtain state-of-the-art results. We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching. We see that FacTool’s implementation offers strong accuracy performance on the FEVER datasets, however the full repository required several hundred lines of code, highlighting the development burden of building these applications without abstractions for semantic-bulk processing. By contrast, each LOTUS program offers comparable or higher accuracy, in relatively few lines of code. We also note that FacTool implements its method without batching, whereas LOTUS, by default, leverages batched LM execution for efficiency. To provide an apples-to-apples comparison, we compare FacTool’s un-batched implementation to the LOTUS programs both with and without batching.
First, we see from the Table LABEL:tab:factchecking that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with 6.5 points higher accuracy and lower execution time without batching. The LOTUS-FacTool implementation with batching further decreases execution time compared to the original FacTool implementation by . We see that the next LOTUS pipeline simply changes a single operation, switching the sem_map to a sem_filter, and maintains similar result quality to LOTUS-FacTool while further reducing execution time by in the batched implementation.
Leveraging LOTUS’ filter operation allows the programmer to further optimize the program using model cascades, which increases accuracy by 3 points and increases the batched execution time by compared to LOTUS-FacTool. Figure 27 highlights the diverse performance trade-offs presented the model cascade optimizations used in LOTUS’ filter sem_filter. The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, Llama 70B, or TinyLlama (Zhang et al., 2024), shown by the circles. We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, or TinyLlama and Llama70B to implement the filter casacade. We generate multiple cascade points, shown by the stars, by varying the confidence threshold used. The plot shows that this filter optimization can substantially reduce execution time, and offer diverse accuracy trade-offs, compared to implementing the pipeline with the oracle model, Llama 70B, alone.
Returning to Table LABEL:tab:factchecking, we find that the last LOTUS pipeline (search-map-join), reduces accuracy and increases execution time. Notably the performance trade-offs of different LOTUS programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines. Each LOTUS program we described can be implemented easily in relatively few lines of code, and the best performing one offers higher accuracy and lower execution time than FacTool’s original, un-batched implementation.
| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
| Semantic Similarity Join | 0.106 | 0.120 | 2.91 | 0.00 |
| LOTUS Semantic Join (map-search-filter pattern) | 0.241 | 0.258 | 2,762 | 7,750 |
| LOTUS Semantic Join (nested-loop pattern) | N/A | N/A | 2,144,560* | 6,092,500 |
| LOTUS Semantic Join (search-filter pattern) | 0.155 | 0.186 | 2,640 | 7,500 |
Estimated under linear-scaling assumption in number of batch calls.
4.2. Application: Extreme Multi-label Classification
4.2.1. Dataset
We evaluate on the Biodex Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and expert-created drug safety reports constructed from each article. The task is to correctly label the drug reactions experience by the patient in each medical article. We sample 250 patient articles for our evaluation. Notably, there are approximately 24,000 possible drug-reaction labels, making this task an extreme multi-label classification task. Due to the large number of possible labels, leveraging an LM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024). We show below that this task can be efficiently modeled and programmed using LOTUS’ semantic join.
4.2.2. Baselines
We consider a simple retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels. We show the pseudocode for this program in Figure 28.
4.2.3. LOTUS Programs
The proposed LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 32. We perform the semantic join using at most 7 map demonstrations and set an LM call budget of 10,000. We use Llama-70b as the LM.
4.2.4. Results
Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while maintaining query efficiency. The table reports the rank-precision@5 (RP@5), rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls required for each program. Since the nested-loop join pattern is prohibitively expensive to run, we show an estimate of the execution time assuming linear scaling in the number of batched calls.
First we compare performance of the LOTUS join program implemented with the map-search-filter pattern to the baseline method, the semantic similarity join. As expected, the LOTUS program offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the retrieval-based similarity join. This highlights the effectiveness of leveraging LMs over the data for complex reasoning-based tasks. We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4 turbo. D’Oosterlinck et al. report 24.73 RP@5 and 27.67 RP@10 for the compiled program, representing comparable result quality to LOTUS’ program, although notably the LOTUS program was not compiled with a prompt optimization system.
We now consider several interesting performance trade-offs presented in the semantic join algorithm. We compare LOTUS map-search-filter join pattern to the naive nested-loop join algorithm and the search-filter join pattern shown in Table 3. We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is untenable and prohibitively costly. By contrast, the search-filter and map-search-filter pattern substantially reduce the LM call budget of the naive algorithm by , using an approximation. While these two approximation patterns have similar efficiency, according to execution time and the number of LM calls, interestingly, they offer substantially different result quality. Specifically, the map-search-filter pattern offers higher RP@5 and higher RP@10 compared to the retrieve-filter pattern on this task. These results highlight the unique opportunities bulk-semantic processing pipelines present for designing new algorithms and optimizations.
4.3. Application: Search & Ranking
4.3.1. Dataset
We evaluate LOTUS’ performance on the search and ranking task using three datasets, including BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, as well as two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generate to evaluate more complex, reasoning-based ranking criteria over the data. For each, we report nDCG@10, as well as the execution time (ET) in seconds.
The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance given each scientific claim. We sample 300 scientific facts for our evaluation, and report the average ranking execution time across these samples.
While the SciFact dataset provides a sorting task based on a simple relevance criterion, our newly proposed benchmarks provide a more complex sorting criteria over a corpus of paper abstracts. Specifically, the ranking task is to find the papers that report the highest accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench respectively.
To generate CIFAR-bench, we took 100 abstracts from the Papers with Code Dataset (pap, [n. d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain the top-10.
We then synthetically generated HellaSwag-bench by prompting Llama-70B to create 200 paper abstracts, each with a specified accuracy value, randomly sampled from . This setup allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth. We note that an alternative approach to these tasks could efficiently leverage sem_map to extract accuracy values on abstracts from either dataset, then perform a structured sort. However, our evaluation focuses on assessing the semantic ranking capabilities of LOTUS’ top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may likely generalize to a wider set of reasoning-based sorting queries, such as ”which paper makes the most outrageous claim”, for which ground truth is less objective to evaluate. For CIFAR-bench and HellaSwag-bench we report results for trials of the ranking task, at temperature , similar to prior works (Khattab et al., 2023)
| Method | nDCG@10 | Execution Time (s) |
| Semantic Search | 0.712 | 0.009 |
| Semantic Search + Reranker | 0.741 | 2.64 |
| LOTUS Top-k (quickselect + sem-index) - Llama-70B | 0.775 | 33.6 |
| LOTUS Top-k (quickselect + sem-index) - GPT-4o | 0.800 | 11.2 |
| Method | CIFAR | HellaSwag | ||
| nDCG@10 | Execution Time (s) | nDCG@10 | Execution Time (s) | |
| Semantic Search | 0.252 | 0.008 | 0.119 | 0.008 |
| Semantic Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
| LOTUS Top-k (quickselect) - Llama 70-B | 0.746 | 41.3 | 0.909 | 59.1 |
4.3.2. Baselines
We consider two simple baselines. The first baseline performs semantic search, as shown in Figure 30, using the E5Model (Wang et al., 2024) for retrieval. The second baseline performs search with re-ranking, as the pseudocode shows in Figure 31, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking.
4.3.3. LOTUS Programs
The proposed LOTUS program performs a semantic top- over the documents, as shown in Figure 32, which shows example pseudocode for the CIFAR-bench dataset. The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets, as shown in the figure. We note that for the SciFact dataset, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk. We report results using both Llama-70B and GPT-4o as the LM.
4.3.4. Results
| Method | Scifact | CIFAR | HellaSwag | ||||||
| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
| Quadratic Sort | 0.836 | 712 | 4950 | 0.868 | 634 | 4950 | 0.966 | 1,803 | 19,900 |
| Heap Top-k | 0.776 | 65.0 | 216 | 0.832 | 99.6 | 350 | 0.907 | 98.9 | 415.2 |
| QuickSelect Top-k | 0.776 | 42.4 | 285 | 0.746 | 41.3 | 303.95 | 0.909 | 59.1 | 620.95 |
| QuickSelect Top-k + Semantic Index | 0.775 | 33.6 | 229 | 0.710 | 39.6 | 307.7 | 0.975 | 63.6 | 672.7 |
Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- operator for complex search tasks. In addition, Table 6 and Figure 33 highlight the rich implementation design space that this task presents. We walk through several note-worthy findings.
We first turn our attention to Table LABEL:tab:ranking_scifact, which presents the results for each bench-marked method on the SciFact dataset, which uses a relevance-based sorting criterion. As expected, both semantic search programs with and without re-ranking present strong baselines. Specifically, the re-ranker model, which is a supervised model trained specifically for the task of relevance-based ranking, increases nDCG@10 by 3 percentage points, while trading off query efficiency compared to the semantic search baseline. Notably, the unsupervised LM-based LOTUS programs outperform the supervised re-ranker’s result quality. The table shows LOTUS semantic top-k program with Llama-70B and GPT-4o, which outperform the semantic search baseline by 6 and 9 percentage points respectively. The LOTUS programs with Llama-70B and GPT-4o also outperform the re-ranker by 3 and 5 points respectively, improving upon the quality of the supervised baseline. As expected, the improved result quality comes at a trade-off to query efficiency due to the cost of LM calls. Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipeline and trade-off result quality and efficiency depending on application-specific requirements.
Turning our attention to Table 5, we study LOTUS’ generality in supporting arbitrary language-based ranking criteria over the dataset. On the CIFAR-bench and HellaSwag-bench datasets, which use a complex sorting criteria, we see that both semantic search baselines with and without re-ranking provide poor result quality with consistently low nDCG@10. The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with 49.4 and 44.8 points higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively. These accuracy gains reflect the powerful reasoning capabilities of LMs efficiently orchestrated over the data. As expected, these significant accuracy gains come at an increase to execution time.
We now analyze the efficiency of LOTUS semantic top- implementation along with it’s proposed optimizations. Table 6 compares several semantic top- algorithms, namely a quick-select top- algorithm, a quick-select top- that leverages the similarity index for pivot selection, a heap-based top- algorithm, and a quadratic sorting algorithm. First, we see that the quadratic algorithm, which performs an LM comparison between each pair of input documents, offers consistently high result quality across each dataset. However, this method is prohibitively expensive, requiring more LM calls and over higher execution time than the alternative implementations. The heap top- and quick-select top- methods offer comparable result quality, but with interesting trade-offs in query efficiency. Notably the quick-select top- method offers lower execution time than the heap-based sorting method across all datasets, despite requiring more LM calls in some cases. This is because the quick-select top- implementation allows for efficient batch-processing in each round of the algorithm, whereas the heap-based top- incurs sequential LM calls during heap updates. For this reason, our current implementation leverages the quick-select top- algorithm, although we envision future iterations may leverage multiple algorithms and allow the user to declaratively trade-off accuracy, query throughput, and cost.
In addition to providing an efficient top- algorithm, the table also demonstrates the use of LOTUS’ similarity index for optimizing top- query performance. The quick-select top- algorithm optimized with the semantic similarity index for pivot selection demonstrates lower execution time at no accuracy loss on SciFact, where the ranking criteria correlates likely with semantic similarity. On the other hand, for the CIFAR-bench and HellaSwag-bench datasets, where the ranking criteria does not correlate with semantic similarity, we see that the similarity index has no significant impact on the accuracy or efficiency top- performance.
Lastly, we analyze the impact of LOTUS’ cascade optimization applied to the quickselect top- implementation. Figure 33 compares the nDCG and execution time of implementing the operator with a single model, either Llama 8B or Llama 70B, to the implementation that leverages these models together using model cascades. We vary the confidence threshold of the cascade to generate several points in the trade-off space. We find that the cascade optimization offers diverse performance trade-offs, which can outperform the single-model oracle baseline, which uses Llama-70B. For instance, one cascade along the Pareto-frontier improves accuracy of the Llama-70B baseline by , while reducing execution time by , demonstrating substantial opportunities for automatically optimizing LOTUS’ semantic query pipelines.
5. Related Work
Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks or applications.
Palimpzest (Liu et al., 2024b) presents a declarative approach to data cleaning and extract-transform-load (ETL) tasks. The authors propose to automatically optimize relational operators with LM calls, and implement two row-wise relational operators, a newly proposed convert operator, which can be transparently optimized to perform entity extraction using LLMs, and an AI-based filter operation, logically similar to LOTUS’ sem_filter. The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis to implement user queries, and proposes several others as future work.
SUQL (Liu et al., 2024c) presents a SQL extension to support conversational agents with knowledge grounding over structured and unstructured data. Specifically, the system extends SQL with two new logically row-wise operators, answer, which prompts an LM to answer the user question over each row, and summary, which prompts an LM to provide a summary to the user over each row. The system can provide automatic optimizations, such as predicate re-ordering using a lazy evaluation approach and proposes to use retrieval to optimize some answer queries.
ZenDB (Lin et al., 2024) and EVAPORATE (Arora et al., 2023) tackle the task of automatically ingesting and extracting semi-structured documents into structured tables that can be queried using standard relational operators and languages. ZenDB extracts structure using a semantic hierarchical tree index, which is integrated with a SQL query engine to support efficient query processing over the extracted attribute values. The systems implements several optimizations, including predicate reordering, push-down, and projection pull-up. Additionally, EVAPORATE performs efficient entity extraction from semi-structured data using LM-based code generation and weak supervision to ensemble candidate functions.
LOTUS, in contrast to these prior works, defines a general-purpose programming model designed to capture broad-ranging applications with a core set of composable semantic operators, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions. In LOTUS’ current implementation, users can use sem_map to perform entity extraction over unstructured text fields, although future work may provide native functionality for entity extraction by integrating systems or optimizations of prior work. Additionally, we believe that several optimizations leveraged in prior work, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
LLM UDFs
Recent research work (Liu et al., 2024a) and existing analytical database vendors, such as
Google BigQuery (ver, [n. d.]),
Databricks (dat, [n. d.]) and
AWS Redshift (noa, 2023a), alternatively offer LLM user-defined functions (UDFs) to the programmer. The LLM UDF programming model provides a lower-level interface, which is limited to logically row-wise LLM execution over the data, equivalent to LOTUS’ sem_map. In contrast, LOTUS’ programming model is declarative and provides a rich set of semantic operators, allowing the system to automatically orchestrate the LM to serve a variety of complex query patterns, including aggregations, ranking, and joins.
Liu et al. (Liu et al., 2024a) study how to optimize LLM UDF functions, demonstrating performance gains with a de-duplication method and a prefix-sharing maximization method that reorders rows and re-formulates parameterized prompts to maximize key-value (KV) cache reuse during query execution. We believe these methods could be effective optimizations in future work at LOTUS’ batched execution layer.
LM Programming Frameworks Recent LM programming frameworks, such as LangChain (lan, [n. d.]), LlamaIndex (lla, [n. d.]), and DSPy (Khattab et al., 2023), have gained significant popularity. These systems provide a set of abstractions for programming with LMs, including utilities for handling prompts and post-processsing LM outputs, and support for common use cases, such as RAG, chat-bots, and function-calling. DSPy focuses on abstracting LM pipelines as programs and provides automatic prompt optimization. In contrast to these systems, LOTUS’ programming model is designed for tasks involving bulk processing data with LLMs. While some of these systems support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized.
ML-based Query Processing Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique opportunities for system design and optimization. MADLib (Hellerstein et al., 2012) extends SQL with efficient abstractions for supervised learning, unsupervised learning, and descriptive statistics. Prior works such as NoScope (Kang et al., 2017), TASTI (Kang et al., [n. d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML models over large datasets, typically in video analytics. Some optimizations, such as model cascades and predicate re-ordering, which were useful in these works are likewise useful for optimizing LOTUS pipelines with language models. However, our setting with natural-language reasoning tasks has new operators with significantly different semantics, such as sem_topk and sem_agg, requiring a new programming model as well as new query execution algorithms and optimizations.
6. Conclusion
In this work, we proposed semantic operators to provide the first declarative and general-purpose interface to serve bulk-semantic processing. We implement these operators in the LOTUS system to seamlessly extend the relational model and allow programmers to easily compose powerful reasoning-based query pipelines over vast corpora of structured and unstructured data. Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of LOTUS’ programming model as well as the efficiency and optimization opportunities of LOTUS’ query engine. For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized to achieve higher performance than existing implementations.
Acknowledgements.
This research was supported in part by affiliate members and other supporters of the Stanford DAWN project, including Meta, Google, and VMware, as well as Cisco, SAP, and a Sloan Fellowship. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.References
- (1)
- dat ([n. d.]) [n. d.]. AI Functions on Databricks. https://docs.databricks.com
- arx ([n. d.]) [n. d.]. arXiv.org ePrint archive. https://arxiv.org
- chr ([n. d.]) [n. d.]. Chroma. https://www.trychroma.com/
- cus ([n. d.]) [n. d.]. Custom Search JSON API | Programmable Search Engine. https://developers.google.com/custom-search/v1/overview
- dis ([n. d.]) [n. d.]. Discovery Insight Platform. https://www.findourview.com
- gai ([n. d.]) [n. d.]. GAIR-NLP/factool: FacTool: Factuality Detection in Generative AI. https://github.com/GAIR-NLP/factool
- int ([n. d.]) [n. d.]. Introducing Meta Llama 3: The most capable openly available LLM to date. https://ai.meta.com/blog/meta-llama-3/
- lan ([n. d.]) [n. d.]. LangChain. https://www.langchain.com/
- ver ([n. d.]) [n. d.]. LLM with Vertex AI only using SQL queries in BigQuery. https://cloud.google.com/blog/products/ai-machine-learning/llm-with-vertex-ai-only-using-sql-queries-in-bigquery
- mos ([n. d.]) [n. d.]. Mosaic AI Vector Search. https://docs.databricks.com
- ope ([n. d.]) [n. d.]. OpenAI Platform. https://platform.openai.com
- pan ([n. d.]) [n. d.]. pandas - Python Data Analysis Library. https://pandas.pydata.org/
- pap ([n. d.]) [n. d.]. Papers with Code - Machine Learning Datasets. https://paperswithcode.com/datasets
- lla ([n. d.]) [n. d.]. Querying - LlamaIndex 0.9.11.post1. https://docs.llamaindex.ai/en/stable/understanding/querying/querying.html
- vec ([n. d.]) [n. d.]. Vector Database for Vector Search. https://www.pinecone.io/
- noa (2023a) 2023a. Large Language Models for sentiment analysis with Amazon Redshift ML (Preview) | AWS Big Data Blog. https://aws.amazon.com/blogs/big-data/large-language-models-for-sentiment-analysis-with-amazon-redshift-ml-preview/ Section: Amazon Redshift.
- noa (2023b) 2023b. Vespa - the big data serving engine. https://vespa.ai/
- noa (2024) 2024. mixedbread-ai/mxbai-rerank-large-v1 · Hugging Face. https://huggingface.co/mixedbread-ai/mxbai-rerank-large-v1
- Adams et al. (2023) Griffin Adams, Alexander Fabbri, Faisal Ladhak, Eric Lehman, and Noémie Elhadad. 2023. From Sparse to Dense: GPT-4 Summarization with Chain of Density Prompting. http://arxiv.org/abs/2309.04269 arXiv:2309.04269 [cs].
- Anantha et al. (2021) Raviteja Anantha, Svitlana Vakulenko, Zhucheng Tu, Shayne Longpre, Stephen Pulman, and Srinivas Chappidi. 2021. Open-Domain Question Answering Goes Conversational via Question Rewriting. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 520–534. https://doi.org/10.18653/v1/2021.naacl-main.44
- Andoni and Indyk (2008) Alexandr Andoni and Piotr Indyk. 2008. Near-optimal hashing algorithms for approximate nearest neighbor in high dimensions. Commun. ACM 51, 1 (Jan. 2008), 117–122. https://doi.org/10.1145/1327452.1327494
- Andoni et al. (2015) Alexandr Andoni, Piotr Indyk, Thijs Laarhoven, Ilya Razenshteyn, and Ludwig Schmidt. 2015. Practical and optimal LSH for angular distance. In Proceedings of the 28th International Conference on Neural Information Processing Systems - Volume 1 (NIPS’15). MIT Press, Cambridge, MA, USA, 1225–1233.
- Andoni and Razenshteyn (2015) Alexandr Andoni and Ilya Razenshteyn. 2015. Optimal Data-Dependent Hashing for Approximate Near Neighbors. In Proceedings of the forty-seventh annual ACM symposium on Theory of Computing (STOC ’15). Association for Computing Machinery, New York, NY, USA, 793–801. https://doi.org/10.1145/2746539.2746553
- Arora et al. (2023) Simran Arora, Brandon Yang, Sabri Eyuboglu, Avanika Narayan, Andrew Hojel, Immanuel Trummer, and Christopher Ré. 2023. Language Models Enable Simple Systems for Generating Structured Views of Heterogeneous Data Lakes. http://arxiv.org/abs/2304.09433 arXiv:2304.09433 [cs].
- Baranchuk et al. (2018) Dmitry Baranchuk, Artem Babenko, and Yury Malkov. 2018. Revisiting the Inverted Indices for Billion-Scale Approximate Nearest Neighbors. https://doi.org/10.48550/arXiv.1802.02422 arXiv:1802.02422 [cs].
- Braverman and Mossel ([n. d.]) Mark Braverman and Elchanan Mossel. [n. d.]. Noisy sorting without resampling. ([n. d.]).
- Chang et al. (2024) Yapei Chang, Kyle Lo, Tanya Goyal, and Mohit Iyyer. 2024. BooookScore: A systematic exploration of book-length summarization in the era of LLMs. http://arxiv.org/abs/2310.00785 arXiv:2310.00785 [cs].
- Chen et al. (2023) Lingjiao Chen, Matei Zaharia, and James Zou. 2023. FrugalGPT: How to Use Large Language Models While Reducing Cost and Improving Performance. http://arxiv.org/abs/2305.05176 arXiv:2305.05176 [cs].
- Chern et al. (2023) I.-Chun Chern, Steffi Chern, Shiqi Chen, Weizhe Yuan, Kehua Feng, Chunting Zhou, Junxian He, Graham Neubig, and Pengfei Liu. 2023. FacTool: Factuality Detection in Generative AI – A Tool Augmented Framework for Multi-Task and Multi-Domain Scenarios. https://doi.org/10.48550/arXiv.2307.13528 arXiv:2307.13528 [cs].
- Desai and Durrett (2020) Shrey Desai and Greg Durrett. 2020. Calibration of Pre-trained Transformers. https://arxiv.org/abs/2003.07892v3
- D’Oosterlinck et al. (2024) Karel D’Oosterlinck, Omar Khattab, François Remy, Thomas Demeester, Chris Develder, and Christopher Potts. 2024. In-Context Learning for Extreme Multi-Label Classification. https://doi.org/10.48550/arXiv.2401.12178 arXiv:2401.12178 [cs].
- D’Oosterlinck et al. (2023) Karel D’Oosterlinck, François Remy, Johannes Deleu, Thomas Demeester, Chris Develder, Klim Zaporojets, Aneiss Ghodsi, Simon Ellershaw, Jack Collins, and Christopher Potts. 2023. BioDEX: Large-Scale Biomedical Adverse Drug Event Extraction for Real-World Pharmacovigilance. https://doi.org/10.48550/arXiv.2305.13395 arXiv:2305.13395 [cs].
- Drozdov et al. (2023) Andrew Drozdov, Honglei Zhuang, Zhuyun Dai, Zhen Qin, Razieh Rahimi, Xuanhui Wang, Dana Alon, Mohit Iyyer, Andrew McCallum, Donald Metzler, and Kai Hui. 2023. PaRaDe: Passage Ranking using Demonstrations with Large Language Models. https://doi.org/10.48550/arXiv.2310.14408 arXiv:2310.14408 [cs].
- Gao et al. (2023) Luyu Gao, Zhuyun Dai, Panupong Pasupat, Anthony Chen, Arun Tejasvi Chaganty, Yicheng Fan, Vincent Y. Zhao, Ni Lao, Hongrae Lee, Da-Cheng Juan, and Kelvin Guu. 2023. RARR: Researching and Revising What Language Models Say, Using Language Models. https://doi.org/10.48550/arXiv.2210.08726 arXiv:2210.08726 [cs].
- Ge et al. (2014) Tiezheng Ge, Kaiming He, Qifa Ke, and Jian Sun. 2014. Optimized Product Quantization. IEEE Transactions on Pattern Analysis and Machine Intelligence 36, 4 (April 2014), 744–755. https://doi.org/10.1109/TPAMI.2013.240 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
- Gionis et al. (1999) Aristides Gionis, Piotr Indyk, and Rajeev Motwani. 1999. Similarity Search in High Dimensions via Hashing. In Proceedings of the 25th International Conference on Very Large Data Bases (VLDB ’99). Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 518–529.
- Gong et al. (2020) Long Gong, Huayi Wang, Mitsunori Ogihara, and Jun Xu. 2020. iDEC: indexable distance estimating codes for approximate nearest neighbor search. Proceedings of the VLDB Endowment 13, 9 (May 2020), 1483–1497. https://doi.org/10.14778/3397230.3397243
- Hellerstein et al. (2012) Joe Hellerstein, Christopher Ré, Florian Schoppmann, Daisy Zhe Wang, Eugene Fratkin, Aleksander Gorajek, Kee Siong Ng, Caleb Welton, Xixuan Feng, Kun Li, and Arun Kumar. 2012. The MADlib Analytics Library or MAD Skills, the SQL. http://arxiv.org/abs/1208.4165 arXiv:1208.4165 [cs].
- Hoare (1961) C. A. R. Hoare. 1961. Algorithm 65: find. Commun. ACM 4, 7 (July 1961), 321–322. https://doi.org/10.1145/366622.366647
- Indyk and Motwani (1998) Piotr Indyk and Rajeev Motwani. 1998. Approximate nearest neighbors: towards removing the curse of dimensionality. In Proceedings of the thirtieth annual ACM symposium on Theory of computing (STOC ’98). Association for Computing Machinery, New York, NY, USA, 604–613. https://doi.org/10.1145/276698.276876
- Izacard et al. (2022) Gautier Izacard, Patrick Lewis, Maria Lomeli, Lucas Hosseini, Fabio Petroni, Timo Schick, Jane Dwivedi-Yu, Armand Joulin, Sebastian Riedel, and Edouard Grave. 2022. Atlas: Few-shot Learning with Retrieval Augmented Language Models. https://doi.org/10.48550/arXiv.2208.03299 arXiv:2208.03299 [cs].
- Jafari et al. (2020) Omid Jafari, Parth Nagarkar, and Jonathan Montaño. 2020. mmLSH: A Practical and Efficient Technique for Processing Approximate Nearest Neighbor Queries on Multimedia Data. In Similarity Search and Applications (Lecture Notes in Computer Science), Shin’ichi Satoh, Lucia Vadicamo, Arthur Zimek, Fabio Carrara, Ilaria Bartolini, Martin Aumüller, Björn Þór Jónsson, and Rasmus Pagh (Eds.). Springer International Publishing, Cham, 47–61. https://doi.org/10.1007/978-3-030-60936-8_4
- Johnson et al. (2017) Jeff Johnson, Matthijs Douze, and Hervé Jégou. 2017. Billion-scale similarity search with GPUs. http://arxiv.org/abs/1702.08734 arXiv:1702.08734 [cs].
- Jégou et al. (2011) Herve Jégou, Matthijs Douze, and Cordelia Schmid. 2011. Product Quantization for Nearest Neighbor Search. IEEE Transactions on Pattern Analysis and Machine Intelligence 33, 1 (Jan. 2011), 117–128. https://doi.org/10.1109/TPAMI.2010.57 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
- Kang et al. (2019) Daniel Kang, Peter Bailis, and Matei Zaharia. 2019. BlazeIt: Optimizing Declarative Aggregation and Limit Queries for Neural Network-Based Video Analytics. http://arxiv.org/abs/1805.01046 arXiv:1805.01046 [cs].
- Kang et al. (2017) Daniel Kang, John Emmons, Firas Abuzaid, Peter Bailis, and Matei Zaharia. 2017. NoScope: Optimizing Neural Network Queries over Video at Scale. http://arxiv.org/abs/1703.02529 arXiv:1703.02529 [cs].
- Kang et al. (2020) Daniel Kang, Edward Gan, Peter Bailis, Tatsunori Hashimoto, and Matei Zaharia. 2020. Approximate selection with guarantees using proxies. Proceedings of the VLDB Endowment 13, 12 (Aug. 2020), 1990–2003. https://doi.org/10.14778/3407790.3407804
- Kang et al. (2022) Daniel Kang, Edward Gan, Peter Bailis, Tatsunori Hashimoto, and Matei Zaharia. 2022. Approximate Selection with Guarantees using Proxies. https://doi.org/10.48550/arXiv.2004.00827 arXiv:2004.00827 [cs].
- Kang et al. ([n. d.]) Daniel Kang, John Guibas, Peter Bailis, Tatsunori Hashimoto, and Matei Zaharia. [n. d.]. Task-agnostic Indexes for Deep Learning-based Queries over Unstructured Data. ([n. d.]).
- Khattab et al. (2021) Omar Khattab, Christopher Potts, and Matei Zaharia. 2021. Relevance-guided Supervision for OpenQA with ColBERT. Transactions of the Association for Computational Linguistics 9 (2021), 929–944. https://doi.org/10.1162/tacl_a_00405 Place: Cambridge, MA Publisher: MIT Press.
- Khattab et al. (2023) Omar Khattab, Arnav Singhvi, Paridhi Maheshwari, Zhiyuan Zhang, Keshav Santhanam, Sri Vardhamanan, Saiful Haq, Ashutosh Sharma, Thomas T. Joshi, Hanna Moazam, Heather Miller, Matei Zaharia, and Christopher Potts. 2023. DSPy: Compiling Declarative Language Model Calls into Self-Improving Pipelines. https://arxiv.org/abs/2310.03714v1
- Khattab and Zaharia (2020) Omar Khattab and Matei Zaharia. 2020. ColBERT: Efficient and Effective Passage Search via Contextualized Late Interaction over BERT. https://arxiv.org/abs/2004.12832v2
- Kwon et al. (2023) Woosuk Kwon, Zhuohan Li, Siyuan Zhuang, Ying Sheng, Lianmin Zheng, Cody Hao Yu, Joseph E. Gonzalez, Hao Zhang, and Ion Stoica. 2023. Efficient Memory Management for Large Language Model Serving with PagedAttention. https://arxiv.org/abs/2309.06180v1
- Lee et al. (2019) Kenton Lee, Ming-Wei Chang, and Kristina Toutanova. 2019. Latent Retrieval for Weakly Supervised Open Domain Question Answering. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Anna Korhonen, David Traum, and Lluís Màrquez (Eds.). Association for Computational Linguistics, Florence, Italy, 6086–6096. https://doi.org/10.18653/v1/P19-1612
- Li et al. (2020) Mingjie Li, Ying Zhang, Yifang Sun, Wei Wang, Ivor W. Tsang, and Xuemin Lin. 2020. I/O Efficient Approximate Nearest Neighbour Search based on Learned Functions. 2020 IEEE 36th International Conference on Data Engineering (ICDE) (April 2020), 289–300. https://doi.org/10.1109/ICDE48307.2020.00032 Conference Name: 2020 IEEE 36th International Conference on Data Engineering (ICDE) ISBN: 9781728129037 Place: Dallas, TX, USA Publisher: IEEE.
- Liang et al. (2022) Percy Liang, Rishi Bommasani, Tony Lee, Dimitris Tsipras, Dilara Soylu, Michihiro Yasunaga, Yian Zhang, Deepak Narayanan, Yuhuai Wu, Ananya Kumar, Benjamin Newman, Binhang Yuan, Bobby Yan, Ce Zhang, Christian Cosgrove, Christopher D. Manning, Christopher Ré, Diana Acosta-Navas, Drew A. Hudson, Eric Zelikman, Esin Durmus, Faisal Ladhak, Frieda Rong, Hongyu Ren, Huaxiu Yao, Jue Wang, Keshav Santhanam, Laurel Orr, Lucia Zheng, Mert Yuksekgonul, Mirac Suzgun, Nathan Kim, Neel Guha, Niladri Chatterji, Omar Khattab, Peter Henderson, Qian Huang, Ryan Chi, Sang Michael Xie, Shibani Santurkar, Surya Ganguli, Tatsunori Hashimoto, Thomas Icard, Tianyi Zhang, Vishrav Chaudhary, William Wang, Xuechen Li, Yifan Mai, Yuhui Zhang, and Yuta Koreeda. 2022. Holistic Evaluation of Language Models. https://arxiv.org/abs/2211.09110v2
- Lin et al. (2024) Yiming Lin, Madelon Hulsebos, Ruiying Ma, Shreya Shankar, Sepanta Zeigham, Aditya G. Parameswaran, and Eugene Wu. 2024. Towards Accurate and Efficient Document Analytics with Large Language Models. http://arxiv.org/abs/2405.04674 arXiv:2405.04674 [cs].
- Liu et al. (2024b) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, and Gerardo Vitagliano. 2024b. A Declarative System for Optimizing AI Workloads. http://arxiv.org/abs/2405.14696 arXiv:2405.14696 [cs].
- Liu et al. (2023) Nelson F. Liu, Kevin Lin, John Hewitt, Ashwin Paranjape, Michele Bevilacqua, Fabio Petroni, and Percy Liang. 2023. Lost in the Middle: How Language Models Use Long Contexts. https://doi.org/10.48550/arXiv.2307.03172 arXiv:2307.03172 [cs].
- Liu et al. (2024a) Shu Liu, Asim Biswal, Audrey Cheng, Xiangxi Mo, Shiyi Cao, Joseph E. Gonzalez, Ion Stoica, and Matei Zaharia. 2024a. Optimizing LLM Queries in Relational Workloads. http://arxiv.org/abs/2403.05821 arXiv:2403.05821 [cs].
- Liu et al. (2024c) Shicheng Liu, Jialiang Xu, Wesley Tjangnaka, Sina J. Semnani, Chen Jie Yu, and Monica S. Lam. 2024c. SUQL: Conversational Search over Structured and Unstructured Data with Large Language Models. https://doi.org/10.48550/arXiv.2311.09818 arXiv:2311.09818 [cs].
- Liu et al. (2021) Wanqi Liu, Hanchen Wang, Ying Zhang, Wei Wang, Lu Qin, and Xuemin Lin. 2021. EI-LSH: An early-termination driven I/O efficient incremental c-approximate nearest neighbor search. The VLDB Journal 30, 2 (March 2021), 215–235. https://doi.org/10.1007/s00778-020-00635-4
- Lu and Kudo (2020) Kejing Lu and Mineichi Kudo. 2020. R2LSH: A Nearest Neighbor Search Scheme Based on Two-dimensional Projected Spaces. In 2020 IEEE 36th International Conference on Data Engineering (ICDE). 1045–1056. https://doi.org/10.1109/ICDE48307.2020.00095 ISSN: 2375-026X.
- Lu et al. (2018) Yao Lu, Aakanksha Chowdhery, Srikanth Kandula, and Surajit Chaudhuri. 2018. Accelerating Machine Learning Inference with Probabilistic Predicates. In Proceedings of the 2018 International Conference on Management of Data (SIGMOD ’18). Association for Computing Machinery, New York, NY, USA, 1493–1508. https://doi.org/10.1145/3183713.3183751
- Ma et al. (2023) Xueguang Ma, Xinyu Zhang, Ronak Pradeep, and Jimmy Lin. 2023. Zero-Shot Listwise Document Reranking with a Large Language Model. https://arxiv.org/abs/2305.02156v1
- Malkov and Yashunin (2018) Yu A. Malkov and D. A. Yashunin. 2018. Efficient and robust approximate nearest neighbor search using Hierarchical Navigable Small World graphs. http://arxiv.org/abs/1603.09320 arXiv:1603.09320 [cs].
- Park et al. (2015) Yongjoo Park, Michael Cafarella, and Barzan Mozafari. 2015. Neighbor-sensitive hashing. Proceedings of the VLDB Endowment 9, 3 (Nov. 2015), 144–155. https://doi.org/10.14778/2850583.2850589
- Patel et al. (2024) Liana Patel, Peter Kraft, Carlos Guestrin, and Matei Zaharia. 2024. ACORN: Performant and Predicate-Agnostic Search Over Vector Embeddings and Structured Data. Proceedings of the ACM on Management of Data 2, 3 (May 2024), 120:1–120:27. https://doi.org/10.1145/3654923
- Pradeep et al. (2023a) Ronak Pradeep, Sahel Sharifymoghaddam, and Jimmy Lin. 2023a. RankVicuna: Zero-Shot Listwise Document Reranking with Open-Source Large Language Models. http://arxiv.org/abs/2309.15088 arXiv:2309.15088 [cs].
- Pradeep et al. (2023b) Ronak Pradeep, Sahel Sharifymoghaddam, and Jimmy Lin. 2023b. RankZephyr: Effective and Robust Zero-Shot Listwise Reranking is a Breeze! http://arxiv.org/abs/2312.02724 arXiv:2312.02724 [cs].
- Qin et al. (2024) Zhen Qin, Rolf Jagerman, Kai Hui, Honglei Zhuang, Junru Wu, Le Yan, Jiaming Shen, Tianqi Liu, Jialu Liu, Donald Metzler, Xuanhui Wang, and Michael Bendersky. 2024. Large Language Models are Effective Text Rankers with Pairwise Ranking Prompting. https://doi.org/10.48550/arXiv.2306.17563 arXiv:2306.17563 [cs].
- Sachan et al. (2022) Devendra Sachan, Mike Lewis, Mandar Joshi, Armen Aghajanyan, Wen-tau Yih, Joelle Pineau, and Luke Zettlemoyer. 2022. Improving Passage Retrieval with Zero-Shot Question Generation. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 3781–3797. https://doi.org/10.18653/v1/2022.emnlp-main.249
- Shah and Wainwright (2016) Nihar B. Shah and Martin J. Wainwright. 2016. Simple, Robust and Optimal Ranking from Pairwise Comparisons. http://arxiv.org/abs/1512.08949 arXiv:1512.08949 [cs, math, stat].
- Sun et al. (2023) Weiwei Sun, Lingyong Yan, Xinyu Ma, Shuaiqiang Wang, Pengjie Ren, Zhumin Chen, Dawei Yin, and Zhaochun Ren. 2023. Is ChatGPT Good at Search? Investigating Large Language Models as Re-Ranking Agents. https://arxiv.org/abs/2304.09542v2
- Sundaram et al. (2013) Narayanan Sundaram, Aizana Turmukhametova, Nadathur Satish, Todd Mostak, Piotr Indyk, Samuel Madden, and Pradeep Dubey. 2013. Streaming similarity search over one billion tweets using parallel locality-sensitive hashing. Proceedings of the VLDB Endowment 6, 14 (Sept. 2013), 1930–1941. https://doi.org/10.14778/2556549.2556574
- Thakur et al. (2021) Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna Gurevych. 2021. BEIR: A Heterogenous Benchmark for Zero-shot Evaluation of Information Retrieval Models. https://doi.org/10.48550/arXiv.2104.08663 arXiv:2104.08663 [cs].
- Thorne et al. (2018) James Thorne, Andreas Vlachos, Christos Christodoulopoulos, and Arpit Mittal. 2018. FEVER: a Large-scale Dataset for Fact Extraction and VERification. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 809–819. https://doi.org/10.18653/v1/N18-1074
- Viola and Jones (2001) P. Viola and M. Jones. 2001. Rapid object detection using a boosted cascade of simple features. In Proceedings of the 2001 IEEE Computer Society Conference on Computer Vision and Pattern Recognition. CVPR 2001, Vol. 1. IEEE Comput. Soc, Kauai, HI, USA, I–511–I–518. https://doi.org/10.1109/CVPR.2001.990517
- Wang et al. (2021) Jianguo Wang, Xiaomeng Yi, Rentong Guo, Hai Jin, Peng Xu, Shengjun Li, Xiangyu Wang, Xiangzhou Guo, Chengming Li, Xiaohai Xu, Kun Yu, Yuxing Yuan, Yinghao Zou, Jiquan Long, Yudong Cai, Zhenxiang Li, Zhifeng Zhang, Yihua Mo, Jun Gu, Ruiyi Jiang, Yi Wei, and Charles Xie. 2021. Milvus: A Purpose-Built Vector Data Management System. In Proceedings of the 2021 International Conference on Management of Data (SIGMOD ’21). Association for Computing Machinery, New York, NY, USA, 2614–2627. https://doi.org/10.1145/3448016.3457550
- Wang et al. (2024) Liang Wang, Nan Yang, Xiaolong Huang, Binxing Jiao, Linjun Yang, Daxin Jiang, Rangan Majumder, and Furu Wei. 2024. Text Embeddings by Weakly-Supervised Contrastive Pre-training. http://arxiv.org/abs/2212.03533 arXiv:2212.03533 [cs].
- Wu et al. (2021) Jeff Wu, Long Ouyang, Daniel M. Ziegler, Nisan Stiennon, Ryan Lowe, Jan Leike, and Paul Christiano. 2021. Recursively Summarizing Books with Human Feedback. https://doi.org/10.48550/arXiv.2109.10862 arXiv:2109.10862 [cs].
- Wu et al. (2024) Shirley Wu, Shiyu Zhao, Michihiro Yasunaga, Kexin Huang, Kaidi Cao, Qian Huang, Vassilis N. Ioannidis, Karthik Subbian, James Zou, and Jure Leskovec. 2024. STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases. https://arxiv.org/abs/2404.13207v2
- Yue et al. (2024) Murong Yue, Jie Zhao, Min Zhang, Liang Du, and Ziyu Yao. 2024. Large Language Model Cascades with Mixture of Thoughts Representations for Cost-efficient Reasoning. http://arxiv.org/abs/2310.03094 arXiv:2310.03094 [cs].
- Yuksekgonul et al. (2024) Mert Yuksekgonul, Federico Bianchi, Joseph Boen, Sheng Liu, Zhi Huang, Carlos Guestrin, and James Zou. 2024. TextGrad: Automatic ”Differentiation” via Text. http://arxiv.org/abs/2406.07496 arXiv:2406.07496 [cs].
- Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah D. Goodman. 2022. STaR: Bootstrapping Reasoning With Reasoning. https://doi.org/10.48550/arXiv.2203.14465 arXiv:2203.14465 [cs].
- Zhang et al. (2024) Peiyuan Zhang, Guangtao Zeng, Tianduo Wang, and Wei Lu. 2024. TinyLlama: An Open-Source Small Language Model. arXiv:2401.02385 [cs.CL] https://arxiv.org/abs/2401.02385
- Zhang et al. (2022) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola. 2022. Automatic Chain of Thought Prompting in Large Language Models. https://doi.org/10.48550/arXiv.2210.03493 arXiv:2210.03493 [cs].
- Zheng et al. (2020) Bolong Zheng, Xi Zhao, Lianggui Weng, Nguyen Quoc Viet Hung, Hang Liu, and Christian S. Jensen. 2020. PM-LSH: A fast and accurate LSH framework for high-dimensional approximate NN search. Proceedings of the VLDB Endowment 13, 5 (Jan. 2020), 643–655. https://doi.org/10.14778/3377369.3377374

### v2

@@
-LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data
+Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data
 Abstract.
 The semantic capabilities of language models (LMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora.
-Unfortunately, existing systems lack high-level abstractions to perform semantic queries at scale.
-We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for semantic queries over datasets (e.g., sorting or aggregating records using natural language criteria).
+Unfortunately, existing systems lack high-level abstractions to perform bulk semantic queries across large corpora.
+We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for bulk semantic queries (e.g., filtering, sorting, joining or aggregating records using natural language criteria).
 Each operator can be implemented and optimized in multiple ways, opening a rich space for execution plans similar to relational operators.
-We implement our operators and several optimizations for them in LOTUS, an open source query engine with a Pandas-like API.
-We demonstrate LOTUS’ effectiveness across a series of real applications, including fact-checking, extreme multi-label classification, and search.
-We find that LOTUS’ programming model is highly expressive, capturing state-of-the-art query pipelines with low development overhead across these diverse applications.
-Specifically, on the FEVER dataset for fact-checking application, LOTUS’ programs can reproduce FacTool, recent state-of-the-art pipeline, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time.
-In the extreme multi-label classification task on the BioDEX dataset, LOTUS reproduces state-of-the art result quality with its join operator, while providing an efficient algorithm that runs faster than a naive join.
-In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods used by prior works.
+We implement our operators in LOTUS, an open source query engine with a DataFrame API.
+Furthermore, we develop several novel optimizations that take advantage of the declarative nature of semantic operators to accelerate semantic filtering, clustering and join operators by up to while offering statistical accuracy guarantees.
+We demonstrate LOTUS’ effectiveness on real AI applications including fact-checking, extreme multi-label classification, and search.
+We show that the semantic operator model is expressive, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality.
+Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while running up to 28 faster.
 LOTUS is publicly available at https://github.com/stanford-futuredata/lotus.
 1. Introduction
-The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI systems that reason over vast knowledge corpora.
+The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI-based analytics systems that reason over vast knowledge corpora.
 Many applications require complex reasoning over large amounts of data, including both unstructured and structured data.
@@
 Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023).
-Likewise organizations may wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business needs and productivity (dis, [n.
+Likewise, organizations wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business (dis, [n.
 d.]).
-Each of these tasks require a form of bulk semantic processing, where the AI system must process large amounts of data in often complex query patterns to perform the reasoning task at hand.
-Supporting the full generality of these applications with efficient and easily programmable query systems would have transformative impact.
-This prospect, however, raises two important and challenging questions: first, how should developers express semantic queries, and secondly, how should we design the underlying query system to achieve high efficiency and accuracy.
-Unfortunately, existing systems are insufficient for serving applications that require bulk semantic processing.
-Many existing LM programming frameworks (lan, [n. d.]; lla, [n. d.])
-and research works (Khattab et al., 2023, 2021; Khattab and Zaharia, 2020; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) provide methods, abstractions and optimizations for retrieval-augmented generation (RAG), which first performs a semantic search over the text corpus, then invokes the LM conditioned on the user question and retrieved documents.
-While RAG is one common and useful query pattern, it is limited to point lookups into the data corpus and assumes the user query can be answered by one or a small set of retrieved documents.
-However, bulk semantic processing pipelines may involve more complex patterns not captured by RAG, such as semantic aggregations or transformations over many documents.
-Alternatively, several systems (ver, [n.
+Each of these tasks require a form of bulk semantic processing, where the analytics system must process large amounts of data and apply semantic-based analysis across a whole dataset.
+Supporting the full generality of these applications with efficient and easy-to-use analytics systems would have a transformative impact, similar to what RDBMSes had for tabular data.
+This prospect, however, raises two challenging questions: first, how should developers express semantic queries, and secondly, how should we design the underlying analytics system to achieve high efficiency and accuracy.
+Unfortunately, existing systems are insufficient for applications that require bulk semantic processing.
+Several systems provide logically row-wise LM operators, either in the form of a general-prupose LLM user-defined function (UDF) (ver, [n.
 d.]; dat, [n.
-d.]; noa, 2023a; Liu et al., 2024b) offer and optimize LLM user-defined functions (UDFs) in SQL.
-This model offers a logically row-wise LM-execution model for batch processing data with arbitrary prompts in composition with arbitrary SQL queries.
-While useful for batch-processing applications, this model cannot support LM-reasoning patterns across rows and provides a low-level of abstraction with a simple LM() function.
+d.]; noa, 2023; Liu et al., 2024b) or specialized row-wise operators (Liu et al., 2024c, a; Lin et al., 2024) for specific applications, such as conversational chat or data cleaning and extract-transform-load (ETL) tasks.
+These programming models fail to support a range of LM-based transformations across rows, such as ranking, grouping or joining records.
+Alternatively, many LM programming frameworks (lan, [n.
+d.]; lla, [n.
+d.]; Khattab et al., 2023) and research works (Khattab et al., 2021; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) support combining LLM calls with retrieval-augmented generation (RAG) and Text2SQL, however these models are limited to point lookups to the data.
+RAG and Text2SQL methods first retrieve data using semantic search or LM-generated SQL calls, and then process the results with an LLM.
+Both are limited to analyses where an LLM can be fed a small number of documents or rows that are placed into a single model call’s context window (e.g., questions answered in one document in the corpus).
+However, bulk semantic processing applications may involve analysis across multiple rows, for clustering, ranking or joining an entire dataset.
 Towards a declarative programming interface for bulk semantic processing, we propose semantic operators, which extend the relational model with AI-based operations that users can compose into powerful, reasoning-based query pipelines over structured and unstructured data.
 These operators include semantic filters, joins, rankings, aggregations, and projections, which take natural language expressions given by the programmer.
-To provide an implementation of these operators, we present the LOTUS (LLMs Over Tables of Unstructured and Structured data) system and programming model.
-LOTUS’ query engine efficiently executes queries with semantic operators using a variety of algorithms and optimizations for each operator, while abstracting away low-level details like model context length limits and choice of algorithms.
+To provide an implementation of these operators, we present the LOTUS system, which exposes these operators through a simple DataFrame-based programming interface.
+LOTUS’ query engine efficiently and accurately executes queries with semantic operators, while abstracting away low-level details like model context length limits and algorithmic choices.
+Finally, we propose novel optimizations for the semantic filter, join, ranking and groupby operators that can improve performance by up to over a naive, expensive implementation while guaranteeing accuracy similar to the naive implementation.
 | Operator | Description |
@@
 | Joins a table against a second table by keeping all pairs of tuples that pass the provided langex predicate. | |
-| Performs a similarity join where each row of the source table is joined with the most semantically similar tuples from the table , using fields and as the left and right join keys, respectively. | |
 | Performs an aggregation over the input tuples according to the langex, which specifies a commutative, associative aggregation function over a list of tuples. | |
 | Ranks each tuple and returns the best according to the langex, which specifies a ranking function that sorts a list of tuples. | |
+| Groups the tuples into categories based on the langex, which specifies a grouping criteria. | |
 | Performs a projection, returning a new column, according to the provided langex. | |
-| Performs a projection according to the langex, returning a column with a list of substrings from the input tuples. | |
-| Performs a semantic similarity clustering over column to create groups. | |
-| Performs a top- search over column using query . | |
-| Creates similarity index for column and saves to path. | |
 Figure 1 begins to demonstrate the power of LOTUS’ declarative programming model and optimized query engine.
-For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we can easily create 3 distinct query pipelines, each written in intuitive LOTUS programs of less than 50 lines of code by composing semantic operators (e.g.
-filters, maps, search and joins).
-The modularity and composability of these semantic operators allows us to quickly explore the design space on this task.
-In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by up to compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), while also providing query efficiency, by default, to maintain modest execution time.
-LOTUS’ optimizer also exploits the rich implementation design space of semantic operators to leverage new algorithmic and optimization opportunities.
-For expensive operators like semantic joins, aggregations, ranking and filters, LOTUS implements novel algorithms that maximize parallel batched-inference opportunities, use model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) with a lightweight scoring function unique to our setting, leverage semantic similarity indices, and perform algorithmic approximations.
+For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we write an intuitive LOTUS program in less than 50 lines of code by composing semantic operators (i.e., filters, maps, and joins).
+The modularity and composability of these semantic operators make it easy to quickly write programs and explore the design space on this task.
+In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by percentage points compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), which was written in over 750 lines of code.
+LOTUS also provides query efficiency by default, yielding lower execution time than the FacTool implementation.
+LOTUS’ optimizer exploits the rich implementation design space of semantic operators to provide novel optimizations for bulk semantic operations.
+To demonstrate the power of programming with semantic operators, we implement optimizations for several costly operators, i.e., semantic filter, join, ranking and group-by.
+Specifically, we develop a novel, lossless optimization for ranking.
+Additionally, for filters, joins and group-bys, we define and implement both a tractable, but expensive, reference algorithm as well as an approximation algorithms with probabilistic guarantees on accuracy with respect to the reference algorithm.
+Our approximation algorithms leverage build on statistical techniques used in prior works (Kang et al., 2021, 2020) with novel and efficient proxy scorers that leverage small LMs or semantic embeddings.
 Figure 1 demonstrates the effectiveness of these methods.
-Compared to the best-performing un-optimized LOTUS program, Program B shown in the figure, the logically equivalent program implemented using LOTUS’ optimizations attains higher accuracy and runs faster.
-We systematically evaluate LOTUS through a series of real applications, including fact-checking, extreme multi-label classification, and search.
-Our results show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications.
-Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time.
-In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, demonstrating the power of LOTUS’ declarative interface.
+The program implemented using LOTUS’ optimizations with recall and precision targets and a error probability attains high accuracy, of the unoptimized program’s accuracy, and runs faster than the unoptimized program.
+We systematically evaluate the expressiveness of the semantic operator model and the efficiency of our algorithms and optimizations with statistical accuracy guarantees through several real applications, including fact-checking, extreme multi-label classification, and search, which had previously been implemented by hand in AI research papers.
+We show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications.
+Furthermore, our algorithms and optimizations can speed up query execution by up to while offering statistical guarantees on accuracy.
+Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering or lower execution time with or without batching, respectively, while providing statistical accuracy guarantees.
+In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, while again offering statistical guarantees on accuracy.
 In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods (Qin et al., 2024) used by prior works.
-2. The LOTUS Programming Model
-We now introduce the LOTUS programming model and show its expressive power in allowing developers to declaratively specify AI-based query pipelines that bulk process large datasets of structured and unstructured data.
-LOTUS extends the relational model with semantic operators, which we show in Table 1.
-These operators can be easily composed together with standard relational operators to build powerful programs that are transparently optimized (Section 3).
-In this section, we describe each semantic operator and our API for them in LOTUS, grounding each in concrete examples.
-While our current API implementation extends Pandas (pan, [n.
-d.]), LOTUS’ semantic operators could be used with other existing relational query languages and APIs, such as SQL.
+Our main contributions are the following:
+We propose semantic operators as a programming model for AI-based analytics.
+Semantic operators (e.g., filters, joins, top-k ranking, aggregations, and group-bys) are declarative and can be implemented in multiple ways, creating a rich design space for execution plans.
+We present efficient algorithms and query optimizations for several costly operators, i.e., semantic filter, join, top-k ranking and group-by.
+We develop a novel, lossless optimization for ranking and approximations for filters, joins and group-bys that provide statistical accuracy guarantees with respect to tractable, but expensive, reference algorithms.
+We provide a reference implementation of semantic operators and optimizations in the open-source LOTUS system.
+2. Semantic Operators
+We define semantic operators as declarative transformations on one or more relations, parameterized by a natural language expression, that can be implemented by a variety of AI-based algorithms.
+Semantic operators provide a seamless extension to the relational model, operating over tables that may contain traditional structured data as well as unstructured fields such as free-form text.
+These new operators can be composed with traditional relational operators to build rich queries that can be transparently optimized, which we discuss in Section 4.
+Table 1 lists a core set of semantic operators we propose in this paper, which cover the common semantic transformations we observe in real-world applications and mirror the key transformation patterns in standard relational operators.
+Specific analytics systems may of course provide additional semantic operators or APIs that combine the ones we discuss here.
+Each semantic operator takes a parameterized natural language expressions (langex for short), which are natural language expressions that specify a function over one or more attributes for the given semantic transformation.
+The langex signature depends on its semantic operator and may represent a predicate, aggregation, comparator function, or projection in natural language.
+Figure 2 shows an example of using semantic operators in LOTUS’ dataframe API to load data about ArXiv papers and perform a summarization task, which finds relevant papers using a top-k search, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest with an aggregation.
+The code in Figure 2 highlights various langex signatures: the sem_filter langex provides a predicate that indicates a filter criteria to apply over the title and abstract attributes, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over abstracts.
+Together, each langex and semantic operator specifies a modular semantic transformation, whose behavior is defined with respect to a world model , which may be a powerful LLM that captures a probability distribution over the vocabulary .
+We briefly overview several semantic operators, both ones whose definitions directly extend their analogs in traditional relational algebra, and those with more complex definitions.
+Semantic Filter is a unary operator over the relation and returns the relation , where the natural language predicate is specified by the langex with respect to model .
+The langex is paramterized by one or more attributes from the relation that are relevant for evaluating the predicate.
+Semantic Join provides a binary operator over relations and to return the relation , once again taking the langex with respect to model .
+Here the langex is parameterized by the left and right join keys and describes a natural language predicate over both.
+Semantic Top-k imposes a ranking111This definition implies that imposes a total and consistent ordering.
+However, this definition can also be softened to assume partial orderings and noisy comparisons with respect to model .
+over the relation and returns .
+Here the signature of the langex provides a general ranking criteria according to one or more attributes of the input relation.
+The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation.
+Semantic Aggregation performs an aggregation over the input relation, returning .
+Here, the the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results.
+We note that the langex with respect to model assumes infinite context and can take an arbitrary number of input tuples.
+Managing finite context limits of the underlying model is an implementation detail left to the system.
+Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents.
+Semantic Group-by takes a langex that specifies a grouping criteria and a target number of groups.
+As a example extending that of Figure 2, we might group-by the topics presented in each paper abstract with a target of 10 group.
+Here, the group-by operator must discover representative group labels and assign each tuple a label.
+In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n.
+d.]).
+Clustering algorithms over points in a metric space typically optimize the potential function tractably using coordinate descent algorithms, such as k-means.
+In the semantic group-by, the clustering task is over unstructured documents with a natural language similarity function specified by the user langex , which imposes a real-valued score between a tuple and a candidate label , and the operator poses the following optimization problem:
+where are group labels, each consisting of tokens in the vocabulary, .
+3. The LOTUS Programming Interface
+In this section we describe the LOTUS API, which implements the semantic operator model as an extension of Pandas (pan, [n.
+d.]).
+We chose a Pandas-like API in our initial system implementation to make it easy for users to integrate LOTUS with popular AI libraries in Python.
+However, semantic operators could also be added to a variety of other data processing APIs and query languages, such as SQL.
 3.1. Datatypes
-LOTUS’ data model consists of tables with structured and unstructured text fields, and our current implementation extends Pandas (pan, [n.
+3.1.1. Data Model
+LOTUS’ data model consists of tables with structured and unstructured text fields.
+LOTUS’ semantic operators can take both of these data-types as inputs.
+For instance, Figure 2 shows an example LOTUS program, where the table schema over ArXiv papers consists of attributes for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date.
+These columns are then passed as parameters to semantic operators, such as sem_search, sem_filter, sem_agg.
+Additionally, LOTUS supports semantic indices over natural-language text columns to provide optimized query processing.
+These indices leverage semantic embeddings over each document in the column to capture semantic similarity using embedding distance metrics.
+Semantic indices can be created off-line with sem_index and a specified retriever model, and then loaded using load_sem_index.
+3.1.2. Parameterized Natural Language Expressions (langex)
+Programmers write each langex in natural language, parameterized by one or more data columns, which are indicated in the brackets within a string, or double brackets within a formatted string, as Figure 2 shows.
+Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user.
+Operator-specific LLM prompts based on these expressions are automatically constructed by LOTUS’ underlying query engine, and the underlying system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
+3.2. Semantic Operators in LOTUS
+We now overview the semantic operators supported in the LOTUS API, which includes the core set of operators described in Section 2, as well as several additional variants provided for convenience.
+Sem_filter, Sem_join, & Sem_sim_join.
+The LOTUS API supports sem_filter and sem_join, both of which take a langex predicate, as described in Section 2.
+In addition, LOTUS provides a join variant, sem_sim_join, where tuples are matched according to their semantic similarity, rather than an arbitrary natural-language predicate.
+Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join, which indicates additional optimization opportunities to the query engine.
+Figure 3 provides an example of the sem_join compared to the sem_sim_join, where the user specifies the left and right table join keys, and a parameter .
+The operator performs a left join such that for each row in the left table, the output table will contain matching rows from the right table with the highest semantically similarity scores.
+Sem_topk & Sem_search .
+LOTUS supports a semantic top-k, which takes the langex ranking criteria, as described in Section 2.
+Programmers can optionally specify a group-by parameter to indicate a subset of columns to group over during ranking, as shown in Figure 4.
+The groupings are defined using standard equality matches over the group-by columns.
+Additionally, as the Figure shows, LOTUS also provides a top-k variant, sem_search, which assumes a semantic similarity-based ranking criteria relative to a natural language query.
+LOTUS also exposes advanced relevance-based re-ranking functionality for search, allowing users to specify the n_rerank parameter during the semantic search.
+The semantic search in this case will first find the top- most relevant documents and then re-rank the top- found documents to return the top n_rerank.
+Sem_agg performs an aggregation over the input relation, with as langex signature that provides a commutative and associative aggregation function, as shown by Figure 2.
+Similar to sem_topk, LOTUS allows users to specify a group-by parameter.
+Sem_group_by creates groups over the input dataframe according to the langex grouping criteria by clustering the data into a target number of groups (e.g., papers_df.sem_group_by("the topic of the abstract", 20)).
+By default, the operator automatically selects group labels, but the user can also optionally specify target labels.
+This operator is useful both for unsupervised discovery of semantic groups, and for semantic classification tasks.
+Sem_map & Sem_extract both perform a natural language projection over an existing column.
+While sem_map projects to an arbitrary text attribute, sem_extract projects each tuple to a list of sub-strings from the source text.
+This is useful for applications, such as entity extraction or fact-checking, where finding snippets or verified quotes may be preferable to synthesized answers.
+4. Execution Plans and Optimizations for Semantic Operators
+Semantic operators create a rich design space of diverse execution plans which have significant consequences on system efficiency and accuracy.
+For example, there are many possible algorithms to join two collections of records using a LLM and a natural language expression (all the standard join algorithms, as well as ML heuristics to approximate the join, e.g., by semantically clustering similar records using an embedding model).
+Specifically our goal is to offer an efficient plan while while guaranteeing high accuracy.
+Defining accuracy for semantic transformations, however, presents a fundamental challenge in the absence of ground truth labels.
+Therefore, we define accuracy of each operator plan with respect to a ”gold” algorithm, which uses a powerful oracle model .
+These ”gold” algorithms are tractable solutions that can achieve high quality results, but may be expensive in terms of execution time and LM call complexity.
+We discuss the design space of these ”gold” algorithms for several core semantic operators we introduced.
+In addition, we present several novel optimizations for semantic filters, joins, top-k and group-by.
+These optimizations are either lossless or provide statistical guarantees on accuracy with respect to a ”gold” algorithm with the expensive model .
+Specifically, our approximations emulate the results of the expensive but high-quality algorithm by using model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) or semantic indexes, building on past work on accelerating ML-based query processing with guarantees (Kang et al., 2022, 2019) but achieving this property for new types of operators.
+Notably, these optimizations can all be performed declaratively given a user’s accuracy targets.
+We show in Section 5 that they significantly accelerate real applications.
+While this paper proposes new optimizations for only a few, expensive semantic operators, we envision that a wide range of other new and traditional optimizations are possible under the semantic operator model.
+For example, several prior works demonstrate performance gains in both logical query plan optimizations (e.g.
+operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g.
+code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)).
+4.1. sem_filter
+4.1.1. Gold Algorithm
+Our gold algorithm for the semantic filter runs batched LLM calls over the tuples in relation .
+Each tuple, , is processed independently using the model , which is prompted to output a boolean value indicating whether the tuple passes the langex predicate.
+4.1.2. Algorithmic Approximation
+We can declaratively optimize semantic filters by providing an approximation which will obtain recall and precision targets and with respect to the gold implementation with probability .
+To achieve reduced cost we leverage the cheaper, but less accurate proxy model , which provides a score indicating whether the tuple passes the predicate, in conjunctions with the oracle LLM .
+This idea is inspired by prior works (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022; Yue et al., 2024), which leverage model cascades for different problem settings.
+Specifically, several works study cascades in the video analytics setting with vision models, which have significantly different properties than LLMs, requiring different proxy scoring mechanisms.
+Additionally, other works, which study cascades for LLMs, use heavy-weight scoring mechanisms to decide whether to use the proxy model, which incurs a heavy execution time penalty, and does not provide accuracy guarantees.
+In contrast, we focus on providing accuracy guarantees for the constrained problem of applying cascades for filters, which allows us to use lighter-weight scoring functions than the more general case studied in prior works.
+Providing accuracy guarantees is important because, in general, we cannot assume the proxy model is accurate.
+In fact, the proxy model, may perform very poorly for some tasks and datasets.
+The goal of our algorithm is to automatically discover the quality of the proxy by sampling and comparing to the gold algorithm with the oracle model.
+If the proxy provides high quality results, our algorithm will exploit it, leading to lower execution time.
+However, if the proxy model performs poorly, we will selectively route to it according to the proxy scores , and a learned threshold on the proxy scores.
+As prior work (Kang et al., 2020) discusses, such sampling-based algorithms to learn a threshold, , introduce multiple-hypothesis testing problems, requiring statistical corrections using confidence intervals, which we carefully apply in our algorithm.
+Our algorithm leverages a small LLM as the proxy model.
+We generate scores using the the log-probabilities corresponding to the True or False output tokens of the proxy model’s answer and re-scaling by the quantiles over the proxy scores of tuples in the relation.
+Algorithm 1 shows the full procedure for performing the approximate semantic filter.
+We begin by using an collecting a sample , which we label with the oracle and proxy models.
+The sampling procedure uses importance sampling and defensively mixes a uniform sample following prior work (Kang et al., 2022).
+Using a the central limit theorem and the normal approximation on the distribution of sample statistics, we then find thresholds and , which will respectively ensure the desired precision target and recall target are met, each with an error probability of .
+These sub-procedures follow the statistical techniques used in prior work (Kang et al., 2022), but applies them to this new setting where we must ensure both targets are met, requiring a correction for hypothesis testing and multiple failure modes.
+Finally the algorithm proceeds to process each tuple in the relation.
+If the tuple’s proxy score is at least , we mark it as passing the predicate.
+If the tuple’s proxy score is less than or equal to , we mark it as failing the predicate, and for tuples with proxy scores between and , we resort to the oracle model to provide a label.
+4.2. sem_join
+4.2.1. Gold Algorithm
+Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left table, and right table, .
+The gold algorithm implements a nested-loop join pattern with efficient batched inference to maximize GPU utilization.
+This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively.
+Each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate.
+This quadratic join algorithm is suitable for small join tables but scales poorly for large tables.
+4.2.2. Algorithmic Approximation
+Similar to semantic filters, we declaratively optimize semantic joins by providing an approximation that obtains recall and precision targets and with respect to the gold algorithm with probability .
+Due to the expensive quadratic scaling of the nested-loop join, rather than using a small LLM as the proxy model, we leverage an even cheaper proxy based on semantic similarity scores, which can be obtained efficiently with a semantic index over the join keys.
+We leverage two possible join approximation algorithms and dynamically chooses the lowest cost one, which will minimize the number of LLM calls.
+The first approximation, sim-filter produces a proxy score , based on embedding similarity.
+We leverage the semantic index to perform batched similarity search and re-calibrate similarity scores according to the quantiles of the distribution over all pairs to obtain proxy scores.
+This approximation is likely to perform efficiently when tuple pairs with high semantic similarity between the the right and left join key are more likely to pass the predicate.
+This correlation phenomenon between predicate matches and embedding similarity has been studied by prior works (Patel et al., 2024) and is not always present.
+Thus, we introduce an alternative approximation suitable when correlation is not present.
+The second approximation, map-sim-filter first performs a projection over the left join key and then uses the projected column to compute proxy scores uses the embedding similarity between the projected value and .
+As before we also re-calibrate the proxy scores taking their quantiles.
+Intuitively, performing the projection is useful when tuple pairs that pass the user’s natural language predicate exhibit low semantic similarity.
+The projection step invokes the LM over each tuple in left join table, prompting it to predict values in the attribute domain of the right join key that match.
+Notably, this LM projection is ungrounded, i.e., it is done without knowledge of the right join key’s attribute domain and can thus be performed in a fully parallelized semantic map operation, which can optionally take user-provided demonstrations.
+Figure 5 provides an example of a semantic join between a table of papers and datasets, where the predicate evaluates whether a given paper abstract uses a specific dataset.
+Here, the map step would invoke the LM over each abstract, instructing the model to output the dataset used, conditioned only on the abstract.
+To dynamically choose between these two approximations, we follow a procedure similar to Algorithm 1 used for semantic filters.
+We begin by importance sampling to collect the sample and construct the set of oracle labels, , over the sample.
+We then obtain , and independently for each approximation algorithm using their respective proxy scores.
+Using the tuned thresholds and known proxy scores over for both algorithms over the dataset, we then determine the exact oracle cost needed to execute either algorithm, and we take the least cost plan with its associated learned thresholds to proceed to evaluate the predicate for each tuple pair.
+4.3. sem_group_by
+Performing this semantic operator entails discovering group labels and classifying each input tuple using the discovered labels.
+In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n.
 d.]).
-Figure 2 shows an example LOTUS program that loads data about ArXiv papers and performs a summarization task, which finds relevant papers, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest.
-We briefly describe two core components of LOTUS programming model: its data model (Section 2.1.1), and its parameterized natural language expressions (Section 2.1.2) for specifying semantic operations over the data.
-2.1.1. Data Model
-LOTUS is designed to seamlessly extend the relational model.
-Each row in the table, represents a logical entity.
-Table columns may contain either structured data or unstructured fields with free-form natural-language text.
-LOTUS’ semantic-relational operators can take both of these data-types as inputs.
-Figure 3 shows an example table, where each rows represents an ArXiv paper document, with fields for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date.
-These columns are then passed as parameters to semantic-relational operators, such as sem_search, sem_filter, sem_agg, as shown in Figure 2.
-Additionally, LOTUS supports semantic similarity indices over natural-language text columns to provide optimized semantic query processing.
-These indices create embeddings over each document in the column to capture semantic similarity using embedding distance metrics.
-Semantic indexes can be created off-line and then loaded using sem_index and load_sem_index.
-Figure 2 provides an example program, where upon reading the ArXiv papers data from a CSV file, the programmer loads a semantic index over the abstract column.
-The program then repeatedly uses the semantically-indexed column in subsequent LOTUS operations, involving semantic search, filtering and aggregation.
-2.1.2. Parameterized Natural Language Expressions (langex)
-A core principle of LOTUS is to provide users with a declarative interface that separates the user-specified, logical query plan from its underlying implementation.
-As such, users program with LOTUS’ semantic operators by writing parameterized natural language expressions (langex111Akin to regular expressions (regex), which specify character patterns to match in text, langex are natural language expressions to programmatically specify reasoning-based patterns over structured data and free-form text.),
-rather than directly prompting an underlying LM.
-Figure 2 shows an example of this, where the programmer provides a langex as the parameter to the sem_filter (line 18) and sem_agg (line 19) operations.
-Programmers write each langex in natural language text, parameterized by one or more data columns, which are indicated in the double brackets within the formatted string.
-The function of these expressions varies according to the semantic operator used and may represent a predicate, aggregation, comparator function, or projection in natural language.
-For instance, as shown in Figure 2, the langex signature of sem_filter provides a predicate that indicates a filter criteria to apply over paper titles and abstracts, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over paper abstracts.
-Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user.
-This makes it simple for the programmer to specify diverse and complex, multi-step query pipelines.
-All operator-specific prompts are automatically handled by LOTUS’ underlying query engine, and the system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
-2.2. Semantic Operators
-We now overview each semantic operator and their corresponding LOTUS API. Table 1 provides a concise summary of each operator.
-Sem_filter returns the subset of rows that pass the filtering condition, specified by the user’s langex.
-As Figure 4 shows, the langex signature provides a semantic predicate over one or more table columns and can be answered by a binary ”True” or ”False” answer.
-Sem_topk ranks a set of rows according to the user-defined criteria, and returns the rows that best match the ranking criteria.
-The signature of the langex provides a general ranking criteria according to one or more columns.
-The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation.
-As Figure 5 shows, the programmer uses the langex to specify arbitrary reasoning-based ranking criteria, such as ranking paper abstracts by the most outrageous claim made.
-Programmers can also optionally specify a group-by parameter to indicate a subset of columns to group over during ranking, as shown in 6.
-The groupings are defined using standard equality matches over the group-by columns.
-To use groupings according to semantic similarity, users can also perform a sem_cluster_by and pass the resulting cluster_id column as the group-by column parameter.
-Sem_join combines data from two tables, evaluating the user’s predicate to return the set of rows from the left and right table that pass.
-As Figure 7 shows, users specify the right join table, a langex, and optionally, the join key of the left and right tables.
-Here the langex contains two or more columns, and provides a predicate over the left and right tables.
-By default, the operator performs an inner join over the two tables, and can alternatively perform left, right or outer joins, if specified.
-Sem_sim_join provides a variant of the semantic join, such that rows are matched according to their semantic similarity, rather than an arbitrary natural-language predicate.
-Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join.
-Figure 8 provides an example, where one table contains papers with an indexed column of abstracts, and the other table contains a list of research interests.
-The user specifies the left and right table join keys, and a parameter .
-The left join key may or may not be indexed, whereas the the right join key must be a semantically indexed column with its index loaded.
-The operator performs a left join such that for each row in the left table, the output table will contain matching rows from the right table with the highest similarity scores.
-The programmer can also optionally specify a return column containing the semantic similarity scores of each joined row.
-Sem_agg performs an aggregation over all rows of the table.
-As Figure 9 shows, the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results.
-Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents.
-Similar to sem_topk, users can also specify a group-by parameter to use.
-We also provide additional flexibility, by allowing the programmer to compose semantic aggregations with sem_partition_by, which provides finer-granularity control over how documents are grouped over in each LM invocation.
-We describe this further below, and show its use in Section 3 for overriding the default commutativity and associativity assumptions for some tasks.
-Sem_partition_by creates a row partitioning over the table, which will be used by sucessive calls to sem_agg to decide which rows to group together within each LM invocation.
-We find in Section 3 that this can non-trivially affect the result quality of aggregation tasks, like summarization.
-As Figure 10 shows, sem_partition_by takes a function, which outputs a group-id for each row.
-LOTUS natively supports a semantic cluster function, which takes the number of clusters to create and a column, but users can also specify arbitrary partition functions.
-The group-ids output by the partition function indicate to the system which rows should be grouped together, in a best-effort manner, during LM invocations.
-The system will aggregate over documents within each group, before merging intermediate results across groups.
-Sem_map performs a natural language projection over an existing column and outputs a new column in the table.
-As shown in Figure 11, the user’s langex specifies what to project.
-This operator provides general functionality and simulates logically row-wise data-flow similar to prior works that use LLM-UDFs in relational languages (Liu et al., 2024a, c).
-These semantic projections are broadly useful for a variety of tasks, such as row-wise summarization, classification, and entity-extraction.
-Sem_extract provides similar functionality to sem_map but provides the answers by returning a list of sub-strings from the source text.
-This functionality is useful for applications, such as entity extraction or fact-checking, where snippet-finding or verified quotes may be preferable to synthesized LLM-based answers.
-As Figure 12, the user’s langex signature specifies a projection, similar to sem_map.
-Sem_index generates a semantic similarity index over the specified data column, as shown in Figure 13.
-To generate the semantic index, users first use the settings module to declare a retrieval model, which will be used for generating semantic embeddings and indexing the column data.
-The sem_index operator takes a column and a local directory path, to which the generated semantic index will be stored.
-As the figure shows, users can separately index multiple columns of the table.
-Load_sem_index re-loads the stored semantic index upon reading the table data from disk, as shown in Figure 14.
-The user specifies the column corresponding to the semantic similarity index and the directory where the index was saved.
-Sem_search performs a top- semantic similarity search over a semantically-indexed column, as Figure 14 shows.
-The user specifies the table column to search over, a query in natural language, a target of the number of results to return, and optionally indicates whether to return similarity scores as a column in the returned table.
-LOTUS also exposes advanced relevance-based re-ranking functionality for search.
-Users can specify a re-ranker model, as shown in Figure 15, and then set the n_rerank parameter during the semantic search.
-The semantic search in this case will first find the top- most relevant documents according to the retriever model, and then re-rank the top- found documents to return the top n_rerank ones using the re-ranker model.
-Sem_cluster_by assigns a group to each row of the table according to semantic similarity.
-This operator is akin to a relational group_by, but uses semantic similarity to determine the groups, rather than equality over the specified column.
-As Figure 16 shows, the user specifies an indexed column to cluster, and the number of clusters to create.
-The returned table augments the input table with an additional column that specifies the assigned cluster_id of each row.
-Users can optionally also return the similarity scores between each row and its cluster centroid in an additional column of the returned table.
-3. Implementing and Optimizing Semantic Operators
-Semantic operators create a rich design space, allowing for a diverse set of algorithmic decisions and optimizations which have significant consequences on system efficiency and accuracy.
-We identify multiple possible algorithms for these operators, noting their performance implications.
-LOTUS’ query engine aims to automatically exploit this rich design space to provide an optimized implementation of each individual operator and composite pipelines.
-Specifically, LOTUS employs a series of novel algorithms designed to leverage the semantics of each operator in order to maximize parallel batched-inference opportunities, use model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) with a lightweight scoring function unique to our setting, leverage the semantic similarity index, and perform algorithmic approximations for expensive operations (e.g.
-joins).
-We see in Section 4 that these decisions can significantly improve efficiency while maintaining high quality results for LOTUS’ query pipelines.
-While our initial investigation studies several operator-specific design decisions unique to the semantic bulk processing setting, we envision a rich set of additional optimization opportunities.
-Several prior works demonstrate performance gains in both logical query plan optimizations (e.g.
-operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g.
-code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)), which we find to be promising opportunities for LOTUS’ implementation, left to future work.
-LOTUS’ current implementation extends the Pandas API (pan, [n.
-d.]).
-We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) for efficient vector similarity search.
-3.1. sem_filter
-LOTUS’ semantic filter runs batched LLM calls over a set of rows, prompting the model to output a boolean value for each row to indicate whether the record passes the user’s natural language predicate.
-Figure 18 shows an example of the operator-specific filtering instruction, which is managed transparently by the system.
-3.1.1. Optimizations
-We leverage model cascades, using an efficient scoring function unique to our setting, to further optimize the filter function.
-Many prior works have studied model cascades for ML tasks (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022), leveraging a relatively cheap and inaccurate proxy model, along with an accurate but expensive oracle model in order to reduce the inference cost by routing easy queries to the small model, and resorting to the large model where necessary.
-Typically, a scoring function provides confidence scores over the proxy model’s outputs, and allows the system to decide when to route examples to the larger model.
-Applying this paradigm to LMs introduces several challenges, including generating a reliable scoring function.
-Several works study this problem, suggesting to fine-tune a smaller LM to score each question along with the answer produced by the weak LLM (Chen et al., 2023), or invoking multiple sampling paths from the small LM and evaluating self-consistency to build a confidence score (Yue et al., 2024).
-While these methods provide general-purpose scoring functions, LOTUS’ scoring function leverages the operator semantics to provide a lighter-weight scoring function, which avoids the execution time penalty of scoring with an additional model or re-sampling from the proxy-model.
-Specifically, LOTUS’ cascade scoring function leverages the binary-label output of each LM invocation.
-The system first batch-processes each record using the cheaper LM, then computes confidence scores by exponentiation the log-probabilities for the output LLM tokens corresponding to the True or False answer.
-For records with confidence scores below the user-defined threshold, the system then passes these records in batch to the larger LM.
-As we show, in Section 4, leveraging model cascades in this way for semantic filtering can significantly improves the system’s efficiency while maintaining high quality results for some tasks.
+Our gold algorithm instead offers a tractable implementation to obtain group labels using a linear pass over the data with the reference model .
+Furthermore, once the labels are discovered, we perform the LM-based classification task and show how to optimize this step using an approximation with probabilistic guarantees.
+4.3.1. Gold Algorithm
+Our semantic group-by algorithm proceeds in two stages.
+The first stage performs the unsupervised label discovery, given the user’s grouping criteria and target number of groups.
+This stage first performs a semantic projection, prompting the LM to predict a label for each input tuple.
+Then, we create a semantic index over the projected column and perform a vector clustering over these predicted labels, using FAISS’ efficient k-means implementation to construct groups.
+Lastly, for each group, we top-k sample documents based on their similarity to the centroid, and few-shot prompt the LM to generate a label describing the given documents according to the user-provided grouping criteria (e.g.
+the topics of the documents) using a semantic aggregation.
+In the second stage of the algorithm, we use the generated labels to perform classification over each document.
+We do this using a semantic map, which prompts the LM to output a label conditioned on the candidate labels and document.
+4.3.2. Algorithmic Approximation
+We provide an efficient approximation, specifically for the classification task in the second stage of the algorithm.
+Similar to semantic filters and join, we perform this optimization declaratively, while guaranteeing a classification accuracy target is met with probability .
+Similar, to semantic joins, we leverage semantic similarity scores as a cheap proxy.
+Specifically, we leverage the index constructed in stage one of the algorithm and compute similarity scores between the embeddings of each document’s un-grounded prediction and the embeddings of each class label.
+The proxy score of the tuple with a predicted label, , for group label is given by .
+Our goal is then to learn a threshold, , such that if the similarity between an embedding and the class label’s embedding is greater than , we return the class label, and otherwise we resort to the more expensive LM-based classification procedure.
+We accomplish this by uniform sampling and running the sub procedure equivalent to the used in Algorithm 1 for semantic filters, where the metric we evaluate is classification accuracy over the groups.
 4.4. sem_topk
 Performing a semantic top- ranking requires logically reasoning across rows, entailing joint reasoning over often large amounts of data.
-Implementing an efficient algorithm introduces many important design decisions pertaining to managing LM context length limits, grouping documents within each LM context window to achieve high quality LM-based comparisons over subsets of the data, and choosing an efficient top-k ranking algorithm to combine LM-based comparisons.
-These decisions have notable performance implications on efficiency and result quality, as we show in Section 4.
 While prior works have studied LM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n.
-d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, LOTUS’ implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low query execution time.
-We leverage several algorithmic design decisions and optimizations to achieve this.
-3.2.1. Algorithms
-Prior LM-based sorting and ranking works suggest several methods for performing LM-based comparisons.
-Specifically, prior works primarily study three classes of methods: point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), and pair-wise ranking methods (Shah and Wainwright, 2016; Qin et al., 2024).
-Point-wise methods output a relevance score for each document independently; while these methods are fully parallelizable across rows, prior work (Desai and Durrett, 2020) shows they provide poor accuracy due to difficulty in calibrating scores across prompts.
-We note that these methods can be implemented in LOTUS with a sem_map, but are not the focus of our semantic top- operator due to their quality limitations.
-By contrast, list-wise methods feed a partial list of 10 to 20 documents to the LM and prompt it to output a ranking over these, which can be aggregated using a sliding window approach.
-Unfortunately, prior works shows these methods are often prone to prediction failures (Sun et al., 2023) and sensitive to document ordering in the prompt (Qin et al., 2024).
-Pairwise-prompting methods instead offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label.
-This method has been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
-To aggregate pairwise comparisons, we consider several ranking algorithms, including a quadratic sorting algorithm, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm.
-The quadratic ranking algorithm obtains a comparison between each pair of input documents, and uses a win rate to determine a ranking over all elements before selecting the best.
-In contrast, the heap-based algorithm maintains a heap of size , and makes a linear pass over the data, updating the heap when more promising elements are found.
-Each time a new element is inserted or removed from the heap, the algorithm performs a series of sequential LM comparisons to update the data structure.
-Lastly the quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot.
-Because each round is fully parallelizable, we perform these LM-based comparisons efficiently in batch before recursing in the next round.
-3.2.2. Discussion
-As the core LM-based computation unit over subsets of the data, LOTUS’ current implementation uses pair-wise rankings, following recent prior work (Qin et al., 2024), which demonstrates its effectiveness.
-We also find this choice useful for implementing further optimizations, such as model cascades, described below.
-Figure 19 shows an example task-instruction prompt, which we use to implement the LM-based comparison by instructing the model to select between two documents according to the user-defined sorting criteria.
-Additionally, for aggregating comparisons, LOTUS’ current implementation leverages the quick-select-based (Hoare, 1961) top-k ranking algorithm, by default.
-Prior work (Qin et al., 2024) studies several alternative algorithms, including the heap-based ranking implementation and the quadratic sorting algorithm.
-We study these alternatives in Section 4, and find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for batched calls leading to lower execution time compared to a heap-based implementation.
-We believe future work may leverage these multiple algorithms to allow the user to declaratively trade-off between performance metrics, like accuracy, execution time, and cost.
+d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, our implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low execution time.
+Here we outline leverage several algorithmic design decisions for the gold algorithm as well as a lossless optimization.
+4.4.1. Gold Algorithm
+Two important algorithmic design decisions for the semantic top-k include how to implement the LM-based comparison and how to aggregate ranking information across LM-based comparisons.
+Our implementation uses pairwise LM-based comparisons for the former, and a quick-select based top-k algorithm (Hoare, 1961) for the latter.
+We briefly describe the reason for these choices and alternatives considered.
+Pairwise-prompting methods offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label.
+The two main classes of alternatives are point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), and list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), both of which have been shown to face quality issues (Sun et al., 2023; Qin et al., 2024; Desai and Durrett, 2020).
+In contrast, pairwise comparisons have been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
+In addition, we studied several alternative rank-aggregation algorithm, including quadratic sorting algorithms, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm, which we discuss in Section 5.
+The quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot.
+Because each round is fully parallelizable, we can efficiently batch these LM-based comparisons before recursing in the next round.
+In section 5, we find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for efficient batched inference, leading to lower execution time, compared to a heap-based implementation.
 4.4.2. Optimizations
-LOTUS’ top- algorithm is amenable to several optimizations.
-First, the top- algorithm’s pair-wise comparisons each invoke the LM to output a binary label, which is amenable to model cascades.
-To implement this, we use a simple an efficient scoring procedure based on log-probabilities of the generated tokens.
-We describe this procedure above for sem_filter (Section 3.1.1) and apply an equivalent procedure here.
-In Section 4, we demonstrate this optimization allows programmers to leverage small, cheap models in conjunction with large, expensive models to obtain high accuracy results at reduced cost.
-Additionally, LOTUS can leverage the semantic index to optimize pivot selection for some queries, rather than resorting to random pivot selection.
+We leverage the semantic index to optimize pivot selection for some queries, while incurring no accuracy loss.
 This optimization is useful when there exists correlation between the rankings imposed by the user’s arbitrary sorting criteria and the rankings imposed by semantic similarity scores.
@@
 This can reduce the number of LM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss.
-We believe fruitful future work will automatically estimate correlation between semantic similarity scores and the user-defined ranking criteria to transparently apply this optimization.
-3.3. sem_join
-Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left and right table.
-Implemented naively, this can be prohibitively expensive, incurring a quadratic number of LM calls with respect to to the size of the left and right join tables.
-LOTUS thus implements several algorithms suitable for a variety of settings, highlighting the rich design space, which we plan to further optimize over in future work.
-Here we describe the design patterns currently implemented, which we find to be useful for real-world tasks that we evaluate in Section 4.
-3.3.1. Algorithms
-The first join algorithm implements the nested-loop join pattern with efficient LM batch processing to maximize GPU utilization rather than naively looping over each pair of rows and invoking the LM.
-This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively.
-As Figure 18 shows, each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate.
-This quadratic join algorithm is suitable for small join tables.
-Alternatively, we implement a map-search-filter join pattern which can be used to approximate the quadratic join algorithm, while incurring fewer LM calls.
-This algorithm first performs a semantic mapping over the left join key to the domain of the right join key.
-In the example provided by Figure 20, which joins paper abstracts with the datasets they use, the map step would invoke the LM over each abstract, instructing the model to output the dataset used.
-This step is un-grounded in the sense that the LM is not given knowledge of right join table, but may optionally leverage user-specified demonstrations.
-Following the semantic projection, the algorithm then leverages the semantic index and performs a similarity search over the right join table to find candidates likely to pass the predicate.
-In the example, this search would retrieve a list of datasets from the right table for each paper, using semantic similarity of the datasets to the LM projection output in the first step.
-The semantic similarity search is parameterized by , which is automatically set based on the user’s specified LM call budget.
-Lastly, the algorithm performs a filter over the candidate pairs and outputs the joined table of tuples pairs that pass.
-This join algorithm partially mimics the information flow of prior work (D’Oosterlinck et al., 2024) and abstracts away intermediate steps by instead allowing users to specify an LM call budget for the join.
-The LM call complexity of this algorithm is .
-Lastly, we implement a search-filter join pattern.
-This algorithm first performs a semantic similarity join using a similarity index on the right join key, with the left join key embedded on the fly.
-In the example provided by Figure 20, this step would retrieve semantically similar datasets from the right table for each paper abstract in the left table.
-The batched semantic-similarity search sets the search parameter according to the user’s specified LM call budget, similar to the map-search-filter algorithm.
-The search results provide candidate tuple pairs, which are then evaluated in batch using the filter operation to gather the final set of rows that pass the user’s language predicate.
-This algorithm obtains an LM call complexity of .
-3.3.2. Discussion
-These join patterns offer performance tradeoffs suitable for different settings.
-The nested-loop algorithm offers an efficient solution when the join tables are sufficiently small, whereas the map-search-filter and search-filter patterns are suitable approximations that can apply efficiently over large tables, due to their linear LM call complexity in table size.
-The expected result quality of either approximation likely varies depending on the presence of predicate clustering and correlation, as defined by prior work (Patel et al., 2024) between the user’s predicate and the semantic embeddings.
-The search-filter pattern is likely to produce high quality results under positive query correlation, where entities that are semantically similar are also likely to pass the predicate.
-On the other hand, the map-search-filter pattern is likely to produce higher quality results when the predicate-embedding correlation is low using the original left join key’s embeddings as queries, but can be increased using the LM-projection over the left join key.
+In the case of no correlation between the langex-based ranking and similarity-based ranking, this method amounts to random pivot selection, and in the worst case of an adversarial pivot, the algorithm will incur one extra pivot round, which can impact execution time, but not degrade quality.
 4.5. sem_agg
-Performing semantic aggregations are inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows.
+Performing semantic aggregations is inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows.
 Thus, the operator’s implementation must efficiently orchestrate the LM over large amounts of data, while managing long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length.
-LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results and provide the programmer with flexibility.
-3.4.1. Algorithms
-LOTUS’ implementation builds on and generalizes the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n.
+LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results.
+4.5.1. Gold Algorithm
+Our implementation builds on the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n.
 d.]; lan, [n.
@@
 These implementations primarily leverage one of two aggregation patterns: either a fold pattern, which produces a sequential, linear pass over the data while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains.
-3.4.2. Discussion
-By default, LOTUS’ aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024).
-However, LOTUS’ partition function can be used to override the default functionality and generalizes the information flow of prior implementation by allowing users to specify arbitrary groupings over the data.
-Each unique group specified by the partition function will be aggregated with as few LM invocations as possible according to the model context length, before being merged with other groups.
-This allows the user to perform a fold pattern, as well as arbitrary user-defined patterns.
-3.4.3. Optimizations
-Preliminary results show that achieving high-quality semantic-aggregations is non-trivial and sensitive to the document’s input ordering and grouping.
-We see qualitative evidence of this in a summarization task of 50 ArXiv paper abstracts, which we show in Figure 21.
-We contrast the summarization results of a naive semantic aggregation (Figure 22) performed over the paper abstracts with the semantic aggregation performed using a partitioning function based on semantic similarity (Figure 23).
-The semantic aggregation using the partitioner first clusters documents based on semantic similarity, then aggregates each cluster before performing aggregations across clusters.
-As the figures show the two methods result in significantly different summaries.
-While the latter demonstrates qualitatively more cohesion and effectively abstracts general themes across papers, the former omits details and tends to list low-level details from individual papers rather than capturing higher-level themes.
-We leave a quantitative study of this to future work and believe that semantic aggregations create a rich design space for optimization.
-3.5. sem_map & sem_extract
-Both LOTUS’ semantic map and extract run batched LM calls over a set of rows, prompting the LM with the user’s arbitrary natural-language expression to generate a new column.
-Both functions can be fully parallelized over rows, and LOTUS implements efficient batched inference with vLLM (Kwon et al., 2023).
-To implement sem_extract, LOTUS prompts the model to answer the user’s langex with direct quotes, and the system then verifies that these snippets returned by the LM match the reference text.
-3.6. Semantic Indexing
-LOTUS supports several different algorithms for its semantic index.
-To generate the semantic similarity index, LOTUS’ sim_index operator first batch processes the user-specified column to generate semantic embeddings with the configured retriever model, then constructs a vector index over the semantic embeddings.
-The current implementation uses FAISS’ flat index by default and writes the index locally to disk.
-User’s can specify the local file path to store the index, and optionally specify alternative vector indices, such as hierarchical navigable small worlds (HNSW) (Malkov and Yashunin, 2018), inverted indices (IVF) (Baranchuk et al., 2018; Ge et al., 2014; Jégou et al., 2011; Johnson et al., 2017), and locality sensitive hash indices (LSH) (Andoni and Indyk, 2008; Andoni and Razenshteyn, 2015; Indyk and Motwani, 1998; Jafari et al., 2020; Li et al., 2020; Liu et al., 2021; Lu and Kudo, 2020; Park et al., 2015; Sundaram et al., 2013; Zheng et al., 2020; Andoni et al., 2015; Gionis et al., 1999; Gong et al., 2020).
-In the future, we envision additionally supporting a wider variety of embeddings stores and indices (Patel et al., 2024; Wang et al., 2021; vec, [n. d.]; noa, 2023b; chr, [n. d.]; mos, [n. d.]).
-LOTUS’ semantic search, similarity join and cluster operations all leverage the similarity index in their implementation.
-sem_search first embeds the user’s query string using the configured retriever model, then performs an efficient top-k search using the loaded FAISS index over the search column.
-sem_sim_join similarly performs a top-k search using the loaded FAISS index over the right key.
-Here the right-key may be a column of natural-language text, which LOTUS will embed using the retriever model on-the-fly, or an indexed column, for which LOTUS will load the previously-generated embeddings.
-The right-key embeddings then serve as the queries to perform batched search over the search column, given by the left key.
-Lastly, LOTUS implements sem_cluster_by using FAISS optimized kmeans library to cluster the user-specified column.
-Here, the column must be previously indexed, and LOTUS uses the generated embeddings to perform the clustering with the user-specified number of cluster centroids.
+Our aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024).
 5. Evaluation
-We now evaluate LOTUS’ programmability and efficiency through three diverse applications: fact-checking (Section 4.1), extreme multi-label classification (Section 4.2), and search and ranking (Section 4.3).
-For each of these applications, we see that state-of-the-art quality results are achievable with low development overhead, using LOTUS programs with a few lines of code and a few or even one semantic operator.
-In addition, our results demonstrate interesting implementation and optimization choices introduced by semantic operators.
-Specifically, we find that:
-On the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce FacTool (Chern et al., 2023), a recent state-of-the-art pipeline, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time.
-On the BioDEX dataset (D’Oosterlinck et al., 2023) for the extreme multi-label classification task, LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that runs up to faster than a naive join, demonstrating the power of LOTUS’ declarative interface.
-On the SciFact dataset (Thakur et al., 2021) and two newly-constructed paper datasets for the search and ranking application, LOTUS’s semantic top- operator achieves higher nDCG@10 than a vanilla retriever and re-ranker, while also running faster than alternative LM-based ranking methods used by prior works (Qin et al., 2024).
-Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama 3 models (int, [n.
+We now evaluate the expressiveness of the semantic operator model as well as the efficiency of our algorithms and optimizations through four diverse applications: fact-checking (Section 5.1), extreme multi-label classification (Section 5.2), search and ranking (Section 5.3), and paper analysis (Section 5.4).
+We find that the semantic operator model is expressive and easy to program with, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality.
+We also find that our optimizations that take advantage of the declarative nature of semantic operators can accelerate semantic filtering, top-k, group-by and join operators by up to while offering statistical accuracy guarantees.
+Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while executing up to 28 faster end-to-end.
+LOTUS’ current implementation extends the Pandas API (pan, [n.
+d.]).
+We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) to support efficient vector search for LOTUS’ semantic search, similarity join and cluster operations.
+Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama-3 models (int, [n.
 d.]), with a batch size of 64 running on vLLM (Kwon et al., 2023).
-For our experiments that use OpenAI’s GPT models (ope, [n.
-d.]), we run with 64-way thread parallelism.
+For our experiments that use OpenAI’s GPT-4o-2024-05-13 model (ope, [n.
+d.]), we control for cost by limiting execution to 64-way thread parallelism.
+We set temperature to , unless otherwise stated.
 5.1. Application: Fact-Checking
-4.1.1. Dataset
+Application.
 We evaluate on FEVER (Thorne et al., 2018), a claim verification dataset.
-We use the development dataset, which contains about 38,000 total claims, of which we sample 500 for our evaluation.
+We use the development dataset, which contains about 38,000 total claims, of which we sample 1,000 for our evaluation.
 Each claim is labeled with one of three labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”, and the task is to correctly determine the label of each claim, leveraging evidence from a corpus of million Wikipedia articles.
 We merge the latter two labels in to a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation.
-4.1.2. Baselines
+Baselines.
 FacTool (Chern et al., 2023) is a recent research work that proposes a multi-step pipeline for fact-checking involving, claim extraction, query generation, tool querying, evidence collection, and verification.
@@
 d.]).
-We evaluate the pipeline with both the default retrieval API, and alternatively test with a ColBERT (Khattab and Zaharia, 2020) index over the document corpus to perform retrieval.
+We evaluate the pipeline with both the default retrieval API, and with the open-source ColBERT (Khattab and Zaharia, 2020) index.
 We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the implemented LOTUS programs.
-| Method | Accuracy | Execution Time, batched (s) | Execution Time, no batching (s) | LoC |
-| FacTool | 83.5 | N/A | 1,174.8 | ¿ 750 |
-| LOTUS-Factool | 90.0 | 111.48 | 979.50 | ¡ 50 |
-| LOTUS-fact-filter | 90.5 | 64.28 | 206.45 | ¡ 50 |
-| LOTUS-fact-filter (+ cascades) | 93 | 34.09 | 168.24 | ¡ 50 |
-| LOTUS-fact-join | 86.5 | 2,394.22 | 12,923.08 | ¡ 50 |
-4.1.3. LOTUS Programs
+LOTUS Programs.
 We compose several intuitive LOTUS programs, each in less than 50 lines of code.
-For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama 70B as the primary LM, and Llama 8B and TinyLlama for cascade optimizations.
+For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama-70B as the primary LM, and Llama-8B as the cascade proxy.
 First, we compose a pipeline designed to directly re-implement FacTool’s information flow in LOTUS.
@@
 d.]), which include 3 demonstrations for generating search queries in the first sem_map, and chain-of-thought prompting in the second sem_map.
-The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the query pipeline, with LOTUS’ semantic filter operation.
+| Method | Accuracy | Execution Time (s), batched | Execution Time (s), no batching | LoC |
+| FacTool | 80.9 | N/A | 5396.11 | ¿ 750 |
+| LOTUS-Factool | 89.9 | 688.90 | 4,454.24 | ¡ 50 |
+| LOTUS-fact-filter | 91.2 | 329.1 | 988.95 | ¡ 50 |
+| LOTUS-fact-filter (opt.)* | 91.0 | 189.88 | 776.37 | ¡ 50 |
+| LOTUS-fact-join | 84.5 | 11,951.35 | 60,364.39 | ¡ 50 |
+The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the program, with a semantic filter.
 We use 3 demonstrations for the filter.
 Figure 7 shows the pseudocode for this program.
-We evaluate this program with and without model cascades for the semantic filter operation.
+We evaluate this program with and without model cascades applied to the semantic filter.
 Lastly, we compose an alternative pipeline, LOTUS-fact-join.
@@
 For the sem_map and join operations, we use 3 demonstrations each.
-4.1.4. Results
-Table LABEL:tab:factchecking demonstrates the powerful abstraction that LOTUS provides, allowing programmers to quickly write and test programs that compose simple operators to obtain state-of-the-art results.
-We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching.
-We see that FacTool’s implementation offers strong accuracy performance on the FEVER datasets, however the full repository required several hundred lines of code, highlighting the development burden of building these applications without abstractions for semantic-bulk processing.
+Results.
+Table 2 demonstrates the powerful abstraction that semantic operators provide, allowing programmers to quickly write and test programs that compose a few operators to obtain state-of-the-art results.
+We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching averaged over 10 runs.
+We see that FacTool’s implementation offers strong accuracy performance on the FEVER dataset, however the full repository was written from scratch in several hundred lines of code, representing a significant development burden.
 By contrast, each LOTUS program offers comparable or higher accuracy, in relatively few lines of code.
@@
 To provide an apples-to-apples comparison, we compare FacTool’s un-batched implementation to the LOTUS programs both with and without batching.
-First, we see from the Table LABEL:tab:factchecking that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with 6.5 points higher accuracy and lower execution time without batching.
-The LOTUS-FacTool implementation with batching further decreases execution time compared to the original FacTool implementation by .
+First, we see from the Table 2 that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with points higher accuracy and lower execution time without batching.
+The LOTUS-FacTool implementation with batching further decreases execution time compared to FacTool by .
 We see that the next LOTUS pipeline simply changes a single operation, switching the sem_map to a sem_filter, and maintains similar result quality to LOTUS-FacTool while further reducing execution time by in the batched implementation.
-Leveraging LOTUS’ filter operation allows the programmer to further optimize the program using model cascades, which increases accuracy by 3 points and increases the batched execution time by compared to LOTUS-FacTool.
-Figure 27 highlights the diverse performance trade-offs presented the model cascade optimizations used in LOTUS’ filter sem_filter.
-The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, Llama 70B, or TinyLlama (Zhang et al., 2024), shown by the circles.
-We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, or TinyLlama and Llama70B to implement the filter casacade.
-We generate multiple cascade points, shown by the stars, by varying the confidence threshold used.
+Leveraging LOTUS’ filter operation allows the programmer to further declaratively optimize the program using an approximation, which maintains high accuracy and reduces the batched execution time by compared to LOTUS-FacTool.
+Figure 9 highlights the diverse performance trade-offs presented by the model cascade optimizations used in LOTUS’ filter sem_filter.
+The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, or Llama 70B,
+shown by the circles.
+We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, to implement the filter with our casacade-based approximation.
+We generate multiple cascade points, shown by the stars, by varying the recall and precision targets of the filter operator, and setting the error probability .
 The plot shows that this filter optimization can substantially reduce execution time, and offer diverse accuracy trade-offs, compared to implementing the pipeline with the oracle model, Llama 70B, alone.
-Returning to Table LABEL:tab:factchecking, we find that the last LOTUS pipeline (search-map-join), reduces accuracy and increases execution time.
-Notably the performance trade-offs of different LOTUS programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines.
+We also verify that the algorithm’s probabilistic guarantees hold by testing 100 trials for recall and precision targets at .
+We find the targets are met in 99 of the 100 trials, satisfying and exceeding the guaranteed 0.8 success rate.
+Returning to Table 2, we find that the last LOTUS pipeline (LOTUS-fact-join), reduces accuracy and increases execution time.
+Notably the performance trade-offs of different AI programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines.
 Each LOTUS program we described can be implemented easily in relatively few lines of code, and the best performing one offers higher accuracy and lower execution time than FacTool’s original, un-batched implementation.
-| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
-| Semantic Similarity Join | 0.106 | 0.120 | 2.91 | 0.00 |
-| LOTUS Semantic Join (map-search-filter pattern) | 0.241 | 0.258 | 2,762 | 7,750 |
-| LOTUS Semantic Join (nested-loop pattern) | N/A | N/A | 2,144,560* | 6,092,500 |
-| LOTUS Semantic Join (search-filter pattern) | 0.155 | 0.186 | 2,640 | 7,500 |
-Estimated under linear-scaling assumption in number of batch calls.
 5.2. Application: Extreme Multi-label Classification
-4.2.1. Dataset
+Dataset.
 We evaluate on the Biodex Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and expert-created drug safety reports constructed from each article.
@@
 We sample 250 patient articles for our evaluation.
-Notably, there are approximately 24,000 possible drug-reaction labels, making this task an extreme multi-label classification task.
+Notably, there are possible drug-reaction labels, making this task an extreme multi-label classification task.
 Due to the large number of possible labels, leveraging an LM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024).
-We show below that this task can be efficiently modeled and programmed using LOTUS’ semantic join.
-4.2.2. Baselines
-We consider a simple retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels.
-We show the pseudocode for this program in Figure 28.
-4.2.3. LOTUS Programs
-The proposed LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 32.
-We perform the semantic join using at most 7 map demonstrations and set an LM call budget of 10,000.
-We use Llama-70b as the LM.
-4.2.4. Results
-Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while maintaining query efficiency.
-The table reports the rank-precision@5 (RP@5), rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls required for each program.
+We show below that this task can be efficiently modeled and programmed using the semantic join.
+Baselines We consider a retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels.
+| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
+| Sem-sim-join | 0.106 | 0.120 | 2.91 | 0.00 |
+| LOTUS Sem-join** | 0.244 | 0.261 | 5,891.6 | 15,107 |
+| nested-loop pattern | N/A | N/A | 2,144,560* | 6,092,500 |
+| map-sim-filter pattern** | 0.244 | 0.261 | 5,891.6 | 15,107 |
+| sim-filter pattern** | 0.154 | 0.191 | 24,206.8 | 63,724 |
+Estimated under linear-scaling assumption in number of batch calls.
+LOTUS Programs.
+Our LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 10.
+We perform the semantic join using at most 7 map demonstrations.
+We use Llama-70b as the LM and the E5Model (Wang et al., 2024) as the retriever for the semantic index.
+Results.
+Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while providing query efficiency compared to naive join implementations.
+The table reports the rank-precision@5 (RP@5) and rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls used.
 Since the nested-loop join pattern is prohibitively expensive to run, we show an estimate of the execution time assuming linear scaling in the number of batched calls.
-First we compare performance of the LOTUS join program implemented with the map-search-filter pattern to the baseline method, the semantic similarity join.
-As expected, the LOTUS program offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the retrieval-based similarity join.
+First we compare performance of the LOTUS join program, declaratively optimized, to the baseline, a semantic similarity join.
+As expected, the optimized LOTUS join offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the similarity join.
 This highlights the effectiveness of leveraging LMs over the data for complex reasoning-based tasks.
-We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4 turbo.
+We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4-turbo.
 D’Oosterlinck et al.
 report 24.73 RP@5 and 27.67 RP@10 for the compiled program, representing comparable result quality to LOTUS’ program, although notably the LOTUS program was not compiled with a prompt optimization system.
-We now consider several interesting performance trade-offs presented in the semantic join algorithm.
-We compare LOTUS map-search-filter join pattern to the naive nested-loop join algorithm and the search-filter join pattern shown in Table 3.
-We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is untenable and prohibitively costly.
-By contrast, the search-filter and map-search-filter pattern substantially reduce the LM call budget of the naive algorithm by , using an approximation.
-While these two approximation patterns have similar efficiency, according to execution time and the number of LM calls, interestingly, they offer substantially different result quality.
-Specifically, the map-search-filter pattern offers higher RP@5 and higher RP@10 compared to the retrieve-filter pattern on this task.
-These results highlight the unique opportunities bulk-semantic processing pipelines present for designing new algorithms and optimizations.
+We now consider several interesting performance trade-offs presented by semantic join algorithms.
+We compare the map-sim-filter join pattern, which was automatically selected in the LOTUS sem-join, to the naive nested-loop join algorithm and the sim-filter join pattern shown in Table 3.
+We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is prohibitively costly.
+By contrast, the sim-filter and map-sim-filter pattern substantially reduce the LM call budget of the naive algorithm by and respectively.
+LOTUS’ optimizer automatically selects the lower cost algorithm, in this case map-sim-filter.
+The selected pattern also results in significantly better accuracy due to its higher quality proxy score for this task.
+Specifically, the map-sim-filter pattern offers higher RP@5 and higher RP@10 compared to the sim-filter.
 5.3. Application: Search & Ranking
-4.3.1. Dataset
-We evaluate LOTUS’ performance on the search and ranking task using three datasets, including BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, as well as two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generate to evaluate more complex, reasoning-based ranking criteria over the data.
-For each, we report nDCG@10, as well as the execution time (ET) in seconds.
-The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance given each scientific claim.
-We sample 300 scientific facts for our evaluation, and report the average ranking execution time across these samples.
-While the SciFact dataset provides a sorting task based on a simple relevance criterion, our newly proposed benchmarks provide a more complex sorting criteria over a corpus of paper abstracts.
-Specifically, the ranking task is to find the papers that report the highest accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench respectively.
+Dataset.
+We evaluate LOTUS’ performance on the search and ranking task using three datasets: BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, and two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generated to evaluate more complex ranking criteria over the data.
+For each, we report nDCG@10 and execution time (ET) in seconds.
+The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance to a given claim.
+We sample 300 claims for our evaluation, and report the average ranking execution time across these samples.
+While the SciFact dataset provides a sorting task based on a simple relevance criterion, our generated benchmarks provide a more complex sorting criteria each over a corpus of paper abstracts.
+The task is to rank papers according to highest reported accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench, respectively.
 To generate CIFAR-bench, we took 100 abstracts from the Papers with Code Dataset (pap, [n.
-d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain the top-10.
+d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain ground truth.
 We then synthetically generated HellaSwag-bench by prompting Llama-70B to create 200 paper abstracts, each with a specified accuracy value, randomly sampled from .
-This setup allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth.
+These datasets allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth an ranking criteria other than relevance-based ones.
 We note that an alternative approach to these tasks could efficiently leverage sem_map to extract accuracy values on abstracts from either dataset, then perform a structured sort.
-However, our evaluation focuses on assessing the semantic ranking capabilities of LOTUS’ top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may likely generalize to a wider set of reasoning-based sorting queries, such as ”which paper makes the most outrageous claim”, for which ground truth is less objective to evaluate.
-For CIFAR-bench and HellaSwag-bench we report results for trials of the ranking task, at temperature , similar to prior works (Khattab et al., 2023)
-| Method | nDCG@10 | Execution Time (s) |
-| Semantic Search | 0.712 | 0.009 |
-| Semantic Search + Reranker | 0.741 | 2.64 |
-| LOTUS Top-k (quickselect + sem-index) - Llama-70B | 0.775 | 33.6 |
-| LOTUS Top-k (quickselect + sem-index) - GPT-4o | 0.800 | 11.2 |
+However, our evaluation focuses on assessing the semantic ranking capabilities of various top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may generalize to a wider set of reasoning-based ranking queries.
+We report results for trials at temperature , similar to prior works (Khattab et al., 2023)
+| Method | nDCG@10 | ET (s) |
+| Search | 0.712 | 0.009 |
+| Search + Reranker | 0.741 | 2.64 |
+| LOTUS Top-k - Llama-70B | 0.775 | 33.6 |
+| LOTUS Top-k - GPT-4o | 0.800 | 11.2 |
 | Method | CIFAR | HellaSwag | ||
-| nDCG@10 | Execution Time (s) | nDCG@10 | Execution Time (s) | |
-| Semantic Search | 0.252 | 0.008 | 0.119 | 0.008 |
-| Semantic Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
-| LOTUS Top-k (quickselect) - Llama 70-B | 0.746 | 41.3 | 0.909 | 59.1 |
-4.3.2. Baselines
-We consider two simple baselines.
-The first baseline performs semantic search, as shown in Figure 30, using the E5Model (Wang et al., 2024) for retrieval.
-The second baseline performs search with re-ranking, as the pseudocode shows in Figure 31, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking.
-4.3.3. LOTUS Programs
-The proposed LOTUS program performs a semantic top- over the documents, as shown in Figure 32, which shows example pseudocode for the CIFAR-bench dataset.
-The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets, as shown in the figure.
-We note that for the SciFact dataset, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk.
+| nDCG@10 | ET (s) | nDCG@10 | ET (s) | |
+| Search | 0.252 | 0.008 | 0.119 | 0.008 |
+| Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
+| LOTUS Top-k - Llama 70B | 0.710 | 39.6 | 0.975 | 63.6 |
+Baselines.
+We consider two simple baselines: a semantic search, using the E5Model (Wang et al., 2024) for retrieval, and a search with re-ranking, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking.
+We use choose these models because they are recent high quality embedding and re-rankers.
+LOTUS Program. Our LOTUS program performs a semantic top- over the documents.
+The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets.
+For SciFact, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk.
 We report results using both Llama-70B and GPT-4o as the LM.
-4.3.4. Results
 | Method | Scifact | CIFAR | HellaSwag | ||||||
@@
 | QuickSelect Top-k + Semantic Index | 0.775 | 33.6 | 229 | 0.710 | 39.6 | 307.7 | 0.975 | 63.6 | 672.7 |
-Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- operator for complex search tasks.
-In addition, Table 6 and Figure 33 highlight the rich implementation design space that this task presents.
+Results.
+Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- for complex search tasks.
+In addition, Table 6 highlights the rich implementation design space that this task presents.
 We walk through several note-worthy findings.
@@
 As expected, the improved result quality comes at a trade-off to query efficiency due to the cost of LM calls.
-Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipeline and trade-off result quality and efficiency depending on application-specific requirements.
+Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipelines and trade-off result quality and efficiency depending on application-specific requirements.
 Turning our attention to Table 5, we study LOTUS’ generality in supporting arbitrary language-based ranking criteria over the dataset.
 On the CIFAR-bench and HellaSwag-bench datasets, which use a complex sorting criteria, we see that both semantic search baselines with and without re-ranking provide poor result quality with consistently low nDCG@10.
-The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with 49.4 and 44.8 points higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively.
+The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with and higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively.
 These accuracy gains reflect the powerful reasoning capabilities of LMs efficiently orchestrated over the data.
 As expected, these significant accuracy gains come at an increase to execution time.
-We now analyze the efficiency of LOTUS semantic top- implementation along with it’s proposed optimizations.
+We now analyze the efficiency of our semantic top- implementation along with it’s proposed optimizations compared to alternatives.
 Table 6 compares several semantic top- algorithms, namely a quick-select top- algorithm, a quick-select top- that leverages the similarity index for pivot selection, a heap-based top- algorithm, and a quadratic sorting algorithm.
@@
 For this reason, our current implementation leverages the quick-select top- algorithm, although we envision future iterations may leverage multiple algorithms and allow the user to declaratively trade-off accuracy, query throughput, and cost.
-In addition to providing an efficient top- algorithm, the table also demonstrates the use of LOTUS’ similarity index for optimizing top- query performance.
+In addition to providing an efficient top- algorithm, the table also demonstrates our use of the similarity index for optimizing top- query performance.
 The quick-select top- algorithm optimized with the semantic similarity index for pivot selection demonstrates lower execution time at no accuracy loss on SciFact, where the ranking criteria correlates likely with semantic similarity.
 On the other hand, for the CIFAR-bench and HellaSwag-bench datasets, where the ranking criteria does not correlate with semantic similarity, we see that the similarity index has no significant impact on the accuracy or efficiency top- performance.
-Lastly, we analyze the impact of LOTUS’ cascade optimization applied to the quickselect top- implementation.
-Figure 33 compares the nDCG and execution time of implementing the operator with a single model, either Llama 8B or Llama 70B, to the implementation that leverages these models together using model cascades.
-We vary the confidence threshold of the cascade to generate several points in the trade-off space.
-We find that the cascade optimization offers diverse performance trade-offs, which can outperform the single-model oracle baseline, which uses Llama-70B.
-For instance, one cascade along the Pareto-frontier improves accuracy of the Llama-70B baseline by , while reducing execution time by , demonstrating substantial opportunities for automatically optimizing LOTUS’ semantic query pipelines.
+5.4. Application: ArXiv Paper Analysis
+Dataset.
+Lastly, we evaluate LOTUS’ use in analyzing a large corpus of documents, specifically focusing on the performance of sem_group_by and its approximation algorithm with accuracy guarantees.
+The dataset consists of 647 recent ArXiv articles we scraped, from the database (cs.DB), information retrieval (cs.IR), cryptography and security (cs.CR), and robotics (cs.RO) domains.
+The task is to group the papers according to the topics they discuss.
+Baselines.
+The sem_group_by consists of two sub-tasks: (a) discovering the groups and labels, and (b) classifying each document using the discovered labels.
+As a baseline for the latter, we consider a semantic similarity join between the labels and each abstract.
+We use the E5 embedding model for retrieval.
+IRIS Programs.
+We use the LOTUS sem_group_by operator to find 5 groups based on the topic abstracts.
+We use Llama-70B as the LM, the E5Model as the retrieval model, and 100 samples for the approximation algorithm described in Section 4.
+| () Advancements in Recommender Systems and Multimodal Data Integration |
+| () Advancements in Generative Information Retrieval Systems |
+| () Advancements in Large Language Models for Various Applications |
+| () Advancements in AI Security and Malware Detection Techniques |
+| () Advancements in Robotic Navigation and Manipulation Techniques |
+Results.
+We first qualitatively analyze the labels discovered, in the first phase of LOTUS’ implementation.
+As Figure LABEL:tab:groupby shows, the discovered labels intuitively align with the ArXiv domains included in the paper scrap, including important and timely topics related to recommendation systems, retrieval, LLM applications, AI security, and robotics.
+Performing this unsupervised group discovery took 44.03 seconds, representing a tractable and efficient implementation.
+Finally, turning our analysis to the classification subprocedure in our sem_group_by implementation, we evaluate the declarative algorithmic approximation.
+Figure 11 compares the execution time and classification accuracy with respect to the oracle implementation for the operator implemented with varying accuracy targets using .
+We see that the oracle implementation, at , offers the slowest execution time.
+On the other hand, the similarity-join baseline has faster execution time, but only accuracy with respect to the LM.
+We find our proxy-based optimization offers diverse performance trade-offs and allows the user to declaratively interpolate between these two extremes.
+We also note that the sampling procedure used to perform this optimization took less than 5 seconds, representing a small relatives cost.
 6. Related Work
-Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks or applications.
+Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks.
 Palimpzest (Liu et al., 2024b) presents a declarative approach to data cleaning and extract-transform-load (ETL) tasks.
 The authors propose to automatically optimize relational operators with LM calls, and implement two row-wise relational operators, a newly proposed convert operator, which can be transparently optimized to perform entity extraction using LLMs, and an AI-based filter operation, logically similar to LOTUS’ sem_filter.
-The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis to implement user queries, and proposes several others as future work.
+The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis and proposes several others as future work.
 SUQL (Liu et al., 2024c) presents a SQL extension to support conversational agents with knowledge grounding over structured and unstructured data.
@@
 Additionally, EVAPORATE performs efficient entity extraction from semi-structured data using LM-based code generation and weak supervision to ensemble candidate functions.
-LOTUS, in contrast to these prior works, defines a general-purpose programming model designed to capture broad-ranging applications with a core set of composable semantic operators, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions.
-In LOTUS’ current implementation, users can use sem_map to perform entity extraction over unstructured text fields, although future work may provide native functionality for entity extraction by integrating systems or optimizations of prior work.
-Additionally, we believe that several optimizations leveraged in prior work, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
-LLM UDFs
-Recent research work (Liu et al., 2024a) and existing analytical database vendors, such as
-Google BigQuery (ver, [n. d.]),
-Databricks (dat, [n. d.]) and
-AWS Redshift (noa, 2023a), alternatively offer LLM user-defined functions (UDFs) to the programmer.
-The LLM UDF programming model provides a lower-level interface, which is limited to logically row-wise LLM execution over the data, equivalent to LOTUS’ sem_map.
-In contrast, LOTUS’ programming model is declarative and provides a rich set of semantic operators, allowing the system to automatically orchestrate the LM to serve a variety of complex query patterns, including aggregations, ranking, and joins.
+In contrast to these works, semantic operators are a general-purpose query model designed to capture broad-ranging applications, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions, as an extension of the relational model.
+For example, users can use sem_map for row-wise entity extraction over unstructured text fields, but can also perform other operations over entire tables, such as sem_join.
+We also propose multiple execution algorithms and novel optimizations for these bulk operators (Section 4).
+Some of the optimizations in prior specialized systems, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
+LLM UDFs.
+Recent research (Liu et al., 2024a) and some existing DBMS vendors, such as
+BigQuery (ver, [n. d.]),
+Databricks (dat, [n.
+d.]) and Redshift (noa, 2023), offer LLM user-defined functions (UDFs) in SQL.
+LLM UDFs are a lower-level programming interface that is limited to row-wise LLM execution over the data, equivalent to our sem_map semantic operator.
+In contrast, our semantic operator programming model is high-level and declarative, allowing the system to automatically select which an algorithm to use and manage calls to the LM to support complex operations, including aggregations, ranking, and joins.
 Liu et al.
-(Liu et al., 2024a) study how to optimize LLM UDF functions, demonstrating performance gains with a de-duplication method and a prefix-sharing maximization method that reorders rows and re-formulates parameterized prompts to maximize key-value (KV) cache reuse during query execution.
-We believe these methods could be effective optimizations in future work at LOTUS’ batched execution layer.
-LM Programming Frameworks Recent LM programming frameworks, such as LangChain (lan, [n.
+(Liu et al., 2024a) study how to optimize LLM UDF calls, demonstrating performance gains with a de-duplication and prefix-sharing method that reorders rows and input fields to maximize key-value (KV) cache reuse during LLM calls.
+This is orthogonal to our optimizations but could likely accelerate batch LLM calls from LOTUS.
+LM Programming Frameworks.
+LM programming frameworks such as LangChain (lan, [n.
 d.]), LlamaIndex (lla, [n.
 d.]), and DSPy (Khattab et al., 2023), have gained significant popularity.
-These systems provide a set of abstractions for programming with LMs, including utilities for handling prompts and post-processsing LM outputs, and support for common use cases, such as RAG, chat-bots, and function-calling.
-DSPy focuses on abstracting LM pipelines as programs and provides automatic prompt optimization.
-In contrast to these systems, LOTUS’ programming model is designed for tasks involving bulk processing data with LLMs.
-While some of these systems support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized.
-ML-based Query Processing Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique opportunities for system design and optimization.
-MADLib (Hellerstein et al., 2012) extends SQL with efficient abstractions for supervised learning, unsupervised learning, and descriptive statistics.
+They provide abstractions for programming with LMs, including utilities for creating prompts and post-processing LM outputs, and modules for common calling patterns such as RAG, chat-bots, and agents.
+DSPy also optimizes a pipeline of LLM calls to maximize a target metric by tuning prompts and model weights.
+However, all these frameworks focus on expressing a pipeline of LLM calls that takes a single small input (or a stream of inputs for chat), as opposed to programs for processing large datasets.
+While some of them support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized.
+In contrast, semantic operators and LOTUS are designed for bulk semantic processing and introduce novel operators and optimizations for this task.
+ML-based Query Processing.
+Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique usage patterns and optimization opportunities.
+MADLib (Hellerstein et al., 2012) extends SQL with abstractions for machine learning and descriptive statistics.
 Prior works such as NoScope (Kang et al., 2017), TASTI (Kang et al., [n.
 d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML models over large datasets, typically in video analytics.
-Some optimizations, such as model cascades and predicate re-ordering, which were useful in these works are likewise useful for optimizing LOTUS pipelines with language models.
-However, our setting with natural-language reasoning tasks has new operators with significantly different semantics, such as sem_topk and sem_agg, requiring a new programming model as well as new query execution algorithms and optimizations.
+Some optimizations proposed in these works, such as model cascades and predicate re-ordering, are also useful for optimizing LOTUS pipelines with language models.
+However, our bulk semantic processing setting requires new operators with significantly different semantics, such as sem_join and sem_groupby, which were not optimized in this prior work.
+These motivate our novel optimizations that can approximate these operators with statistical guarantees (Section 4).
+Text2SQL A large body of prior work looks at Text2SQL (Yu et al., 2018; Yaghmazadeh et al., 2017; Zhang et al., 2024; Zelle and Mooney, 1996), using LMs to translate natural language queries to executable SQL code.
+These methods serve semantic queries that are expressible in traditioanl relational operators.
+By contrast, semantic operators provide new AI-based transformations, such as aggregations for summarizing large document sets, or rankings, joins and filters according to natural language expressions, which are not expressible in SQL.
+We note an interesting line of future work may build a natural language interface on top of semantic operators.
 7. Conclusion
@@
 We implement these operators in the LOTUS system to seamlessly extend the relational model and allow programmers to easily compose powerful reasoning-based query pipelines over vast corpora of structured and unstructured data.
-Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of LOTUS’ programming model as well as the efficiency and optimization opportunities of LOTUS’ query engine.
-For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized to achieve higher performance than existing implementations.
+Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of the semantic operator model as well as new opportunities for efficient algorithms and optimizations.
+For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized with accuracy guarantees to achieve higher performance than existing implementations.
 Acknowledgements.
@@
 - arx ([n. d.]) [n. d.]. arXiv.org ePrint archive. https://arxiv.org
-- chr ([n. d.]) [n. d.]. Chroma. https://www.trychroma.com/
 - cus ([n. d.]) [n. d.]. Custom Search JSON API | Programmable Search Engine. https://developers.google.com/custom-search/v1/overview
@@
 - ver ([n. d.]) [n. d.]. LLM with Vertex AI only using SQL queries in BigQuery. https://cloud.google.com/blog/products/ai-machine-learning/llm-with-vertex-ai-only-using-sql-queries-in-bigquery
-- mos ([n. d.]) [n. d.]. Mosaic AI Vector Search. https://docs.databricks.com
 - ope ([n. d.]) [n. d.]. OpenAI Platform. https://platform.openai.com
@@
 - lla ([n. d.]) [n. d.]. Querying - LlamaIndex 0.9.11.post1. https://docs.llamaindex.ai/en/stable/understanding/querying/querying.html
-- vec ([n. d.]) [n. d.]. Vector Database for Vector Search. https://www.pinecone.io/
-- noa (2023a) 2023a.
+- noa (2023) 2023.
 Large Language Models for sentiment analysis with Amazon Redshift ML (Preview) | AWS Big Data Blog.
 https://aws.amazon.com/blogs/big-data/large-language-models-for-sentiment-analysis-with-amazon-redshift-ml-preview/ Section: Amazon Redshift.
-- noa (2023b) 2023b. Vespa - the big data serving engine. https://vespa.ai/
 - noa (2024) 2024. mixedbread-ai/mxbai-rerank-large-v1 · Hugging Face. https://huggingface.co/mixedbread-ai/mxbai-rerank-large-v1
@@
 https://doi.org/10.18653/v1/2021.naacl-main.44
-- Andoni and Indyk (2008) Alexandr Andoni and Piotr Indyk.
-2008.
-Near-optimal hashing algorithms for approximate nearest neighbor in high dimensions.
-Commun.
-ACM 51, 1 (Jan.
-2008), 117–122.
-https://doi.org/10.1145/1327452.1327494
-- Andoni et al.
-(2015) Alexandr Andoni, Piotr Indyk, Thijs Laarhoven, Ilya Razenshteyn, and Ludwig Schmidt.
-2015.
-Practical and optimal LSH for angular distance.
-In Proceedings of the 28th International Conference on Neural Information Processing Systems - Volume 1 (NIPS’15).
-MIT Press, Cambridge, MA, USA, 1225–1233.
-- Andoni and Razenshteyn (2015) Alexandr Andoni and Ilya Razenshteyn.
-2015.
-Optimal Data-Dependent Hashing for Approximate Near Neighbors.
-In Proceedings of the forty-seventh annual ACM symposium on Theory of Computing (STOC ’15).
-Association for Computing Machinery, New York, NY, USA, 793–801.
-https://doi.org/10.1145/2746539.2746553
 - Arora et al.
@@
 http://arxiv.org/abs/2304.09433 arXiv:2304.09433 [cs].
-- Baranchuk et al.
-(2018) Dmitry Baranchuk, Artem Babenko, and Yury Malkov.
-2018.
-Revisiting the Inverted Indices for Billion-Scale Approximate Nearest Neighbors.
-https://doi.org/10.48550/arXiv.1802.02422 arXiv:1802.02422 [cs].
 - Braverman and Mossel ([n. d.]) Mark Braverman and Elchanan Mossel. [n. d.]. Noisy sorting without resampling. ([n. d.]).
@@
 https://doi.org/10.48550/arXiv.2307.13528 arXiv:2307.13528 [cs].
+- Dasgupta ([n. d.]) Sanjoy Dasgupta. [n. d.]. The hardness of k-means clustering. ([n. d.]).
 - Desai and Durrett (2020) Shrey Desai and Greg Durrett. 2020. Calibration of Pre-trained Transformers. https://arxiv.org/abs/2003.07892v3
@@
 https://doi.org/10.48550/arXiv.2210.08726 arXiv:2210.08726 [cs].
-- Ge et al.
-(2014) Tiezheng Ge, Kaiming He, Qifa Ke, and Jian Sun.
-2014.
-Optimized Product Quantization.
-IEEE Transactions on Pattern Analysis and Machine Intelligence 36, 4 (April 2014), 744–755.
-https://doi.org/10.1109/TPAMI.2013.240 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
-- Gionis et al.
-(1999) Aristides Gionis, Piotr Indyk, and Rajeev Motwani.
-1999.
-Similarity Search in High Dimensions via Hashing.
-In Proceedings of the 25th International Conference on Very Large Data Bases (VLDB ’99).
-Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 518–529.
-- Gong et al.
-(2020) Long Gong, Huayi Wang, Mitsunori Ogihara, and Jun Xu.
-2020.
-iDEC: indexable distance estimating codes for approximate nearest neighbor search.
-Proceedings of the VLDB Endowment 13, 9 (May 2020), 1483–1497.
-https://doi.org/10.14778/3397230.3397243
 - Hellerstein et al.
@@
 - Hoare (1961) C. A. R. Hoare. 1961. Algorithm 65: find. Commun. ACM 4, 7 (July 1961), 321–322. https://doi.org/10.1145/366622.366647
-- Indyk and Motwani (1998) Piotr Indyk and Rajeev Motwani.
-1998.
-Approximate nearest neighbors: towards removing the curse of dimensionality.
-In Proceedings of the thirtieth annual ACM symposium on Theory of computing (STOC ’98).
-Association for Computing Machinery, New York, NY, USA, 604–613.
-https://doi.org/10.1145/276698.276876
 - Izacard et al.
@@
 https://doi.org/10.48550/arXiv.2208.03299 arXiv:2208.03299 [cs].
-- Jafari et al.
-(2020) Omid Jafari, Parth Nagarkar, and Jonathan Montaño.
-2020.
-mmLSH: A Practical and Efficient Technique for Processing Approximate Nearest Neighbor Queries on Multimedia Data.
-In Similarity Search and Applications (Lecture Notes in Computer Science), Shin’ichi Satoh, Lucia Vadicamo, Arthur Zimek, Fabio Carrara, Ilaria Bartolini, Martin Aumüller, Björn Þór Jónsson, and Rasmus Pagh (Eds.).
-Springer International Publishing, Cham, 47–61.
-https://doi.org/10.1007/978-3-030-60936-8_4
 - Johnson et al. (2017) Jeff Johnson, Matthijs Douze, and Hervé Jégou. 2017. Billion-scale similarity search with GPUs. http://arxiv.org/abs/1702.08734 arXiv:1702.08734 [cs].
-- Jégou et al.
-(2011) Herve Jégou, Matthijs Douze, and Cordelia Schmid.
-2011.
-Product Quantization for Nearest Neighbor Search.
-IEEE Transactions on Pattern Analysis and Machine Intelligence 33, 1 (Jan.
-2011), 117–128.
-https://doi.org/10.1109/TPAMI.2010.57 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
 - Kang et al.
@@
 https://doi.org/10.48550/arXiv.2004.00827 arXiv:2004.00827 [cs].
+- Kang et al.
+(2021) Daniel Kang, John Guibas, Peter Bailis, Tatsunori Hashimoto, Yi Sun, and Matei Zaharia.
+2021.
+Accelerating approximate aggregation queries with expensive predicates.
+Proceedings of the VLDB Endowment 14, 11 (July 2021), 2341–2354.
+https://doi.org/10.14778/3476249.3476285
 - Kang et al.
@@
 https://doi.org/10.18653/v1/P19-1612
-- Li et al.
-(2020) Mingjie Li, Ying Zhang, Yifang Sun, Wei Wang, Ivor W.
-Tsang, and Xuemin Lin.
-2020.
-I/O Efficient Approximate Nearest Neighbour Search based on Learned Functions.
-2020 IEEE 36th International Conference on Data Engineering (ICDE) (April 2020), 289–300.
-https://doi.org/10.1109/ICDE48307.2020.00032 Conference Name: 2020 IEEE 36th International Conference on Data Engineering (ICDE) ISBN: 9781728129037 Place: Dallas, TX, USA Publisher: IEEE.
 - Liang et al.
@@
 https://doi.org/10.48550/arXiv.2311.09818 arXiv:2311.09818 [cs].
-- Liu et al.
-(2021) Wanqi Liu, Hanchen Wang, Ying Zhang, Wei Wang, Lu Qin, and Xuemin Lin.
-2021.
-EI-LSH: An early-termination driven I/O efficient incremental c-approximate nearest neighbor search.
-The VLDB Journal 30, 2 (March 2021), 215–235.
-https://doi.org/10.1007/s00778-020-00635-4
-- Lu and Kudo (2020) Kejing Lu and Mineichi Kudo.
-2020.
-R2LSH: A Nearest Neighbor Search Scheme Based on Two-dimensional Projected Spaces.
-In 2020 IEEE 36th International Conference on Data Engineering (ICDE).
-1045–1056.
-https://doi.org/10.1109/ICDE48307.2020.00095 ISSN: 2375-026X.
 - Lu et al.
@@
 - Ma et al. (2023) Xueguang Ma, Xinyu Zhang, Ronak Pradeep, and Jimmy Lin. 2023. Zero-Shot Listwise Document Reranking with a Large Language Model. https://arxiv.org/abs/2305.02156v1
-- Malkov and Yashunin (2018) Yu A.
-Malkov and D.
-A.
-Yashunin.
-2018.
-Efficient and robust approximate nearest neighbor search using Hierarchical Navigable Small World graphs.
-http://arxiv.org/abs/1603.09320 arXiv:1603.09320 [cs].
-- Park et al.
-(2015) Yongjoo Park, Michael Cafarella, and Barzan Mozafari.
-2015.
-Neighbor-sensitive hashing.
-Proceedings of the VLDB Endowment 9, 3 (Nov.
-2015), 144–155.
-https://doi.org/10.14778/2850583.2850589
 - Patel et al.
@@
 https://arxiv.org/abs/2304.09542v2
-- Sundaram et al.
-(2013) Narayanan Sundaram, Aizana Turmukhametova, Nadathur Satish, Todd Mostak, Piotr Indyk, Samuel Madden, and Pradeep Dubey.
-2013.
-Streaming similarity search over one billion tweets using parallel locality-sensitive hashing.
-Proceedings of the VLDB Endowment 6, 14 (Sept.
-2013), 1930–1941.
-https://doi.org/10.14778/2556549.2556574
 - Thakur et al.
@@
 - Wang et al.
-(2021) Jianguo Wang, Xiaomeng Yi, Rentong Guo, Hai Jin, Peng Xu, Shengjun Li, Xiangyu Wang, Xiangzhou Guo, Chengming Li, Xiaohai Xu, Kun Yu, Yuxing Yuan, Yinghao Zou, Jiquan Long, Yudong Cai, Zhenxiang Li, Zhifeng Zhang, Yihua Mo, Jun Gu, Ruiyi Jiang, Yi Wei, and Charles Xie.
-2021.
-Milvus: A Purpose-Built Vector Data Management System.
-In Proceedings of the 2021 International Conference on Management of Data (SIGMOD ’21).
-Association for Computing Machinery, New York, NY, USA, 2614–2627.
-https://doi.org/10.1145/3448016.3457550
-- Wang et al.
 (2024) Liang Wang, Nan Yang, Xiaolong Huang, Binxing Jiao, Linjun Yang, Daxin Jiang, Rangan Majumder, and Furu Wei.
@@
 https://arxiv.org/abs/2404.13207v2
+- Yaghmazadeh et al.
+(2017) Navid Yaghmazadeh, View Profile, Yuepeng Wang, View Profile, Isil Dillig, View Profile, Thomas Dillig, and View Profile.
+2017.
+SQLizer: query synthesis from natural language.
+Proceedings of the ACM on Programming Languages 1, OOPSLA (Oct.
+2017), 1–26.
+https://doi.org/10.1145/3133887 Publisher: Association for Computing Machinery.
+- Yu et al.
+(2018) Tao Yu, Zifan Li, Zilin Zhang, Rui Zhang, and Dragomir Radev.
+2018.
+TypeSQL: Knowledge-based Type-Aware Neural Text-to-SQL Generation.
+https://doi.org/10.48550/arXiv.1804.09769 arXiv:1804.09769 [cs].
 - Yue et al.
@@
 - Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah D. Goodman. 2022. STaR: Bootstrapping Reasoning With Reasoning. https://doi.org/10.48550/arXiv.2203.14465 arXiv:2203.14465 [cs].
-- Zhang et al. (2024) Peiyuan Zhang, Guangtao Zeng, Tianduo Wang, and Wei Lu. 2024. TinyLlama: An Open-Source Small Language Model. arXiv:2401.02385 [cs.CL] https://arxiv.org/abs/2401.02385
+- Zelle and Mooney (1996) John M Zelle and Raymond J Mooney. 1996. 1996-Learning to Parse Database Queries Using Inductive Logic Programming. (1996).
+- Zhang et al.
+(2024) Bin Zhang, Yuxiao Ye, Guoqing Du, Xiaoru Hu, Zhishuai Li, Sun Yang, Chi Harold Liu, Rui Zhao, Ziyue Li, and Hangyu Mao.
+2024.
+Benchmarking the Text-to-SQL Capability of Large Language Models: A Comprehensive Evaluation.
+https://doi.org/10.48550/arXiv.2403.02951 arXiv:2403.02951 [cs].
 - Zhang et al.
@@
 https://doi.org/10.48550/arXiv.2210.03493 arXiv:2210.03493 [cs].
-- Zheng et al.
-(2020) Bolong Zheng, Xi Zhao, Lianggui Weng, Nguyen Quoc Viet Hung, Hang Liu, and Christian S.
-Jensen.
-2020.
-PM-LSH: A fast and accurate LSH framework for high-dimensional approximate NN search.
-Proceedings of the VLDB Endowment 13, 5 (Jan.
-2020), 643–655.
-https://doi.org/10.14778/3377369.3377374

### v3

@@
-Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data
+Semantic Operators: A Declarative Model for Rich, AI-based Data Processing
 Abstract.
-The semantic capabilities of language models (LMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora.
-Unfortunately, existing systems lack high-level abstractions to perform bulk semantic queries across large corpora.
-We introduce semantic operators, a declarative programming interface that extends the relational model with composable AI-based operations for bulk semantic queries (e.g., filtering, sorting, joining or aggregating records using natural language criteria).
-Each operator can be implemented and optimized in multiple ways, opening a rich space for execution plans similar to relational operators.
-We implement our operators in LOTUS, an open source query engine with a DataFrame API.
-Furthermore, we develop several novel optimizations that take advantage of the declarative nature of semantic operators to accelerate semantic filtering, clustering and join operators by up to while offering statistical accuracy guarantees.
-We demonstrate LOTUS’ effectiveness on real AI applications including fact-checking, extreme multi-label classification, and search.
-We show that the semantic operator model is expressive, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality.
-Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while running up to 28 faster.
-LOTUS is publicly available at https://github.com/stanford-futuredata/lotus.
+The semantic capabilities of large language models (LLMs) have the potential to enable rich analytics and reasoning over vast knowledge corpora.
+Unfortunately, existing systems either empirically optimize expensive LLM-powered operations with no performance guarantees, or serve a limited set of row-wise LLM operations, providing limited robustness, expressiveness and usability.
+We introduce semantic operators, the first formalism for declarative and general-purpose AI-based transformations based on natural language specifications (e.g., filtering, sorting, joining or aggregating records using natural language criteria).
+Each operator opens a rich space for execution plans, similar to relational operators.
+Our model specifies the expected behavior of each operator with a high-quality gold algorithm, and we develop an optimization framework that reduces cost, while providing accuracy guarantees with respect to a gold algorithm.
+Using this approach, we propose several novel optimizations to accelerate semantic filtering, joining, group-by and top-k operations by up to .
+We implement semantic operators in the LOTUS system and demonstrate LOTUS’ effectiveness on real, bulk-semantic processing applications, including fact-checking, biomedical multi-label classification, search, and topic analysis.
+We show that the semantic operator model is expressive, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that match or exceed quality of recent LLM-based analytic systems by up to , while offering accuracy guarantees.
+Overall, LOTUS programs match or exceed the accuracy of state-of-the-art AI pipelines for each task while running up to faster than the highest-quality baselines.
+LOTUS is publicly available at https://github.com/lotus-data/lotus.
 1. Introduction
-The powerful semantic capabilities of modern language models (LMs) create exciting opportunities for building AI-based analytics systems that reason over vast knowledge corpora.
+The powerful semantic capabilities of modern language models (LLMs) create exciting opportunities for building AI-based analytics systems that reason over vast knowledge corpora.
 Many applications require complex reasoning over large amounts of data, including both unstructured and structured data.
-For example a researcher reviewing recent ArXiv (arx, [n.
-d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset.
+For example a researcher reviewing recent ArXiv (arx, [n.d.]) preprints may want to quickly obtain a summary of relevant papers from the past week, or find the papers that report the best performance for a particular task and dataset.
 Similarly, a medical professional may automatically extract biomedical characteristics and candidate diagnoses from many patient reports (D’Oosterlinck et al., 2023).
-Likewise, organizations wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business (dis, [n.
-d.]).
+Likewise, organizations wish to automatically digest lengthy transcripts from internal meeting transcripts and chat histories to validate hypotheses about their business (dis, [n.d.]).
 Each of these tasks require a form of bulk semantic processing, where the analytics system must process large amounts of data and apply semantic-based analysis across a whole dataset.
@@
 This prospect, however, raises two challenging questions: first, how should developers express semantic queries, and secondly, how should we design the underlying analytics system to achieve high efficiency and accuracy.
-Unfortunately, existing systems are insufficient for applications that require bulk semantic processing.
-Several systems provide logically row-wise LM operators, either in the form of a general-prupose LLM user-defined function (UDF) (ver, [n.
-d.]; dat, [n.
-d.]; noa, 2023; Liu et al., 2024b) or specialized row-wise operators (Liu et al., 2024c, a; Lin et al., 2024) for specific applications, such as conversational chat or data cleaning and extract-transform-load (ETL) tasks.
-These programming models fail to support a range of LM-based transformations across rows, such as ranking, grouping or joining records.
-Alternatively, many LM programming frameworks (lan, [n.
-d.]; lla, [n.
-d.]; Khattab et al., 2023) and research works (Khattab et al., 2021; Lee et al., 2019; Anantha et al., 2021; Gao et al., 2023; Izacard et al., 2022; Zelikman et al., 2022; Zhang et al., 2022) support combining LLM calls with retrieval-augmented generation (RAG) and Text2SQL, however these models are limited to point lookups to the data.
-RAG and Text2SQL methods first retrieve data using semantic search or LM-generated SQL calls, and then process the results with an LLM.
-Both are limited to analyses where an LLM can be fed a small number of documents or rows that are placed into a single model call’s context window (e.g., questions answered in one document in the corpus).
-However, bulk semantic processing applications may involve analysis across multiple rows, for clustering, ranking or joining an entire dataset.
-Towards a declarative programming interface for bulk semantic processing, we propose semantic operators, which extend the relational model with AI-based operations that users can compose into powerful, reasoning-based query pipelines over structured and unstructured data.
-These operators include semantic filters, joins, rankings, aggregations, and projections, which take natural language expressions given by the programmer.
-To provide an implementation of these operators, we present the LOTUS system, which exposes these operators through a simple DataFrame-based programming interface.
-LOTUS’ query engine efficiently and accurately executes queries with semantic operators, while abstracting away low-level details like model context length limits and algorithmic choices.
-Finally, we propose novel optimizations for the semantic filter, join, ranking and groupby operators that can improve performance by up to over a naive, expensive implementation while guaranteeing accuracy similar to the naive implementation.
-| Operator | Description |
-| Returns the tuples in a table that pass the provided langex predicate. | |
-| Joins a table against a second table by keeping all pairs of tuples that pass the provided langex predicate. | |
-| Performs an aggregation over the input tuples according to the langex, which specifies a commutative, associative aggregation function over a list of tuples. | |
-| Ranks each tuple and returns the best according to the langex, which specifies a ranking function that sorts a list of tuples. | |
-| Groups the tuples into categories based on the langex, which specifies a grouping criteria. | |
-| Performs a projection, returning a new column, according to the provided langex. | |
-Figure 1 begins to demonstrate the power of LOTUS’ declarative programming model and optimized query engine.
-For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we write an intuitive LOTUS program in less than 50 lines of code by composing semantic operators (i.e., filters, maps, and joins).
-The modularity and composability of these semantic operators make it easy to quickly write programs and explore the design space on this task.
-In doing so, we find the un-optimized LOTUS programs can reproduce and improve accuracy on this task by percentage points compared to a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), which was written in over 750 lines of code.
-LOTUS also provides query efficiency by default, yielding lower execution time than the FacTool implementation.
-LOTUS’ optimizer exploits the rich implementation design space of semantic operators to provide novel optimizations for bulk semantic operations.
-To demonstrate the power of programming with semantic operators, we implement optimizations for several costly operators, i.e., semantic filter, join, ranking and group-by.
-Specifically, we develop a novel, lossless optimization for ranking.
-Additionally, for filters, joins and group-bys, we define and implement both a tractable, but expensive, reference algorithm as well as an approximation algorithms with probabilistic guarantees on accuracy with respect to the reference algorithm.
-Our approximation algorithms leverage build on statistical techniques used in prior works (Kang et al., 2021, 2020) with novel and efficient proxy scorers that leverage small LMs or semantic embeddings.
-Figure 1 demonstrates the effectiveness of these methods.
-The program implemented using LOTUS’ optimizations with recall and precision targets and a error probability attains high accuracy, of the unoptimized program’s accuracy, and runs faster than the unoptimized program.
-We systematically evaluate the expressiveness of the semantic operator model and the efficiency of our algorithms and optimizations with statistical accuracy guarantees through several real applications, including fact-checking, extreme multi-label classification, and search, which had previously been implemented by hand in AI research papers.
-We show that LOTUS’ programming model is highly expressive, capturing high quality and state-of-the-art query pipelines with low development overhead for these wide-ranging applications.
-Furthermore, our algorithms and optimizations can speed up query execution by up to while offering statistical guarantees on accuracy.
-Specifically, on the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce a recent state-of-the-art pipeline (Chern et al., 2023), as shown in Figure 1, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering or lower execution time with or without batching, respectively, while providing statistical accuracy guarantees.
-In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that provides lower execution time than the naive algorithm, while again offering statistical guarantees on accuracy.
-In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the vanilla retriever and re-ranker, while also providing query efficiency, with lower execution time than LM-based ranking methods (Qin et al., 2024) used by prior works.
-Our main contributions are the following:
-We propose semantic operators as a programming model for AI-based analytics.
-Semantic operators (e.g., filters, joins, top-k ranking, aggregations, and group-bys) are declarative and can be implemented in multiple ways, creating a rich design space for execution plans.
-We present efficient algorithms and query optimizations for several costly operators, i.e., semantic filter, join, top-k ranking and group-by.
-We develop a novel, lossless optimization for ranking and approximations for filters, joins and group-bys that provide statistical accuracy guarantees with respect to tractable, but expensive, reference algorithms.
-We provide a reference implementation of semantic operators and optimizations in the open-source LOTUS system.
-2. Semantic Operators
-We define semantic operators as declarative transformations on one or more relations, parameterized by a natural language expression, that can be implemented by a variety of AI-based algorithms.
-Semantic operators provide a seamless extension to the relational model, operating over tables that may contain traditional structured data as well as unstructured fields such as free-form text.
-These new operators can be composed with traditional relational operators to build rich queries that can be transparently optimized, which we discuss in Section 4.
-Table 1 lists a core set of semantic operators we propose in this paper, which cover the common semantic transformations we observe in real-world applications and mirror the key transformation patterns in standard relational operators.
-Specific analytics systems may of course provide additional semantic operators or APIs that combine the ones we discuss here.
-Each semantic operator takes a parameterized natural language expressions (langex for short), which are natural language expressions that specify a function over one or more attributes for the given semantic transformation.
-The langex signature depends on its semantic operator and may represent a predicate, aggregation, comparator function, or projection in natural language.
-Figure 2 shows an example of using semantic operators in LOTUS’ dataframe API to load data about ArXiv papers and perform a summarization task, which finds relevant papers using a top-k search, filters according to whether the paper claims to outperform a specific baseline, and then summarizes the remaining papers into a single digest with an aggregation.
-The code in Figure 2 highlights various langex signatures: the sem_filter langex provides a predicate that indicates a filter criteria to apply over the title and abstract attributes, while sem_agg takes a langex that provides an associative aggregation expression, which here indicates a many-to-one summarization task over abstracts.
-Together, each langex and semantic operator specifies a modular semantic transformation, whose behavior is defined with respect to a world model , which may be a powerful LLM that captures a probability distribution over the vocabulary .
-We briefly overview several semantic operators, both ones whose definitions directly extend their analogs in traditional relational algebra, and those with more complex definitions.
-Semantic Filter is a unary operator over the relation and returns the relation , where the natural language predicate is specified by the langex with respect to model .
-The langex is paramterized by one or more attributes from the relation that are relevant for evaluating the predicate.
-Semantic Join provides a binary operator over relations and to return the relation , once again taking the langex with respect to model .
+Unfortunately, existing systems are insufficient for bulk semantic processing, either limiting their expressiveness to row-wise LLM execution models or focusing on empirical optimizations without any formalism or accuracy guarantees, leading to unexpected behaviors and frequent failures.
+First, several systems provide logically row-wise LLM operators, either in the form of a general-purpose AI user-defined functions (UDF) (MotherDuck, [n.d.]; WilliamDAssafMSFT, 2024; noa, [n.d.]d; ver, [n.d.]; dat, [n.d.]; noa, 2023; Liu et al., 2024a) or specialized row-wise operators (Liu et al., 2024c, b; Lin et al., 2024) for specific applications, such as conversational chat, data cleaning and extract-transform-load (ETL) tasks.
+These programming models fail to support a range of LLM-based transformations across rows, such as ranking, grouping or joining records.
+Alternatively, more recent LLM-based analytics systems, such as DocETL (Shankar et al., 2024) and UQE (Dai et al., 2024), empirically optimize LLM-based operations with no accuracy guarantees, lacking any formalism to define correct behavior, which hinders their robustness and usability, as we show in Section 5.
+These obstacles highlight a core challenge of integrating semantic-based processing within reliable query systems due to the inherent ambiguity of natural language instructions and LLM outputs.
+Towards robust, declarative query systems for bulk semantic processing, we propose the semantic operator model, which extends the relational model with AI-based operations.
+Semantic operators provide the first formalism for general-purpose and multi-row, AI-based transformation, including semantic filters, joins, top-k rankings, aggregations, and projections.
+Each operator takes a concise natural language signature, given by the programmer, and its behavior is fully specified by a tractable, high-quality ”gold” algorithm, which indicates how to orchestrate the underlying AI model over the data.
+Our optimization approach then exploits the rich design space of semantic operator execution plans to reduce cost, while providing statistical accuracy guarantees with respect to the gold algorithm (i.e., guarantees that the output of the optimized operator will be similar to that of the gold algorithm).
+We propose several novel optimizations for the semantic filter, join, top-k and group-by operators.
+Our methods include approximation algorithms, building on statistical techniques used in prior works (Kang et al., 2021, 2020) with novel and efficient proxy scorers using small LLMs or semantic embeddings.
+We implement semantic operators in LOTUS (LLMs Over Tables of Unstructured and Structured data), an open source system that exposes these operators through a simple DataFrame-based programming interface.
+| Operator | Description | Definition | Gold Algorithm |
+| Returns the tuples that pass the langex predicate. | Compute | ||
+| Joins a table against a second table by keeping all tuple pairs that pass the langex predicate. | Compute | ||
+| Aggregates input tuples according to the langex reducer function. | Perform a reduce algorithm, recursively computing , | ||
+| Returns an ordered list of the best tuples according to the langex ranking criteria. | Perform top-k sorting algorithm using pairwise comparisons, | ||
+| Groups the tuples into categories based on the langex grouping criteria. | Obtain centers with a clustering algorithm, and perform pointwise assignments | ||
+| Performs the projection specified by the langex. | Compute | ||
+Figure 1 illustrates the power of LOTUS’ declarative programming model and optimized query engine.
+For a fact-checking task on the FEVER dataset (Thorne et al., 2018), we write an intuitive LOTUS program to re-implement a recent state-of-the-art fact-checking pipeline, FacTool (Chern et al., 2023), which was originally implemented in over 750 lines of code.
+Our un-optimized LOTUS program composes semantic operators (filters, maps and joins) and achieves higher accuracy on this task.
+LOTUS also provides query efficiency by default, yielding lower execution time than the original FacTool implementation.
+Furthermore, the optimized LOTUS program, configured with accuracy targets and a allowed error probability, runs faster and obtains accuracy compared to the un-optimized LOTUS program.
+We systematically evaluate LOTUS on four real bulk-semantic processing applications: fact-checking, biomedical multi-label classification, search, and topic analysis.
+These include applications expressible and studied by recent LLM-based analytics systems (Shankar et al., 2024; Dai et al., 2024), where we compare against those systems, as well as others, where we compare against hand-designed pipelines from the AI literature.
+We demonstrate that the semantic operator model is effective, capturing state-of-the-art AI pipelines in a few operator calls that reliably match or exceed quality of recent LLM-based analytics systems by up to , while offering accuracy guarantees and running up to faster than the best-quality baselines.
+Furthermore, our gold algorithms provide high-quality results, and our optimizations speed up query execution by up to with respect to them.
+Specifically, our LOTUS program re-implements FacTool’s recent state-of-the-art fact-checking pipeline (Chern et al., 2023), as shown in Figure 1, and achieves higher accuracy with lower execution time on the FEVER dataset (Thorne et al., 2018), outperforming alternative LLM-powered analytics systems on both accuracy and efficiency.
+In the extreme multi-label classification task on the BioDEX dataset (D’Oosterlinck et al., 2023), LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) and the LLM-powered analytics baselines, while providing accuracy guarantees and speedups with respect to our gold algorithm.
+In the search and ranking application, LOTUS allows a simple composition of operators to achieve higher nDCG@10 than the next best performing baselines, while also providing query efficiency, with lower execution time than alternative high-quality algorithms.
+In the topic analysis task, LOTUS efficiently processes hundreds of ArXiv papers to discover a taxonomy of key topics in dozens of seconds, while once again providing configurable speedups with statistical accuracy guarantees.
+Overall, our main contributions are the following:
+We propose semantic operators, the first formalized programming model for general-purpose AI-based transformations with natural language specifications.
+We define an optimization framework for semantic operators with statistical accuracy guarantees, and novel optimizations for several operators, including semantic filters, joins, top-k and group-by, yielding up to speedups.
+We extensively evaluate the expressiveness of the semantic operator model, the quality of our gold algorithms, and the benefit of our optimizations, including comparisons to DocETL, UQE, and AI UDFs.
+2. The Semantic Operator Model
+Semantic operators provide a declarative interface for semantic-based manipulation and access to data.
+Extending the data independence model of relational systems (Codd, 1970), these new AI-based operators introduce model-data independence, which separates the application logic from the underlying ML-based algorithm and access pattern that specifies how data is passed to model invocations.
+Unlike relational operators, semantic operators are parametrized by natural language expressions and rely on AI-based computation, making their behavior inherently ambiguous and imprecise.
+We address this key challenge by presenting a formalism for defining semantic operators, their behavior and correct optimizations, in this section.
+In Section 3, we build on this formalism to provide optimized execution plans with accuracy guarantees, similar to relational query optimizers.
+2.1. Defining Semantic Operators
+Definition.
+A semantic operator is a declarative transformation over one or more datasets, parameterized by a natural language expression.
+Each semantic operator can be implemented by many AI-based algorithms, and its behavior is defined with respect to a gold algorithm.
+Table 1 lists a core set of semantic operators, which cover common semantic transformations in real-world applications and mirror key transformations in relational operators.
+Specific systems may, of course, provide additional semantic operators beyond the ones we discuss here.
+Each semantic operator takes a parameterized natural language expressions (langex for short), which are natural language expressions that specify a function over one or more attributes.
+As Figure 2 demonstrates, the langex signature varies for different semantic transformations.
+The figure shows a LOTUS program,
+which performs a sem_join, sem_map and sem_agg over datasets of research papers and user projects.
+While the sem_join langex signature provides a natural language predicate, the sem_map langex signature indicates a natural language projection, and the sem_agg langex is a commutative, associative aggregator expression, which here indicates a summarization task over abstracts.
+We provide a high-level definition of each semantic transformation with respect to the user langex and a world model111In practice, the world model, may be the strongest LLM a practioner has available., , which captures a probability distribution over the vocabulary .
+For example, semantic filter returns , where is the input relation and represents a natural language predicate evaluated on tuple with model M.
+Notably, the definition of each semantic operator can be implemented by multiple AI-based algorithms, and different decisions as to how to invoke the model over the relation have consequences on the algorithm’s result quality.
+Thus, to fully specify the ground truth behavior of each semantic operator, we define a gold algorithm, which is a computable and tractable ML algorithm that produces high-quality results and specifies the model’s access pattern over the data, .
+2.2. Defining Correct Optimizations for Semantic Operators
+Semantic operators create a rich design space of diverse execution plans.
+While gold algorithms define a tractable, high-quality implementation, they are often expensive, with the complexity of LLM calls scaling linearly or quadratically in the dataset cardinality.
+For example, the semantic filter’s gold algorithm entails invoking the LLM in a linear pass over each tuple.
+However, an alternate AI-based algorithm may reduce cost, while offering close results.
+In the case of semantic filters, close results, can be measured according to accuracy metrics, in this case recall and precision, and guaranteed with high probability compared to the accuracy obtained by the gold algorithm.
+As such, we define correct optimizations for semantic operators222Our optimization approach can naturally extend to multi-operator execution plans by composition.
+In multi-operator programs, the user may configure an accuracy target and failure probability per operator, or for the full program, in which case the optimizer assigns a target accuracy and failure budget to each operator.
+below.
+Definition.
+A correct optimization for a given semantic operator, and a gold algorithm for that operator, reduces cost while providing statistical accuracy guarantees with respect to the gold algorithm.
+Specifically, the optimization should ensure an accuracy target, , is met with probability .
+In general, optimized semantic operator plans allows for both lossless optimizations and approximations with respect to a gold algorithm.
+This formulation builds on approaches in approximate query processing and assumes some error is often tolerable, which we find (Section 5) is reasonable for achieving high-quality results, for semantic processing, which is inherently non-exact.
+2.3. Core Semantic Operators
+We now overview several core semantic operators, providing the operator’s definition and at least one gold algorithm for each.
+We discuss design decisions for our gold algorithms based on state-of-the-art algorithms studied in the AI literature, well-known failure cases, such as long-context challenges (Liu et al., 2023), and our empirical evaluation in Section 5.
+Semantic Filter is a unary operator over the relation and returns the relation , where the langex provides a natural language predicate over one or more attributes.
+Gold Algorithm.
+Our gold algorithm runs batched LLM calls over all tuples in relation .
+Each model invocation, , prompts the LLM with a single tuple , the langex predicate, and an operator-specific instruction to generate a boolean value.
+This simple choice avoids well-studied long-context issues (Liu et al., 2023) by processing rows independently rather than batching multiple tuples within a single prompt invocation.
+Semantic Join provides a binary operator over relations and to return the relation .
 Here the langex is parameterized by the left and right join keys and describes a natural language predicate over both.
-Semantic Top-k imposes a ranking111This definition implies that imposes a total and consistent ordering.
+Gold Algorithm.
+The gold algorithm implements a nested-loop join pattern, performing a single predicate evaluation over a tuple pair with each model invocation , similar to filters This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively.
+This quadratic join algorithm is suitable for small join tables but scales poorly for large tables.
+Semanic Top-k imposes a ranking333This definition implies that imposes a total and consistent ordering.
 However, this definition can also be softened to assume partial orderings and noisy comparisons with respect to model .
-over the relation and returns .
-Here the signature of the langex provides a general ranking criteria according to one or more attributes of the input relation.
-The underlying system can use this langex to impose a ranking over any subset of rows, according to the chosen implementation.
-Semantic Aggregation performs an aggregation over the input relation, returning .
-Here, the the langex signature provides a commutative and associative aggregation function, which can be applied over any subset of rows to produce an intermediate results.
-We note that the langex with respect to model assumes infinite context and can take an arbitrary number of input tuples.
+over the relation and returns the ordered sequence, .
+Here, the langex signature provides a general ranking criteria according to one or more attributes.
+Gold Algorithm.
+Two important algorithmic design decisions for the semantic top-k include how to implement LLM-based comparison and how to aggregate ranking information from these comparisons.
+Our gold algorithm uses pairwise LLM comparisons for the former, and a quick-select top-k algorithm (Hoare, 1961) for the latter.
+We briefly describe the reason for these choices and alternatives considered.
+Our decisions build on prior works, which have studied LLM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n.d.]) with the goal of achieving high quality results in a modest complexity of LLM calls or comparisons.
+First, pairwise-prompting methods offer a simple and high-quality approach that feeds a single pair of tuples to each LLM invocation, , prompting the model to compare the two inputs and output a binary label.
+The two main classes of alternatives are point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), and list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), both of which have been shown to face quality issues (Sun et al., 2023; Qin et al., 2024; Desai and Durrett, 2020).
+We verify these limitations in Section 5.
+In contrast, pairwise comparisons have been shown to be effective and relatively robust to input ordering (Qin et al., 2024).
+In addition, we consider several possible rank-aggregation algorithms, including quadratic sorting algorithms, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm.
+All three alternatives represent candidate gold algorithms.
+Our evaluation in Section 5 demonstrates that each of these sorting algorithms yield high-quality results, comparable to one another.
+However, the quick-select-based algorithm offers an efficient implementation with at least an order of magnitude fewer LLM calls than the quadratic sorting algorithm and more opportunities for efficient batched inference, leading to lower execution time, compared to a heap-based implementation.
+The quick-select top-k algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining tuples to the pivot tuple to determine the rank of the pivot.
+Because each round is fully parallelizable, we can efficiently batch these LLM-based comparisons before recursing in the next round.
+Semantic Aggregation performs a many-to-one reduce over the input relation, returning .
+Here, the the langex signature provides a commutative and associative aggregation function444We note that the ordering of inputs within an LLM prompt invocation can in fact affect results quality for some tasks.
+Thus, programmers may wish to override the commutativity and associativity assumptions of the semantic aggregation langex.
+For this, the LOTUS API exposes a partioner function which groups and orders inputs within LLM prompts according to an arbitrary user-specification., which can be applied over any subset of rows to produce an intermediate results.
+We note that the langex iteself is model-agnostic, assuming infinite context.
 Managing finite context limits of the underlying model is an implementation detail left to the system.
-Semantic aggregations can be useful for many tasks involving many-to-one reduction patterns, such as summarization or question-answering over multiple documents.
-Semantic Group-by takes a langex that specifies a grouping criteria and a target number of groups.
-As a example extending that of Figure 2, we might group-by the topics presented in each paper abstract with a target of 10 group.
-Here, the group-by operator must discover representative group labels and assign each tuple a label.
-In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n.
-d.]).
+Gold Algorithm.
+Our gold algorithm must efficiently orchestrate the LLM to manage long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length.
+We select a hierarchical reduce pattern for the gold algorithm.
+Our choice builds on the LLM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n.d.]; lan, [n.d.]), which we briefly overview.
+Prior works primarily implement one of two aggregation patterns: either a fold pattern, which performs a sequential, linear pass over the data, while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains.
+Both represent candidate gold algorithms, however, the hierarchical pattern has been shown to produce higher quality results for commutative, associative aggregation tasks, like summarization, in prior work (Chang et al., 2024) and allows for greater parallelism during query processing, making it our default choice.
+Semantic Group-by takes a langex that specifies a projection from a tuple to an unknown group label, as well as a target number of groups, which specifies the desired granularity of group labels.
+As a example, a user might group-by the topics presented in a set of ArXiv papers, wishing to find 10 key groups.
+The group-by operator must discover representative group labels and assign a label to each each tuple.
+In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n.d.]).
 Clustering algorithms over points in a metric space typically optimize the potential function tractably using coordinate descent algorithms, such as k-means.
-In the semantic group-by, the clustering task is over unstructured documents with a natural language similarity function specified by the user langex , which imposes a real-valued score between a tuple and a candidate label , and the operator poses the following optimization problem:
+For the semantic group-by, the clustering task is over unstructured fields with a natural language similarity function specified by , which imposes a real-valued score between a tuple and a candidate label .
+This operator poses the following optimization problem:
 where are group labels, each consisting of tokens in the vocabulary, .
-3. The LOTUS Programming Interface
-In this section we describe the LOTUS API, which implements the semantic operator model as an extension of Pandas (pan, [n.
-d.]).
+Gold Algorithm.
+Since this operator, by definition, entails a clustering problem, known to be NP-hard (Dasgupta, [n.d.]), our gold algorithm instead offers a computable implementation to obtain group labels using a tractable clustering algorithm to discover group labels followed by a point-wise classification step.
+Both steps use a linear LLM pass over the relation.
+Specifically, the first stage performs an LLM-based clustering algorithm to discover centers , given user langex.
+Our gold algorithm first performs a semantic projection, prompting the LLM to predict a label for each input tuple.
+Then, we embed these candidate labels and perform an efficient vector clustering using a k-means algorithm to construct groups.
+For each group, we top-k sample tuples based on centroid-similarity scores, and use a semantic aggregation that prompts the model to generate an appropriate label over each group.
+In the second stage of our gold algorithm, we use the generated labels, and invoke the model to perform point-wise assignments .
+3. Optimized Execution Plans for Semantic Operators
+In this section, we present novel optimizations for several costly operators, including semantic filters, joins, top-k and group-by.
+Each optimization provides statistical accuracy guarantees with respect to the gold algorithm, building from Section 2.
+While this paper focuses on novel optimizations for several expensive semantic operators, we envision a rich space of future work exploring new optimizations and applications of traditional optimizations to semantic operator programs.
+For example, several prior works demonstrate performance gains in both logical query plan optimizations (e.g.
+operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g.
+code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)).
+3.1. Optimizing Semantic Filter
+We provide an approximation for semantic filters that obtains recall and precision targets and with respect to the gold implementation with probability .
+To achieve reduced cost we leverage the cheaper, but less accurate proxy model , which outputs a score indicating whether the tuple passes the predicate.
+This idea is inspired by prior works (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022; Yue et al., 2024), which leverage model cascades for different problem settings.
+Specifically, several works study cascades in the video analytics setting with vision models, which have significantly different properties than LLMs, requiring different proxy scoring mechanisms.
+Other works study cascades for LLMs, but use heavy-weight scoring mechanisms to decide whether to use the proxy model and do not provide accuracy guarantees.
+In contrast, we focus on providing accuracy guarantees for the constrained problem of applying cascades for filters, which allows us to use lighter-weight scoring functions than the more general case studied in prior works.
+In general, we cannot assume the proxy model is accurate.
+In fact, the proxy model may perform very poorly for some tasks and records.
+The goal of our algorithm is to automatically discover the quality of the proxy at query time by sampling and comparing to the gold algorithm with the oracle model.
+Our algorithm will exploit the proxy when it is likely to provide high-quality outputs, and otherwise defer the oracle model.
+To do this we must learn a decision rule, with learned thresholds on the proxy scores using sampling.
+As prior work (Kang et al., 2020) discusses, such sampling-based algorithms introduce multiple-hypothesis testing problems, requiring statistical corrections using confidence intervals, which we carefully apply in our algorithm.
+We consider a small LLM as the proxy model and generate scores using log-probabilities555This optimizations assumes access to log-probabilities, which is available in common model providers, such as OpenAI, and serving systems, such as vLLM.
+In the absence of available log-probabilities, alternative uncertainty quantification methods, such as prompt-based methods, may be more suitable.
+corresponding to the True or False output tokens of the proxy model’s predicate evaluations, re-scaling by the quantiles over all generated log-probabilities over the relation.
+Algorithm 1 shows the full procedure for performing the approximate semantic filter.
+We begin by collecting a sample of tuples , and labeling each sample with the oracle and proxy models.
+The sampling procedure uses importance sampling and defensively mixes a uniform sample following prior work (Kang et al., 2022).
+Using the central limit theorem and the normal approximation on the distribution of sample statistics, we then learn a decision rule by finding thresholds and , which will respectively ensure the precision target and recall target are met, each with an error probability of .
+The statistical sub-procedures to choose each threshold follows prior work (Kang et al., 2022), but applies them to this new setting where we must ensure both targets are met, requiring a correction for hypothesis testing and multiple failure modes.
+Finally, using the learned decision rule, our algorithm proceeds to process each tuple in the relation.
+If the tuple’s proxy score is at least , we mark it as passing the predicate.
+If the tuple’s proxy score is less than or equal to , we mark it as failing the predicate, and for tuples with proxy scores between and , we resort to the oracle model to provide a label.
+3.2. Optimizing Semantic Join
+Similar to semantic filters, we provide an approximation for semantic joins that obtains recall and precision targets and with respect to the gold algorithm with probability .
+Due to the expensive quadratic scaling of the nested-loop join, rather than using a small LLM as the proxy model, we consider an even cheaper proxy based on semantic similarity scores using embeddings.
+We leverage two possible proxy algorithms and dynamically choose the lowest cost one.
+The first approximation, sim-filter produces a proxy score , based on embedding similarity.
+We perform batched similarity search between the right and left join keys and re-calibrate similarity scores according to their quantiles to obtain proxy scores.
+This approximation is likely to perform efficiently when tuple pairs with high semantic similarity between the the right and left join key are more likely to pass the predicate.
+This correlation phenomenon between predicate matches and embedding similarity has been studied by prior works (Patel et al., 2024b) and is not always present.
+Thus, we introduce an alternative approximation, often suitable when correlation is not present.
+The second approximation, project-sim-filter first performs a projection over the left join key and then uses the projected column to compute proxy scores uses the embedding similarity between the projected value and .
+As before, we also re-calibrate the proxy scores taking their quantiles.
+Intuitively, performing the projection is useful when tuple pairs that pass the user’s natural language predicate exhibit low semantic similarity.
+The projection step invokes the LLM over each tuple in left join table, prompting it to predict values for the right join key.
+Notably, this LLM projection is ungrounded, i.e., it is done without knowledge of the right join key’s attribute domain and can thus be performed in a fully parallelized semantic map operation.
+Figure 3 provides an example of a semantic join between a table of papers and datasets, where the predicate evaluates whether a given paper abstract uses a specific dataset.
+Here, the map step would invoke the LLM over each abstract, instructing the model to output the dataset used, conditioned only on the abstract.
+To dynamically choose between these two approximations, we follow a procedure similar to Algorithm 1 used for semantic filters.
+We begin by importance sampling to collect the sample and construct the set of oracle labels, , over the sample.
+We then obtain tresholds and independently for each approximation algorithm using their respective proxy scores, and .
+We use the learned thresholds for each candidate plan to then determine the exact oracle cost needed to execute either algorithm, and we take the least cost plan with its associated learned thresholds to proceed to evaluate the predicate for each tuple pair.
+3.3. Optimizing Semantic Group-by
+We provide an efficient approximation that guarantees a classification accuracy target, , is met with probability .
+We follow the clustering algorithm in the first stage of the gold algorithm to discover centers .
+We then use a proxy-based approximation for point wise assignments in the second stage of the algorithm.
+Similar, to semantic joins, we consider semantic similarity scores as a cheap proxy, although other proxy models are also feasible.
+Specifically, we leverage the embeddings constructed during the clustering stage and compute similarity scores between the the candidate label of each tuple and the discovered centers .
+The proxy score of the tuple with a predicted label, , for a discovered center is given by .
+Our goal is then to learn a threshold, , such that if the proxy score between a tuple and center is greater than , we return the center as label for the tuple, and otherwise we resort to the more expensive LLM-based classification procedure.
+We accomplish this by uniform sampling and running the sub-procedure equivalent to the used in Algorithm 1 for semantic filters, where the metric we evaluate is classification accuracy over the groups.
+3.4. Optimizing Semantic Top-k
+We leverage the embedding similarity scores to optimize pivot selection for some queries, while incurring no accuracy loss.
+This optimization is useful when there exists correlation between the ranking imposed by the user’s arbitrary sorting criteria and the ranking imposed by semantic similarity scores.
+In this case, we can sort tuples based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot.
+This can reduce the number of LLM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss.
+In the case of no correlation between the langex-based ranking and similarity-based ranking, this method amounts to random pivot selection, and in the worst case of an adversarial pivot, the algorithm will incur one extra pivot round, which can impact execution time, but not degrade quality.
+4. The LOTUS Systems
+In this section we describe the LOTUS system, which implements the semantic operator model as an extension of Pandas (pan, [n.d.]).
 We chose a Pandas-like API in our initial system implementation to make it easy for users to integrate LOTUS with popular AI libraries in Python.
 However, semantic operators could also be added to a variety of other data processing APIs and query languages, such as SQL.
+LOTUS leverages vLLM (Kwon et al., 2023) to perform efficient batched inference, and uses FAISS (Douze et al., 2024; Johnson et al., 2017) to support efficient vector search for LOTUS’ semantic search, similarity join and cluster operations, with indices stored and maintained locally on disk, by default.
 4.1. Datatypes
-3.1.1. Data Model
-LOTUS’ data model consists of tables with structured and unstructured text fields.
-LOTUS’ semantic operators can take both of these data-types as inputs.
-For instance, Figure 2 shows an example LOTUS program, where the table schema over ArXiv papers consists of attributes for the paper’s title, ArXiv URL, abstract, ArXiv domain categories, and the publication date.
-These columns are then passed as parameters to semantic operators, such as sem_search, sem_filter, sem_agg.
+LOTUS’ data model consists of tables with structured and unstructured fields (e.g., text or images). LOTUS’ semantic operators can take both of these data-types as parameters to a langex.
 Additionally, LOTUS supports semantic indices over natural-language text columns to provide optimized query processing.
-These indices leverage semantic embeddings over each document in the column to capture semantic similarity using embedding distance metrics.
-Semantic indices can be created off-line with sem_index and a specified retriever model, and then loaded using load_sem_index.
-3.1.2. Parameterized Natural Language Expressions (langex)
-Programmers write each langex in natural language, parameterized by one or more data columns, which are indicated in the brackets within a string, or double brackets within a formatted string, as Figure 2 shows.
-Notably, these language expressions are sufficiently versatile and easy to program with, providing an intuitive and higher-level interface to the user.
-Operator-specific LLM prompts based on these expressions are automatically constructed by LOTUS’ underlying query engine, and the underlying system can leverage existing prompt optimization techniques (Khattab et al., 2023; Yuksekgonul et al., 2024) on- or off-line.
+These indices leverage semantic embeddings over each document in the column and capture semantic similarity using embedding distance metrics.
+Semantic indices can be created off-line with sem_index and a specified retriever model, and then loaded from disk using load_sem_index.
 4.2. Semantic Operators in LOTUS
 We now overview the semantic operators supported in the LOTUS API, which includes the core set of operators described in Section 2, as well as several additional variants provided for convenience.
+Each operator takes optional parameters to specify a target accuracy and error probability, which the optimizer will use to transparently perform optimizations.
 Sem_filter, Sem_join, & Sem_sim_join.
@@
 In addition, LOTUS provides a join variant, sem_sim_join, where tuples are matched according to their semantic similarity, rather than an arbitrary natural-language predicate.
-Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join, which indicates additional optimization opportunities to the query engine.
+Akin to an equi-join in standard relational algebra, the semantic similarity join is a specialized semantic join, which indicates additional optimization opportunities to the query engine by leveraging vector similarity search.
 Figure 4 provides an example of the sem_join compared to the sem_sim_join, where the user specifies the left and right table join keys, and a parameter .
@@
 The semantic search in this case will first find the top- most relevant documents and then re-rank the top- found documents to return the top n_rerank.
-Sem_agg performs an aggregation over the input relation, with as langex signature that provides a commutative and associative aggregation function, as shown by Figure 2.
+Sem_agg performs an aggregation over the input relation, with a langex signature that provides a commutative and associative aggregation function, as shown by Figure 2.
 Similar to sem_topk, LOTUS allows users to specify a group-by parameter.
-Sem_group_by creates groups over the input dataframe according to the langex grouping criteria by clustering the data into a target number of groups (e.g., papers_df.sem_group_by("the topic of the abstract", 20)).
-By default, the operator automatically selects group labels, but the user can also optionally specify target labels.
+Sem_group_by creates groups over the input dataframe according to the langex grouping criteria and target number of groups.
+By default, the operator discovers suitable group labels, but the user can also optionally specify target labels.
 This operator is useful both for unsupervised discovery of semantic groups, and for semantic classification tasks.
@@
 While sem_map projects to an arbitrary text attribute, sem_extract projects each tuple to a list of sub-strings from the source text.
-This is useful for applications, such as entity extraction or fact-checking, where finding snippets or verified quotes may be preferable to synthesized answers.
-4. Execution Plans and Optimizations for Semantic Operators
-Semantic operators create a rich design space of diverse execution plans which have significant consequences on system efficiency and accuracy.
-For example, there are many possible algorithms to join two collections of records using a LLM and a natural language expression (all the standard join algorithms, as well as ML heuristics to approximate the join, e.g., by semantically clustering similar records using an embedding model).
-Specifically our goal is to offer an efficient plan while while guaranteeing high accuracy.
-Defining accuracy for semantic transformations, however, presents a fundamental challenge in the absence of ground truth labels.
-Therefore, we define accuracy of each operator plan with respect to a ”gold” algorithm, which uses a powerful oracle model .
-These ”gold” algorithms are tractable solutions that can achieve high quality results, but may be expensive in terms of execution time and LM call complexity.
-We discuss the design space of these ”gold” algorithms for several core semantic operators we introduced.
-In addition, we present several novel optimizations for semantic filters, joins, top-k and group-by.
-These optimizations are either lossless or provide statistical guarantees on accuracy with respect to a ”gold” algorithm with the expensive model .
-Specifically, our approximations emulate the results of the expensive but high-quality algorithm by using model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) or semantic indexes, building on past work on accelerating ML-based query processing with guarantees (Kang et al., 2022, 2019) but achieving this property for new types of operators.
-Notably, these optimizations can all be performed declaratively given a user’s accuracy targets.
-We show in Section 5 that they significantly accelerate real applications.
-While this paper proposes new optimizations for only a few, expensive semantic operators, we envision that a wide range of other new and traditional optimizations are possible under the semantic operator model.
-For example, several prior works demonstrate performance gains in both logical query plan optimizations (e.g.
-operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g.
-code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)).
-4.1. sem_filter
-4.1.1. Gold Algorithm
-Our gold algorithm for the semantic filter runs batched LLM calls over the tuples in relation .
-Each tuple, , is processed independently using the model , which is prompted to output a boolean value indicating whether the tuple passes the langex predicate.
-4.1.2. Algorithmic Approximation
-We can declaratively optimize semantic filters by providing an approximation which will obtain recall and precision targets and with respect to the gold implementation with probability .
-To achieve reduced cost we leverage the cheaper, but less accurate proxy model , which provides a score indicating whether the tuple passes the predicate, in conjunctions with the oracle LLM .
-This idea is inspired by prior works (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022; Yue et al., 2024), which leverage model cascades for different problem settings.
-Specifically, several works study cascades in the video analytics setting with vision models, which have significantly different properties than LLMs, requiring different proxy scoring mechanisms.
-Additionally, other works, which study cascades for LLMs, use heavy-weight scoring mechanisms to decide whether to use the proxy model, which incurs a heavy execution time penalty, and does not provide accuracy guarantees.
-In contrast, we focus on providing accuracy guarantees for the constrained problem of applying cascades for filters, which allows us to use lighter-weight scoring functions than the more general case studied in prior works.
-Providing accuracy guarantees is important because, in general, we cannot assume the proxy model is accurate.
-In fact, the proxy model, may perform very poorly for some tasks and datasets.
-The goal of our algorithm is to automatically discover the quality of the proxy by sampling and comparing to the gold algorithm with the oracle model.
-If the proxy provides high quality results, our algorithm will exploit it, leading to lower execution time.
-However, if the proxy model performs poorly, we will selectively route to it according to the proxy scores , and a learned threshold on the proxy scores.
-As prior work (Kang et al., 2020) discusses, such sampling-based algorithms to learn a threshold, , introduce multiple-hypothesis testing problems, requiring statistical corrections using confidence intervals, which we carefully apply in our algorithm.
-Our algorithm leverages a small LLM as the proxy model.
-We generate scores using the the log-probabilities corresponding to the True or False output tokens of the proxy model’s answer and re-scaling by the quantiles over the proxy scores of tuples in the relation.
-Algorithm 1 shows the full procedure for performing the approximate semantic filter.
-We begin by using an collecting a sample , which we label with the oracle and proxy models.
-The sampling procedure uses importance sampling and defensively mixes a uniform sample following prior work (Kang et al., 2022).
-Using a the central limit theorem and the normal approximation on the distribution of sample statistics, we then find thresholds and , which will respectively ensure the desired precision target and recall target are met, each with an error probability of .
-These sub-procedures follow the statistical techniques used in prior work (Kang et al., 2022), but applies them to this new setting where we must ensure both targets are met, requiring a correction for hypothesis testing and multiple failure modes.
-Finally the algorithm proceeds to process each tuple in the relation.
-If the tuple’s proxy score is at least , we mark it as passing the predicate.
-If the tuple’s proxy score is less than or equal to , we mark it as failing the predicate, and for tuples with proxy scores between and , we resort to the oracle model to provide a label.
-4.2. sem_join
-4.2.1. Gold Algorithm
-Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left table, and right table, .
-The gold algorithm implements a nested-loop join pattern with efficient batched inference to maximize GPU utilization.
-This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively.
-Each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate.
-This quadratic join algorithm is suitable for small join tables but scales poorly for large tables.
-4.2.2. Algorithmic Approximation
-Similar to semantic filters, we declaratively optimize semantic joins by providing an approximation that obtains recall and precision targets and with respect to the gold algorithm with probability .
-Due to the expensive quadratic scaling of the nested-loop join, rather than using a small LLM as the proxy model, we leverage an even cheaper proxy based on semantic similarity scores, which can be obtained efficiently with a semantic index over the join keys.
-We leverage two possible join approximation algorithms and dynamically chooses the lowest cost one, which will minimize the number of LLM calls.
-The first approximation, sim-filter produces a proxy score , based on embedding similarity.
-We leverage the semantic index to perform batched similarity search and re-calibrate similarity scores according to the quantiles of the distribution over all pairs to obtain proxy scores.
-This approximation is likely to perform efficiently when tuple pairs with high semantic similarity between the the right and left join key are more likely to pass the predicate.
-This correlation phenomenon between predicate matches and embedding similarity has been studied by prior works (Patel et al., 2024) and is not always present.
-Thus, we introduce an alternative approximation suitable when correlation is not present.
-The second approximation, map-sim-filter first performs a projection over the left join key and then uses the projected column to compute proxy scores uses the embedding similarity between the projected value and .
-As before we also re-calibrate the proxy scores taking their quantiles.
-Intuitively, performing the projection is useful when tuple pairs that pass the user’s natural language predicate exhibit low semantic similarity.
-The projection step invokes the LM over each tuple in left join table, prompting it to predict values in the attribute domain of the right join key that match.
-Notably, this LM projection is ungrounded, i.e., it is done without knowledge of the right join key’s attribute domain and can thus be performed in a fully parallelized semantic map operation, which can optionally take user-provided demonstrations.
-Figure 5 provides an example of a semantic join between a table of papers and datasets, where the predicate evaluates whether a given paper abstract uses a specific dataset.
-Here, the map step would invoke the LM over each abstract, instructing the model to output the dataset used, conditioned only on the abstract.
-To dynamically choose between these two approximations, we follow a procedure similar to Algorithm 1 used for semantic filters.
-We begin by importance sampling to collect the sample and construct the set of oracle labels, , over the sample.
-We then obtain , and independently for each approximation algorithm using their respective proxy scores.
-Using the tuned thresholds and known proxy scores over for both algorithms over the dataset, we then determine the exact oracle cost needed to execute either algorithm, and we take the least cost plan with its associated learned thresholds to proceed to evaluate the predicate for each tuple pair.
-4.3. sem_group_by
-Performing this semantic operator entails discovering group labels and classifying each input tuple using the discovered labels.
-In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n.
-d.]).
-Our gold algorithm instead offers a tractable implementation to obtain group labels using a linear pass over the data with the reference model .
-Furthermore, once the labels are discovered, we perform the LM-based classification task and show how to optimize this step using an approximation with probabilistic guarantees.
-4.3.1. Gold Algorithm
-Our semantic group-by algorithm proceeds in two stages.
-The first stage performs the unsupervised label discovery, given the user’s grouping criteria and target number of groups.
-This stage first performs a semantic projection, prompting the LM to predict a label for each input tuple.
-Then, we create a semantic index over the projected column and perform a vector clustering over these predicted labels, using FAISS’ efficient k-means implementation to construct groups.
-Lastly, for each group, we top-k sample documents based on their similarity to the centroid, and few-shot prompt the LM to generate a label describing the given documents according to the user-provided grouping criteria (e.g.
-the topics of the documents) using a semantic aggregation.
-In the second stage of the algorithm, we use the generated labels to perform classification over each document.
-We do this using a semantic map, which prompts the LM to output a label conditioned on the candidate labels and document.
-4.3.2. Algorithmic Approximation
-We provide an efficient approximation, specifically for the classification task in the second stage of the algorithm.
-Similar to semantic filters and join, we perform this optimization declaratively, while guaranteeing a classification accuracy target is met with probability .
-Similar, to semantic joins, we leverage semantic similarity scores as a cheap proxy.
-Specifically, we leverage the index constructed in stage one of the algorithm and compute similarity scores between the embeddings of each document’s un-grounded prediction and the embeddings of each class label.
-The proxy score of the tuple with a predicted label, , for group label is given by .
-Our goal is then to learn a threshold, , such that if the similarity between an embedding and the class label’s embedding is greater than , we return the class label, and otherwise we resort to the more expensive LM-based classification procedure.
-We accomplish this by uniform sampling and running the sub procedure equivalent to the used in Algorithm 1 for semantic filters, where the metric we evaluate is classification accuracy over the groups.
-4.4. sem_topk
-Performing a semantic top- ranking requires logically reasoning across rows, entailing joint reasoning over often large amounts of data.
-While prior works have studied LM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n.
-d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, our implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low execution time.
-Here we outline leverage several algorithmic design decisions for the gold algorithm as well as a lossless optimization.
-4.4.1. Gold Algorithm
-Two important algorithmic design decisions for the semantic top-k include how to implement the LM-based comparison and how to aggregate ranking information across LM-based comparisons.
-Our implementation uses pairwise LM-based comparisons for the former, and a quick-select based top-k algorithm (Hoare, 1961) for the latter.
-We briefly describe the reason for these choices and alternatives considered.
-Pairwise-prompting methods offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label.
-The two main classes of alternatives are point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), and list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), both of which have been shown to face quality issues (Sun et al., 2023; Qin et al., 2024; Desai and Durrett, 2020).
-In contrast, pairwise comparisons have been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
-In addition, we studied several alternative rank-aggregation algorithm, including quadratic sorting algorithms, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm, which we discuss in Section 5.
-The quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot.
-Because each round is fully parallelizable, we can efficiently batch these LM-based comparisons before recursing in the next round.
-In section 5, we find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for efficient batched inference, leading to lower execution time, compared to a heap-based implementation.
-4.4.2. Optimizations
-We leverage the semantic index to optimize pivot selection for some queries, while incurring no accuracy loss.
-This optimization is useful when there exists correlation between the rankings imposed by the user’s arbitrary sorting criteria and the rankings imposed by semantic similarity scores.
-In this case, LOTUS can sort the document set based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot.
-This can reduce the number of LM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss.
-In the case of no correlation between the langex-based ranking and similarity-based ranking, this method amounts to random pivot selection, and in the worst case of an adversarial pivot, the algorithm will incur one extra pivot round, which can impact execution time, but not degrade quality.
-4.5. sem_agg
-Performing semantic aggregations is inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows.
-Thus, the operator’s implementation must efficiently orchestrate the LM over large amounts of data, while managing long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length.
-LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results.
-4.5.1. Gold Algorithm
-Our implementation builds on the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n.
-d.]; lan, [n.
-d.]).
-These implementations primarily leverage one of two aggregation patterns: either a fold pattern, which produces a sequential, linear pass over the data while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains.
-Our aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024).
+This is useful for applications, such as entity extraction, where finding snippets or verified quotes may be preferable to synthesized answers.
 5. Evaluation
-We now evaluate the expressiveness of the semantic operator model as well as the efficiency of our algorithms and optimizations through four diverse applications: fact-checking (Section 5.1), extreme multi-label classification (Section 5.2), search and ranking (Section 5.3), and paper analysis (Section 5.4).
-We find that the semantic operator model is expressive and easy to program with, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality.
-We also find that our optimizations that take advantage of the declarative nature of semantic operators can accelerate semantic filtering, top-k, group-by and join operators by up to while offering statistical accuracy guarantees.
-Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while executing up to 28 faster end-to-end.
-LOTUS’ current implementation extends the Pandas API (pan, [n.
-d.]).
-We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) to support efficient vector search for LOTUS’ semantic search, similarity join and cluster operations.
-Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama-3 models (int, [n.
-d.]), with a batch size of 64 running on vLLM (Kwon et al., 2023).
-For our experiments that use OpenAI’s GPT-4o-2024-05-13 model (ope, [n.
-d.]), we control for cost by limiting execution to 64-way thread parallelism.
-We set temperature to , unless otherwise stated.
-5.1. Application: Fact-Checking
-Application.
-We evaluate on FEVER (Thorne et al., 2018), a claim verification dataset.
-We use the development dataset, which contains about 38,000 total claims, of which we sample 1,000 for our evaluation.
-Each claim is labeled with one of three labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”, and the task is to correctly determine the label of each claim, leveraging evidence from a corpus of million Wikipedia articles.
+We systematically evaluate LOTUS on four bulk-semantic processing applications from the AI and data literature: fact-checking, biomedical multi-label classification, search, and topic analysis.
+Some of these applications can be expressed by recent LLM-based analytics systems, like AI UDFs, UQE and DocETL, allowing us to compare against them; while other applications are not supported by the baseline systems, but LOTUS supports them.
+For each application, we additionally compare against strong hand-designed pipelines from the literature.
+We evaluate the LOTUS optimizer, comparing the performance of optimizations for semantic filters, joins, top-k and group-by with our gold algorithms.
+Lastly, we analyze the statistical accuracy guarantees provided by LOTUS’ optimizations.
+Overall, we find the following:
+Compared to AI-based analytics systems, like AI UDFs, UQE, and DocETL, LOTUS attains similar or up to higher accuracy, while running up to faster and offering statistical accuracy guarantees that the other systems do not provide.
+Compared to hand-written pipelines from the AI literature for fact-checking, biomedical classification, and search, LOTUS matches state-of-the-art quality, or exceeds it by up to , in programs involving a few semantic operators.
+The LOTUS optimizer can substantially reduce cost of semantic operator programs, by up to compared to high-quality gold algorithms, and provides accuracy guarantees, which hold across repeated trials.
+To report controlled latency numbers, we run each baseline and experiment locally on 4 80GB A100 GPUs using Llama-3-70B (noa, [n.d.]c) and E5 embeddings (Wang et al., 2024), with a batch size of 64 running on vLLM (Kwon et al., 2023), unless otherwise stated.
+To run DocETL baselines and reproduce some of their benchmarks (Shankar et al., 2024), we additionally use OpenAI (noa, [n.d.]e) models, including GPT-4o-mini-2024-07-18, GPT-4o-2024-08-06 model, and text-embedding-3-small, where noted.
+For our experiments that use OpenAI models , we control for cost by limiting execution to 64-way thread parallelism.
+For reproducibility, we set temperature to for all methods and baselines, unless otherwise stated.
+5.0.1. Benchmarked Methods
+We briefly overview the main methods we benchmark and tested parameters.
+AI UDFs.
+Many database vendors (ver, [n.d.]; dat, [n.d.]; noa, 2023; MotherDuck, [n.d.]; WilliamDAssafMSFT, 2024; noa, [n.d.]d) now support AI UDFs, which include map-like, row-wise LLM operations and primitives for vector search.
+To control for serving infrastructure, we implement these programs using LOTUS running on vLLM, and restrict our API to sem_map, sem_search and sem_sim_join.
+UQE.
+UQE (Dai et al., 2024) proposes an optimized LLM-powered filter method, which is relevant to this work.
+UQE also optimizes non-LLM aggregates (e.g.
+COUNT) integrated with LLM-based queries, but this is beyond the scope of our evaluation.
+Since the code is not open-source, we implement the UQE filter in about 100 line of code in python.
+The UQE filter method takes an LLM call budget and provides a best-effort embedding-based approximation.
+In contrast, LOTUS provides a general proxy-oracle approximation method, where the proxy may be embedding-based or LLM-based.
+To provide a fair comparison between UQE and LOTUS, we normalize the latency budget when comparing against LOTUS programs that use LLM-based proxies, and we normalize the LLM call budget when comparing against LOTUS programs that use embedding-based proxies.
+We also hyper-parameter tune the mini-batch size used for UQE’s active learning algorithm.
+DocETL.
+DocETL (Shankar et al., 2024) uses an LLM agent as the query optimizer to perform rewrites to programs, allowing programs to specify separate LLMs for the optimizer and execution.
+Our experiments find that the optimizer often fails, and among runs where the optimizer succeeds, performance varies.
+Based on personal communication with the authors, we run the DocETL baselines multiple times.
+We obtain three successful runs and report the average performance of these three successes.
+We also try using both Llama-70B as the DocETL optimizer, following our setup for each other baseline, and GPT-4o-mini as the DocETL optimizer, following the DocETL preprint (Shankar et al., 2024).
+However, we find both fail to produce a successful run in Section 5.2.
+After reaching out to the authors, we confirm a mistake in the pre-print and find that GPT-4o is required for the DocETL optimizer for our evaluation on BioDEX in Section 5.2.
+We also note we found several bugs in the DocETL codebase, which we have done our best to fix and have also coordinated with the authors about.
+Since the codebase is an evolving artifact, we report the commit number we used for our evaluation666https://github.com/ucbepic/docetl/commit/93050998077a9eb6fc1ee99dc96d1e0a222a987f.
+LOTUS.
+We set default accuracy targets of and failure probabilities of .
+We provide a detailed ablation, varying these parameters in Section 5.5.
+Additionally, our sampling size uses of the data, similar to other works (Shankar et al., 2024), and a minimum sample size of for smaller datasets, following prior work (Kang et al., 2020).
+5.1. Fact-Checking
+Fact-checking systems determine whether a given claim is correct, typically based on verifiability with a knowledge corpus.
+FacTool (Chern et al., 2023) is a recent open-source research work that provides a multi-step fact-checking pipeline, involving claim extraction, query generation, tool querying, evidence collection, and verification.
+We consider the task of using constructing the FacTool pipeline, which was originally written in over 750 lines of code.
+We use the FEVER (Thorne et al., 2018) dataset, a claim verification dataset, for our evaluation.
+Claim verification is based on a corpus of million Wikipedia articles, and each claim is labeled with one of three gold labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”.
 We merge the latter two labels into a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation.
+To control for our API spending budget, we sample 1,000 claims from the development dataset.
 Baselines.
-FacTool (Chern et al., 2023) is a recent research work that proposes a multi-step pipeline for fact-checking involving, claim extraction, query generation, tool querying, evidence collection, and verification.
-We use FactTool’s open source codebase (gai, [n.
-d.]) to measure its performance.
-FactTool’s pipeline, by default, performs retrieval with a Google Search API (cus, [n.
-d.]).
+For all baselines we use ColBERT (Khattab and Zaharia, 2020) as the retreiver model777FactTool’s pipeline, by default, performs retrieval with a Google Search API (noa, [n.d.]a).
 We evaluate the pipeline with both the default retrieval API, and with the open-source ColBERT (Khattab and Zaharia, 2020) index.
-We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the implemented LOTUS programs.
-LOTUS Programs.
-We compose several intuitive LOTUS programs, each in less than 50 lines of code.
-For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama-70B as the primary LM, and Llama-8B as the cascade proxy.
-First, we compose a pipeline designed to directly re-implement FacTool’s information flow in LOTUS.
-Figure 6 shows the pseudocode for the LOTUS-FacTool pipeline.
-The two tables are shown by wiki_df, which stores the Wikipedia articles, and claim_df, which stores the claims from the FEVER dataset.
-After loading the semantic index to wiki_df, the pipeline first performs a semantic map over each claim to generate two search queries, which are then used to perform a semantic similarity join over the corpus of Wikipedia articles.
-The program then concatenates the context retrieved for each claim, and performs a sem_map to output whether the claim is true or false, along with a revised claim if the claim is false.
-We use the same prompts found in FacTool (gai, [n.
-d.]), which include 3 demonstrations for generating search queries in the first sem_map, and chain-of-thought prompting in the second sem_map.
-| Method | Accuracy | Execution Time (s), batched | Execution Time (s), no batching | LoC |
+We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the other baselines.
+and Llama-70B served with vLLM.
+We run FactTool’s open source codebase (noa, [n.d.]b) to measure its performance.
+The AI UDF baseline uses a simple semantic map-search-map dataflow, following FacTool’s pipeline.
+First each claim is mapped to two search queries.
+Then the generated queries are used for search over the Wikipedia corpus.
+Lastly, each claim appended with retrieved context is mapped to truth label, along with a reasoning.
+We use the same prompts found in FacTool (noa, [n.d.]b), which include 3 demonstrations for generating search queries in the first semantic mapping, and chain-of-thought prompting in the second semantic mapping.
+The UQE baseline follows the first two steps of the AI UDF baseline, but replaces the final step with UQE’s embedding-based LLM-powered filter.
+Our LOTUS program similarly follows a map-search-filter pipeline and uses a Llama-8B proxy model.
+To provide a fair comparison, we tune the UQE LLM call budget to normalize execution time with the optimized LOTUS program.
+Since DocETL does not provide search primitives, we are unable to benchmark it in this application.
+| Method | Accuracy | ET (s), with batching | ET (s), no batching | LoC |
 | FacTool | 80.9 | N/A | 5396.11 | ¿ 750 |
-| LOTUS-Factool | 89.9 | 688.90 | 4,454.24 | ¡ 50 |
-| LOTUS-fact-filter | 91.2 | 329.1 | 988.95 | ¡ 50 |
-| LOTUS-fact-filter (opt.)* | 91.0 | 189.88 | 776.37 | ¡ 50 |
-| LOTUS-fact-join | 84.5 | 11,951.35 | 60,364.39 | ¡ 50 |
-The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the program, with a semantic filter.
-We use 3 demonstrations for the filter.
-Figure 7 shows the pseudocode for this program.
-We evaluate this program with and without model cascades applied to the semantic filter.
-Lastly, we compose an alternative pipeline, LOTUS-fact-join.
-As the pseudocode shows in Figure 8, the pipeline first performs a semantic map over each claim to obtain a set of sub-claims.
-From this, we create the claimed_facts_df, which separates each sub-claim into a different row.
-Next, the pipeline uses these sub-claims to perform a semantic similarity join over the Wikipedia corpus, then performs a semantic map over each retrieved article to generate the important facts in each one.
-Lastly, the program performs a semantic join between the sub-claims and the facts described in the retrieved passages.
-If the returned table contains a supporting fact for each sub-claim, the claim is labeled as ”Supported”.
-For the sem_map and join operations, we use 3 demonstrations each.
+| AI UDF: map, search, map | 89.9 | 688.9 | 4,454.2 | ¡ 50 |
+| AI UDF map, search + UQE filter | 66.0 | 184.4 | 738.3 | 150 |
+| LOTUS map, search, filter (unopt.) | 91.2 | 329.1 | 989.0 | ¡ 50 |
+| LOTUS map, search, filter (opt.) | 91.0 | 190.0 | 776.37 | ¡ 50 |
 Results.
-Table 2 demonstrates the powerful abstraction that semantic operators provide, allowing programmers to quickly write and test programs that compose a few operators to obtain state-of-the-art results.
-We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching averaged over 10 runs.
-We see that FacTool’s implementation offers strong accuracy performance on the FEVER dataset, however the full repository was written from scratch in several hundred lines of code, representing a significant development burden.
-By contrast, each LOTUS program offers comparable or higher accuracy, in relatively few lines of code.
-We also note that FacTool implements its method without batching, whereas LOTUS, by default, leverages batched LM execution for efficiency.
-To provide an apples-to-apples comparison, we compare FacTool’s un-batched implementation to the LOTUS programs both with and without batching.
-First, we see from the Table 2 that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with points higher accuracy and lower execution time without batching.
-The LOTUS-FacTool implementation with batching further decreases execution time compared to FacTool by .
-We see that the next LOTUS pipeline simply changes a single operation, switching the sem_map to a sem_filter, and maintains similar result quality to LOTUS-FacTool while further reducing execution time by in the batched implementation.
-Leveraging LOTUS’ filter operation allows the programmer to further declaratively optimize the program using an approximation, which maintains high accuracy and reduces the batched execution time by compared to LOTUS-FacTool.
-Figure 9 highlights the diverse performance trade-offs presented by the model cascade optimizations used in LOTUS’ filter sem_filter.
-The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, or Llama 70B,
-shown by the circles.
-We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, to implement the filter with our casacade-based approximation.
-We generate multiple cascade points, shown by the stars, by varying the recall and precision targets of the filter operator, and setting the error probability .
-The plot shows that this filter optimization can substantially reduce execution time, and offer diverse accuracy trade-offs, compared to implementing the pipeline with the oracle model, Llama 70B, alone.
-We also verify that the algorithm’s probabilistic guarantees hold by testing 100 trials for recall and precision targets at .
-We find the targets are met in 99 of the 100 trials, satisfying and exceeding the guaranteed 0.8 success rate.
-Returning to Table 2, we find that the last LOTUS pipeline (LOTUS-fact-join), reduces accuracy and increases execution time.
-Notably the performance trade-offs of different AI programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines.
-Each LOTUS program we described can be implemented easily in relatively few lines of code, and the best performing one offers higher accuracy and lower execution time than FacTool’s original, un-batched implementation.
-5.2. Application: Extreme Multi-label Classification
-Dataset.
-We evaluate on the Biodex Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and expert-created drug safety reports constructed from each article.
-The task is to correctly label the drug reactions experience by the patient in each medical article.
-We sample 250 patient articles for our evaluation.
+We report the accuracy of each method, an estimate of lines of code (LoC), and average execution time over 10 runs in seconds, both with and without batching to provide a fair comparison with FacTool, which provides a sequential implementation only.
+Overall, Table 2 demonstrates that the optimized LOTUS program reproduces state-of-the art accuracy, comparable to that of FacTool, with faster unbatched execution and faster batched execution, and in less than 50 lines of user code, while also outperforming each baseline.
+Comparing the optimized LOTUS program to the AI UDF baseline, we see that both acheive comparable accuracy in relatively few lines of code.
+However, the LOTUS sem_filter can be more heavily optimized compared to the generic AI UDF map operation.
+This allows the LOTUS program to attain faster execution, attributable to the sem_filter’s short LLM generations, which leads to cheaper decoding, and our proxy-based optimization for filters.
+We also observe that the LOTUS program outperforms the UQE program by accuracy at an equivalent latency budget.
+While our filter optimization adaptively learns when to leverage the proxy in order to meet an accuracy target, UQE’s embedding-based approximation provides no accuracy guarantees.
+Finally, we compare the optimized LOTUS program to the un-optimized LOTUS program.
+We observe that the optimized LOTUS program achieves accuracy relative to the un-optimized one, while reducing batched execution time by .
+Figure 6 highlights the diverse performance trade-offs achievable by our sem_filter approximation.
+The plot shows the accuracy and execution time of the LOTUS program using the gold algorithm with Llama 70B, or using the proxy only, shown respectively by the green and red circles.
+By varying the filter recall and precision target, our cascade-based approximation offers diverse trade-offs between accuracy and execution time, shown by the stars.
+5.2. Biomedical Multi-label Classification
+Biomedical classification entails processing complex, lengthy patient documents.
+We consider the extreme multi-label classification task on the BioDEX Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and ground-truth expert-created drug safety reports constructed from each article.
+The task is to classify the patient’s drug reactions, given each medical article.
 Notably, there are possible drug-reaction labels, making this task an extreme multi-label classification task.
-Due to the large number of possible labels, leveraging an LM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024).
-We show below that this task can be efficiently modeled and programmed using the semantic join.
-Baselines We consider a retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels.
+Due to the large number of possible labels, leveraging an LLM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024).
+We show below that this task can be efficiently modeled and optimized using the semantic join.
+Similar to prior work (D’Oosterlinck et al., 2023) we sample 250 patient articles for our evaluation to control for our API spending budget.
+Baselines.
+The search baseline performs batched similarity search over each patient article and the set of reaction labels.
+For the baseline LLM-based analytics systems, we implement a LLM-powered join over the patient articles and set of candidate reaction labels, followed by a ranking step, if supported.
+The AI UDF program entails a naive, nested-loop join pattern using row-wise LLM operators; however we note this is prohibitively costly and we report estimated latency.
+While UQE does not explicitly study joins, we implement a UQE join using its optimized filter.
+The DocETL program uses a join and reduce, which uses listwise ranking, following the preprint (Shankar et al., 2024).
+We note that we are unable to successfully run DocETL using Llama-70B or GPT-4o-mini for its optimizer so results (Table 3, 4) are shown using GPT-4o for the DocETL optimizer.
+LOTUS supports both join and ranking.
+Although LOTUS provides a sem_topk operator, we limit our ranking step to list-wise ranking to maintain a fair comparison with DocETL, and revisit ranking in Section 5.3.
+Our main results are shown in Table 3, using Llama-70B and E5 embeddings and sampling size of .
+We also provide an additional set of benchmarks in Table 4 comparing LOTUS and DocETL following the models, prompts and parameters used in the original DocETL preprint.
+The DocETL setup uses GPT-4o-mini, text-embedding-3-small embeddings and a sampling size of 500.
+Once again, DocETL fails when we attempt to use GPT-4o-mini for the optimizer, so we use GPT-4o.
+DocETL’s join takes an LLM call budget, which we control to compare against the other baselines.
 | Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
-| Sem-sim-join | 0.106 | 0.120 | 2.91 | 0.00 |
-| LOTUS Sem-join** | 0.244 | 0.261 | 5,891.6 | 15,107 |
-| nested-loop pattern | N/A | N/A | 2,144,560* | 6,092,500 |
-| map-sim-filter pattern** | 0.244 | 0.261 | 5,891.6 | 15,107 |
-| sim-filter pattern** | 0.154 | 0.191 | 24,206.8 | 63,724 |
-Estimated under linear-scaling assumption in number of batch calls.
-LOTUS Programs.
-Our LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 10.
-We perform the semantic join using at most 7 map demonstrations.
-We use Llama-70b as the LM and the E5Model (Wang et al., 2024) as the retriever for the semantic index.
+| Search | 0.106 | 0.120 | 2.91 | 0.00 |
+| AI UDF * | N/A | N/A | 2,144,560 | 6,092,500 |
+| UQE | 0.115 | 0.114 | 6,559 | 15,000 |
+| DocETL Join (avg of 3 successes)** | 0.180 | 0.219 | 2050 | 13,185 |
+| DocETL Join + Rank (avg of 3 successes)** | 0.262 | 0.282 | 2,342 | 13,433 |
+| LOTUS Join | 0.212 | 0.213 | 2,340 | 5,644 |
+| LOTUS Join + Rank | 0.265 | 0.280 | 2,503 | 5,869 |
+| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
+| DocETL Join + Rank (avg of 3 successes)** | 0.268 | 0.287 | 583 | 35,669 |
+| LOTUS Join + Rank | 0.289 | 0.296 | 647 | 32,026 |
+* AI UDF is infeasible to run and latency is estimated under linear-scaling assumption in number of batched LLM calls.
+** The DocETL baselines requires using GPT-4o as its optimizer and still exhibit frequent failures, with, on average, 1 in 3 runs failing.
+We rerun the system multiple times and report results averaged only over 3 successful runs.
 Results.
-Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while providing query efficiency compared to naive join implementations.
-The table reports the rank-precision@5 (RP@5) and rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls used.
-Since the nested-loop join pattern is prohibitively expensive to run, we show an estimate of the execution time assuming linear scaling in the number of batched calls.
-First we compare performance of the LOTUS join program, declaratively optimized, to the baseline, a semantic similarity join.
-As expected, the optimized LOTUS join offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the similarity join.
-This highlights the effectiveness of leveraging LMs over the data for complex reasoning-based tasks.
-We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4-turbo.
+Table 3 demonstrates the LOTUS programs consistently match or exceed the accuracy of all baselines, while also providing performance guarantees, unlike DocETL, which requires multiple reruns and manual selection of the optimizer LLM to produce high-quality results.
+The table reports the rank-precision@5 (RP@5) and rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LLM calls.
+The LOTUS join with ranking program achieves higher rank-precision than the standalone LOTUS join program, and we show both.
+We begin by informally comparing these accuracy results to that of D’Oosterlinck et al.
+(D’Oosterlinck et al., 2024), a state-of-the art AI pipeline that composed a multi-step DSPy (Khattab et al., 2023) program compiled using a combination of Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4-turbo.
 D’Oosterlinck et al.
-report 24.73 RP@5 and 27.67 RP@10 for the compiled program, representing comparable result quality to LOTUS’ program, although notably the LOTUS program was not compiled with a prompt optimization system.
-We now consider several interesting performance trade-offs presented by semantic join algorithms.
-We compare the map-sim-filter join pattern, which was automatically selected in the LOTUS sem-join, to the naive nested-loop join algorithm and the sim-filter join pattern shown in Table 3.
-We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is prohibitively costly.
-By contrast, the sim-filter and map-sim-filter pattern substantially reduce the LM call budget of the naive algorithm by and respectively.
-LOTUS’ optimizer automatically selects the lower cost algorithm, in this case map-sim-filter.
-The selected pattern also results in significantly better accuracy due to its higher quality proxy score for this task.
-Specifically, the map-sim-filter pattern offers higher RP@5 and higher RP@10 compared to the sim-filter.
-5.3. Application: Search & Ranking
-Dataset.
-We evaluate LOTUS’ performance on the search and ranking task using three datasets: BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, and two new benchmarks, CIFAR-bench, and HellaSwag-bench, which we generated to evaluate more complex ranking criteria over the data.
-For each, we report nDCG@10 and execution time (ET) in seconds.
-The SciFact dataset consists of a set of scientific claims and a corpus of articles, where the task is to rank articles by relevance to a given claim.
-We sample 300 claims for our evaluation, and report the average ranking execution time across these samples.
-While the SciFact dataset provides a sorting task based on a simple relevance criterion, our generated benchmarks provide a more complex sorting criteria each over a corpus of paper abstracts.
-The task is to rank papers according to highest reported accuracy on CIFAR-10 and HellaSwag in CIFAR-bench and HellaSwag-bench, respectively.
-To generate CIFAR-bench, we took 100 abstracts from the Papers with Code Dataset (pap, [n.
-d.]) that state performance on CIFAR-10 in the abstract, and we and manually labeled their accuracy to obtain ground truth.
-We then synthetically generated HellaSwag-bench by prompting Llama-70B to create 200 paper abstracts, each with a specified accuracy value, randomly sampled from .
-These datasets allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth an ranking criteria other than relevance-based ones.
-We note that an alternative approach to these tasks could efficiently leverage sem_map to extract accuracy values on abstracts from either dataset, then perform a structured sort.
-However, our evaluation focuses on assessing the semantic ranking capabilities of various top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may generalize to a wider set of reasoning-based ranking queries.
-We report results for trials at temperature , similar to prior works (Khattab et al., 2023)
-| Method | nDCG@10 | ET (s) |
-| Search | 0.712 | 0.009 |
-| Search + Reranker | 0.741 | 2.64 |
-| LOTUS Top-k - Llama-70B | 0.775 | 33.6 |
-| LOTUS Top-k - GPT-4o | 0.800 | 11.2 |
-| Method | CIFAR | HellaSwag | ||
-| nDCG@10 | ET (s) | nDCG@10 | ET (s) | |
-| Search | 0.252 | 0.008 | 0.119 | 0.008 |
-| Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
-| LOTUS Top-k - Llama 70B | 0.710 | 39.6 | 0.975 | 63.6 |
+report 24.73 RP@5 and 27.67 RP@10 for the compiled program, which is comparable to result quality achieved by the simple LOTUS join with ranking program.
+Turning our attention to the baselines measured in Table 3, we see that, first, the search baseline offers low results quality, with the LOTUS programs achieving over higher RP@5 and RP@10.
+Next, The AI UDF baseline, using row-wise LLM calls, is prohibitively expensive, and reflects the cost of our gold algorithm for joins.
+Its quadratic scaling of LLM call complexity with respect to dataset size leads to higher estimated cost relative to the optimized LOTUS join program.
+UQE, which uses an embedding-based optimization, obtains similar accuracy to the search baseline, remaining well below the RP@5 and RP@10 of the LOTUS programs despite a generous LLM call budget, indicating that the UQE method has lower sample efficiency on this task.
+Next, we turn to a detailed comparison of DocETL and LOTUS, shown in Table 3 and Table 4.
+LOTUS consistently matches or exceeds the accuracy of DocETL’s successful runs.
+Notably, the DocETL agentic optimizer requires GPT-4o and frequently fails, requiring multiple reruns, while LOTUS’ optimizer provides statistical accuracy guarantees, which we provide a detailed analysis on in Section 5.5.
+Table 4 also shows that the LOTUS and DocETL programs use a similar number of total LLM calls and execution time, however the token consumption per call of the DocETL programs are lower on average due to its optimized plan, which leads to modestly lower execution time despite modestly higher LLM calls.
+| Method | RP@5 | RP@10 | Execution Time (s) | # LM Calls |
+| LOTUS Join Plan 1 | 0.1541 | 0.170 | 12,563 | 27,687 |
+| LOTUS Join Plan 2 (chosen) | 0.212 | 0.213 | 2,116 | 5,290 |
+| Gold Algorithm | N/A | N/A | 2,144,560 | 6,092,500 |
+Lastly, we turn to Table 5 and study the candidate plans produced by the LOTUS join program.
+Plan 1, the sim-filter, requires more LLM calls to meet the recall and precision targets, compared to Plan 2, which is the project-sim-filter plan.
+This reflects that Plan 2’s proxy offers a stronger signal for predicate evaluations.
+We see that the LOTUS optimizer selects Plan 2 to execute since it is lower cost, requiring fewer LLM calls.
+The selected pattern also results in significantly better accuracy due to its higher quality proxy signal for this task.
+Specifically, the project-sim-filter pattern offers higher RP@5 and higher RP@10 compared to the sim-filter, with fewer LLM calls than the gold algorithm.
+5.3. Search & Ranking
+Relevance-based ranking has been widely studied in the context of information retrieval.
+In addition, our conversations with LOTUS users reveal a common need for ranking based on complex natural language criteria (e.g., ranking customer reviews based on how frustrated they sound).
+In this application, we assess LOTUS’ ranking capabilities on both types of ranking tasks with two datasets: BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, and a new dataset, HellaSwag-bench, which we generated to assess more complex ranking tasks.
+For both we report average nDCG@10, a standard ranking metric, and execution time (ET).
+The SciFact dataset consists of scientific claims and a corpus of articles, where the task is to rank articles by relevance to a given claim.
+We sample 300 claims for our evaluation.
+HellaSwag-bench consists of 200 synthetic paper abstracts888Each paper abstract in HellaSwag-bench is synthetically generated by prompting Llama-70b to write a research abstract that claims a specified accuracy value, which we randomly sample from ., each reporting an accuracy on the HellaSwag dataset (Zellers et al., 2019), and the task is to rank abstracts according to their reported accuracy.
+This dataset provides an objective ground truth, while focusing on a reasoning-based ranking criteria, rather than a relevance-based one999While the Hellaswag-bench ranking task might be efficiently solved with a sem_map, to extract accuracy values, followed by a structured sorting, our evaluation focuses on assessing semantic ranking algorithms.
+We find this benchmark useful for understanding performance trade-offs, which may generalize to a wider set of reasoning-based ranking queries..
+We report results for trials at temperature , similar to prior works (Khattab et al., 2023).
 Baselines.
-We consider two simple baselines: a semantic search, using the E5Model (Wang et al., 2024) for retrieval, and a search with re-ranking, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking.
-We use choose these models because they are recent high quality embedding and re-rankers.
-LOTUS Program. Our LOTUS program performs a semantic top- over the documents.
-The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets.
-For SciFact, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk.
-We report results using both Llama-70B and GPT-4o as the LM.
-| Method | Scifact | CIFAR | HellaSwag | ||||||
-| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
-| Quadratic Sort | 0.836 | 712 | 4950 | 0.868 | 634 | 4950 | 0.966 | 1,803 | 19,900 |
-| Heap Top-k | 0.776 | 65.0 | 216 | 0.832 | 99.6 | 350 | 0.907 | 98.9 | 415.2 |
-| QuickSelect Top-k | 0.776 | 42.4 | 285 | 0.746 | 41.3 | 303.95 | 0.909 | 59.1 | 620.95 |
-| QuickSelect Top-k + Semantic Index | 0.775 | 33.6 | 229 | 0.710 | 39.6 | 307.7 | 0.975 | 63.6 | 672.7 |
+In addition to the standard search baseline, we add a re-ranking method using the MixedBread cross-encoder (noa, 2024), a high quality re-ranker that is common in relevance-based retrieval.
+The AI UDF baseline performs a pointwise-ranking by prompting the LLM to assign a relevance score to each row, similar to prior work (Zhuang et al., 2024).
+The DocETL baseline uses the system’s LLM reduce operation for ranking, as described by the pre-print (Shankar et al., 2024).
+Our LOTUS program uses the sem_topk operator with a semantic index on the document corpus of both datasets.
+For the reranker, AI UDF, DocETL, and LOTUS programs on SciFact, we perform search to retrieve 100 articles, before re-ranking them with each method.
+| Method | SciFact | HellaSwag-bench | ||||
+| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
+| Search | 0.712 | 0.009 | 0 | 0.119 | 0.008 | 0 |
+| Reranker | 0.741 | 2.64 | 0 | 0.461 | 2.36 | 0 |
+| AI UDF | 0.457 | 9.83 | 100 | 0.091 | 19.7 | 200 |
+| DocETL - Llama 70B optimizer (avg. of 3 sucesses)* | N/A | N/A | N/A | 0.246 | 14.1 | 3 |
+| DocETL - GPT 4o optimizer* | N/A | N/A | N/A | N/A | N/A | N/A |
+| LOTUS | 0.765 | 36.3 | 213.2 | 0.919 | 57.0 | 506.4 |
+* DocETL is unable to acheive any successful runs using GPT-4o for its optimizer on SciFact and HellaSwag-bench, nor using Llama-70B for its optimizer on SciFact
+| Method | SciFact | HellaSwag-bench | ||||
+| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
+| Quadratic Topk | 0.768 | 801.2 | 4950 | 0.886 | 2116.6 | 19,900 |
+| Heap Topk | 0.765 | 60.0 | 192.6 | 0.896 | 59.4 | 241.5 |
+| Quickselect Topk | 0.771 | 40.5 | 237.0 | 0.893 | 49.4 | 448.1 |
+| LOTUS | 0.765 | 36.3 | 213.2 | 0.919 | 57.0 | 506.4 |
 Results.
-Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- for complex search tasks.
-In addition, Table 6 highlights the rich implementation design space that this task presents.
-We walk through several note-worthy findings.
-We first turn our attention to Table LABEL:tab:ranking_scifact, which presents the results for each bench-marked method on the SciFact dataset, which uses a relevance-based sorting criterion.
-As expected, both semantic search programs with and without re-ranking present strong baselines.
-Specifically, the re-ranker model, which is a supervised model trained specifically for the task of relevance-based ranking, increases nDCG@10 by 3 percentage points, while trading off query efficiency compared to the semantic search baseline.
-Notably, the unsupervised LM-based LOTUS programs outperform the supervised re-ranker’s result quality.
-The table shows LOTUS semantic top-k program with Llama-70B and GPT-4o, which outperform the semantic search baseline by 6 and 9 percentage points respectively.
-The LOTUS programs with Llama-70B and GPT-4o also outperform the re-ranker by 3 and 5 points respectively, improving upon the quality of the supervised baseline.
-As expected, the improved result quality comes at a trade-off to query efficiency due to the cost of LM calls.
-Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipelines and trade-off result quality and efficiency depending on application-specific requirements.
-Turning our attention to Table 5, we study LOTUS’ generality in supporting arbitrary language-based ranking criteria over the dataset.
-On the CIFAR-bench and HellaSwag-bench datasets, which use a complex sorting criteria, we see that both semantic search baselines with and without re-ranking provide poor result quality with consistently low nDCG@10.
-The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with and higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively.
-These accuracy gains reflect the powerful reasoning capabilities of LMs efficiently orchestrated over the data.
-As expected, these significant accuracy gains come at an increase to execution time.
-We now analyze the efficiency of our semantic top- implementation along with it’s proposed optimizations compared to alternatives.
-Table 6 compares several semantic top- algorithms, namely a quick-select top- algorithm, a quick-select top- that leverages the similarity index for pivot selection, a heap-based top- algorithm, and a quadratic sorting algorithm.
-First, we see that the quadratic algorithm, which performs an LM comparison between each pair of input documents, offers consistently high result quality across each dataset.
-However, this method is prohibitively expensive, requiring more LM calls and over higher execution time than the alternative implementations.
-The heap top- and quick-select top- methods offer comparable result quality, but with interesting trade-offs in query efficiency.
-Notably the quick-select top- method offers lower execution time than the heap-based sorting method across all datasets, despite requiring more LM calls in some cases.
-This is because the quick-select top- implementation allows for efficient batch-processing in each round of the algorithm, whereas the heap-based top- incurs sequential LM calls during heap updates.
-For this reason, our current implementation leverages the quick-select top- algorithm, although we envision future iterations may leverage multiple algorithms and allow the user to declaratively trade-off accuracy, query throughput, and cost.
-In addition to providing an efficient top- algorithm, the table also demonstrates our use of the similarity index for optimizing top- query performance.
-The quick-select top- algorithm optimized with the semantic similarity index for pivot selection demonstrates lower execution time at no accuracy loss on SciFact, where the ranking criteria correlates likely with semantic similarity.
-On the other hand, for the CIFAR-bench and HellaSwag-bench datasets, where the ranking criteria does not correlate with semantic similarity, we see that the similarity index has no significant impact on the accuracy or efficiency top- performance.
-5.4. Application: ArXiv Paper Analysis
-Dataset.
-Lastly, we evaluate LOTUS’ use in analyzing a large corpus of documents, specifically focusing on the performance of sem_group_by and its approximation algorithm with accuracy guarantees.
-The dataset consists of 647 recent ArXiv articles we scraped, from the database (cs.DB), information retrieval (cs.IR), cryptography and security (cs.CR), and robotics (cs.RO) domains.
-The task is to group the papers according to the topics they discuss.
-Baselines.
-The sem_group_by consists of two sub-tasks: (a) discovering the groups and labels, and (b) classifying each document using the discovered labels.
-As a baseline for the latter, we consider a semantic similarity join between the labels and each abstract.
-We use the E5 embedding model for retrieval.
-IRIS Programs.
-We use the LOTUS sem_group_by operator to find 5 groups based on the topic abstracts.
-We use Llama-70B as the LM, the E5Model as the retrieval model, and 100 samples for the approximation algorithm described in Section 4.
-| () Advancements in Recommender Systems and Multimodal Data Integration |
-| () Advancements in Generative Information Retrieval Systems |
-| () Advancements in Large Language Models for Various Applications |
-| () Advancements in AI Security and Malware Detection Techniques |
-| () Advancements in Robotic Navigation and Manipulation Techniques |
+We evaluated the performance of the LOTUS program against baselines, in comparison to the gold algorithm, and to alternative high-quality ranking algorithms we considered as gold algorithms for our sem_topk.
+First Table 6 demonstrates that LOTUS achieves higher accuracy than the next best-performing baselines on Scifact and HellaSwag-bench.
+Specifically, as expected, the re-ranker offers competitive accuracy compared to the LOTUS program on Scifact, which focuses on relevance-based retrieval, the supervised task the re-ranker model was trained on.
+However, on HellaSwag-bench, the LOTUS program significantly outperforms the re-ranker on this reasoning-based task.
+We also observe that the point-wise AI UDF method and the list-wise ranking, used by DocETL, offer substantially lower accuracy, unable to outperform the re-ranking baseline.
+This reflects the limitations of point-wise and list-wise ranking, which has been studied in prior work (Qin et al., 2024).
+Notably, the DocETL baseline altogether fails to produce any successful execution plans using the GPT-4o optimizer on Scifact and Hellaswag-bench and using the Llama-70B optimizer on Scifact.
+Turning our attention to Table 7, we compare the quick-select top-k gold algorithm to several alternative candidates.
+The quadratic top-k, heap top-k, and quick-select top-k offer comparable accuracy, making each a viable gold algorithm.
+However, the three candidate choices provide significant differences in efficiency.
+The quadratic algorithm, which performs an LLM comparison between each pair of input documents, requires more LM calls and over higher execution time than the heap top-k and quick-select top-k.
+In addition, the heap top- and quick-select top- methods offer interesting trade-offs in LLM call complexity and execution time.
+Notably the quick-select top- offers lower execution time than the heap-based sorting method, despite requiring more LLM calls in some cases.
+This is because the quick-select top- implementation allows for efficient batched processing in each round of the algorithm, whereas the heap-based top- incurs sequential LLM calls during heap updates.
+LOTUS currently implements the quick-select topk- by default, favoring lower execution time, although we future iterations may leverage multiple algorithms.
+Finally, the Table 7 compares the quick-select top-k with the LOTUS quick-select top-k using embedding-based pivot selection.
+We demonstrate that the optimization is lossless, and can decrease latency by .
+Specifically, this optimization is useful where document rank correlates with semantic similarity, which is likely on the Scifact dataset.
+| () | Advancements in Recommender Systems and Multimodal Data Integration |
+| () | Advancements in Generative Information Retrieval Systems |
+| () | Advancements in Large Language Models for Various Applications |
+| () | Advancements in AI Security and Malware Detection Techniques |
+| () | Advancements in Robotic Navigation and Manipulation Techniques |
+5.4. Topic Analysis on ArXiv Papers
+A common unsupervised discovery task requires grouping a large document corpus by topics and assigning descriptive labels to each group.
+We consider this task over a dataset of 647 recent ArXiv articles, which we scraped from the database (cs.DB), information retrieval (cs.IR), cryptography and security (cs.CR), and robotics (cs.RO) domains.
+We aim to discover key topic labels.
+We can succinctly represent this task using the sem_group_by operator with LOTUS.
+AI UDFs, DocETL and standalone search systems do not serve operations for unsupervised semantic-based group discovery.
+While UQE (Dai et al., 2024) studies a tractable optimization for serving non-LLM aggregations, e.g., COUNT, with LLM-based group-bys using stratified sampling, it does not focus on implementing, optimizing, and evaluating a standalone group-by operator, but suggests a standalone group-by implementation similar to our sem_group_by gold algorithm.
+Thus, in this section, we focus our analysis on understanding the performance of our sem_group_by gold algorithm and analyzing our optimization with respect to it.
+We use Llama-70B and E5 embeddings.
+For approximations, we use a sample size of 100 and .
 Results.
-We first qualitatively analyze the labels discovered, in the first phase of LOTUS’ implementation.
-As Figure LABEL:tab:groupby shows, the discovered labels intuitively align with the ArXiv domains included in the paper scrap, including important and timely topics related to recommendation systems, retrieval, LLM applications, AI security, and robotics.
-Performing this unsupervised group discovery took 44.03 seconds, representing a tractable and efficient implementation.
-Finally, turning our analysis to the classification subprocedure in our sem_group_by implementation, we evaluate the declarative algorithmic approximation.
-Figure 11 compares the execution time and classification accuracy with respect to the oracle implementation for the operator implemented with varying accuracy targets using .
-We see that the oracle implementation, at , offers the slowest execution time.
-On the other hand, the similarity-join baseline has faster execution time, but only accuracy with respect to the LM.
-We find our proxy-based optimization offers diverse performance trade-offs and allows the user to declaratively interpolate between these two extremes.
-We also note that the sampling procedure used to perform this optimization took less than 5 seconds, representing a small relatives cost.
+The sem_group_by consists of two sub-tasks: (a) discovering representative group labels, and (b) classifying each document to a discovered label.
+We first qualitatively analyze the labels discovered, in the first sub-task of the LOTUS implementation, which is equivalent to gold algorithm.
+As Figure 7 shows, the discovered labels intuitively align with recent research topics from the cs.DB, cs.IR, cs.CR, and cs.RO ArXiv domains related to recommendation systems, retrieval, LLM applications, AI security, and robotics.
+Performing this unsupervised group discovery took 44.03 seconds, representing a tractable implementation.
+Turning to the classification sub-task, we compare the performance of the gold algorithm to our approximation, which uses an embedding-based proxy here.
+Figure 8 compares the execution time and classification accuracy of the oracle-based gold algorithm (green circle), our approximations using both the Llama-70B oracle LLM and embedding-based proxy at varied accuracy targets (blue stars), and the embedding-based proxy alone (red circle).
+We see that the proxy-only baseline is faster than the oracle, but about less accurate.
+Our approximation algorithm allows users to declaratively interpolate between these two extremes by varying the accuracy target.
+We also note that the sampling-based optimization procedure used to perform this optimization took less than 5 seconds, representing a small relative cost.
+5.5. Accuracy Guarantees Evaluation
+Our approximation methods are designed to adaptively learn when to leverage a proxy to provide an accuracy guarantee, despite variations in proxy accuracy.
+To demonstrate the accuracy guarantees provided by our semantic operator optimizations, we study the sem_filter performance on the fact-checking performance with varied proxy models, recall and precision targets and failure probability parameters.
+Figure 9(a) and Figure 9(b) demonstrate the average accuracy observed over 20 trials for varied recall targets and precision targets.
+We evaluate the filter approximation with both the Llama-8B proxy (solid lines) and a weaker proxy, with TinyLlama-1B (dashed lines) with a failure probability of (red) and (blue).
+As expected, the average recall and precision values increase with increasing accuracy targets for either metric.
+Figure 9(c) demonstrates that decreasing the recall and precision target tends reduce the number of calls routed to the oracle.
+As expected, at a fixed accuracy target, our approximation requires more oracle calls when using the weaker proxy, in this case TinyLlama.
+Finally, in Figure 9(d) we record the percentage of 50 trials that do not meet the recall and precision target, both set to 0.9, using the TinyLlama proxy.
+The results confirm that the observed failure probabilities increase with the user-configured delta values, and are lower than the configured failure probabilities, as expected due to our conservative implementation.
 6. Related Work
-Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks.
-Palimpzest (Liu et al., 2024b) presents a declarative approach to data cleaning and extract-transform-load (ETL) tasks.
-The authors propose to automatically optimize relational operators with LM calls, and implement two row-wise relational operators, a newly proposed convert operator, which can be transparently optimized to perform entity extraction using LLMs, and an AI-based filter operation, logically similar to LOTUS’ sem_filter.
-The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis and proposes several others as future work.
-SUQL (Liu et al., 2024c) presents a SQL extension to support conversational agents with knowledge grounding over structured and unstructured data.
-Specifically, the system extends SQL with two new logically row-wise operators, answer, which prompts an LM to answer the user question over each row, and summary, which prompts an LM to provide a summary to the user over each row.
-The system can provide automatic optimizations, such as predicate re-ordering using a lazy evaluation approach and proposes to use retrieval to optimize some answer queries.
-ZenDB (Lin et al., 2024) and EVAPORATE (Arora et al., 2023) tackle the task of automatically ingesting and extracting semi-structured documents into structured tables that can be queried using standard relational operators and languages.
-ZenDB extracts structure using a semantic hierarchical tree index, which is integrated with a SQL query engine to support efficient query processing over the extracted attribute values.
-The systems implements several optimizations, including predicate reordering, push-down, and projection pull-up.
-Additionally, EVAPORATE performs efficient entity extraction from semi-structured data using LM-based code generation and weak supervision to ensemble candidate functions.
-In contrast to these works, semantic operators are a general-purpose query model designed to capture broad-ranging applications, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions, as an extension of the relational model.
-For example, users can use sem_map for row-wise entity extraction over unstructured text fields, but can also perform other operations over entire tables, such as sem_join.
-We also propose multiple execution algorithms and novel optimizations for these bulk operators (Section 4).
-Some of the optimizations in prior specialized systems, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
-LLM UDFs.
-Recent research (Liu et al., 2024a) and some existing DBMS vendors, such as
-BigQuery (ver, [n. d.]),
-Databricks (dat, [n.
-d.]) and Redshift (noa, 2023), offer LLM user-defined functions (UDFs) in SQL.
-LLM UDFs are a lower-level programming interface that is limited to row-wise LLM execution over the data, equivalent to our sem_map semantic operator.
-In contrast, our semantic operator programming model is high-level and declarative, allowing the system to automatically select which an algorithm to use and manage calls to the LM to support complex operations, including aggregations, ranking, and joins.
-Liu et al.
-(Liu et al., 2024a) study how to optimize LLM UDF calls, demonstrating performance gains with a de-duplication and prefix-sharing method that reorders rows and input fields to maximize key-value (KV) cache reuse during LLM calls.
-This is orthogonal to our optimizations but could likely accelerate batch LLM calls from LOTUS.
-LM Programming Frameworks.
-LM programming frameworks such as LangChain (lan, [n.
-d.]), LlamaIndex (lla, [n.
-d.]), and DSPy (Khattab et al., 2023), have gained significant popularity.
-They provide abstractions for programming with LMs, including utilities for creating prompts and post-processing LM outputs, and modules for common calling patterns such as RAG, chat-bots, and agents.
-DSPy also optimizes a pipeline of LLM calls to maximize a target metric by tuning prompts and model weights.
-However, all these frameworks focus on expressing a pipeline of LLM calls that takes a single small input (or a stream of inputs for chat), as opposed to programs for processing large datasets.
-While some of them support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized.
-In contrast, semantic operators and LOTUS are designed for bulk semantic processing and introduce novel operators and optimizations for this task.
-ML-based Query Processing.
-Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique usage patterns and optimization opportunities.
-MADLib (Hellerstein et al., 2012) extends SQL with abstractions for machine learning and descriptive statistics.
-Prior works such as NoScope (Kang et al., 2017), TASTI (Kang et al., [n.
-d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML models over large datasets, typically in video analytics.
+Data Systems for Row-wise LLM Operations.
+Many prior works focus on map-like, row-wise LLM-based operators, either with generic AI user-defined functions (UDFs) or specialized operators (Liu et al., 2024b; Lin et al., 2024; Arora et al., 2023; Liu et al., 2024c).
+AI UDF systems (MotherDuck, [n.d.]; WilliamDAssafMSFT, 2024; noa, [n.d.]d; dat, [n.d.]; noa, 2023; Liu et al., 2024a) provide a low-level, non-declarative programming interface for row-wise LLM invocations.
+These systems also integrate vector search, equivalent to our sem_search or sem_sim_join, and support SQL queries that combine row-wise AI UDFs with non-LLM operations (e.g.
+average, count).
+Alternatively, several recent works implement optimized row-wise LLM operations for specialized tasks, including data cleaning, extract-transform-load (ETL) and conversational agents.
+ZenDB (Lin et al., 2024), EVAPORATE (Arora et al., 2023) and Palimpzest (Liu et al., 2024b) provide LLM-based functions for extracting semi-structured documents into structured tables, logically similar to our sem_extract, and Palimpzest also implements an LLM-powered filter, logically similar to our sem_filter.
+SUQL (Liu et al., 2024c) augments SQL with a two logically row-wise operators, answer, for questions-answering over each row, and summary, for row-wise LLM-based summarization, to serve knowledge-grounded conversational agents.
+In contrast to these works, semantic operators are a general-purpose query model for semantic transformations over datasets, including both logically row-wise ones and more complex ones (e.g., joins, aggregations and ranking).
+Our detailed evaluation (Section 5) demonstrates the performance limitations of row-wise LLM execution models.
+Best-Effort LLM-Based Analytics Systems.
+Several recent systems (Anderson et al., 2024; Dai et al., 2024; Shankar et al., 2024) go beyond row-wise LLM operators.
+However, these systems lack a formalism for LLM-based operators, providing no performance guarantees for their execution.
+Following a preprint (Patel et al., 2024a) of this work, DocETL (Shankar et al., 2024) proposes to use LLM agents as the query optimizer over LLM-powered operators.
+This is a promising direction for future work; however, our evaluation (Section 5) demonstrates that DocETL’s agentic approach leads to high variance in performance, no accuracy guarantees, and requires multiple reruns to achieve comparable accuracy to ours, with approximately 1 in 3 runs failing due to optimizer errors, indicating that more work is needed to make these methods robust.
+UQE (Dai et al., 2024) also studies optimizations for LLM-powered operators, proposing an embedding-based approximation for LLM-powered filters.
+In contrast to our methods, which adaptively learn how to leverage a given proxy to meet accuracy targets, UQE’s approach provides best-effort performance with no accuracy guarantees.
+UQE also studies optimizations for non-LLM aggregations (e.g., average, count) combined with LLM-based filter and group-by operators using stratified sampling, which is complementary to this work.
+General ML-based Query Processing.
+Prior works study the use of non-LLM machine learning (ML) in databases.
+In contrast, our work integrates LLMs, which motivates our new formalisms for language-based operators and novel optimizations.
+MADLib (Hellerstein et al., 2012) extends SQL with abstractions for descriptive statistics and machine learning (e.g., regression and classification).
+NoScope (Kang et al., 2017), TASTI (Kang et al., [n.d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML-based predicates (e.g., using object detectors) over large datasets, typically in video analytics.
 Some optimizations proposed in these works, such as model cascades and predicate re-ordering, are also useful for optimizing LOTUS pipelines with language models.
-However, our bulk semantic processing setting requires new operators with significantly different semantics, such as sem_join and sem_groupby, which were not optimized in this prior work.
-These motivate our novel optimizations that can approximate these operators with statistical guarantees (Section 4).
-Text2SQL A large body of prior work looks at Text2SQL (Yu et al., 2018; Yaghmazadeh et al., 2017; Zhang et al., 2024; Zelle and Mooney, 1996), using LMs to translate natural language queries to executable SQL code.
-These methods serve semantic queries that are expressible in traditioanl relational operators.
-By contrast, semantic operators provide new AI-based transformations, such as aggregations for summarizing large document sets, or rankings, joins and filters according to natural language expressions, which are not expressible in SQL.
-We note an interesting line of future work may build a natural language interface on top of semantic operators.
+Table Question Answering.
+A large body of work, including retrieval-augmented generation (Lewis et al., 2021) and Text2SQL (Yu et al., 2018; Yaghmazadeh et al., 2017; Zhang et al., 2024; Zelle and Mooney, 1996), serves natural language interfaces for question-answering over databases.
+Typically, the LLM system invokes external tools, e.g., search APIs or SQL programs.
+Interestingly, these agentic workflows can instead invoke semantic operator programs as the underlying API for data processing.
+Recent work (Biswal et al., 2024) demonstrates the promise of this approach to outperform baseline TableQA methods.
 7. Conclusion
-In this work, we proposed semantic operators to provide the first declarative and general-purpose interface to serve bulk-semantic processing.
-We implement these operators in the LOTUS system to seamlessly extend the relational model and allow programmers to easily compose powerful reasoning-based query pipelines over vast corpora of structured and unstructured data.
-Our results across a diverse set of applications, including fact-checking, extreme multi-label classification, and search, demonstrate the generality and effectiveness of the semantic operator model as well as new opportunities for efficient algorithms and optimizations.
-For each task, we find that LOTUS programs capture high quality and state-of-the-art query pipelines with low development overhead, and that they can be automatically optimized with accuracy guarantees to achieve higher performance than existing implementations.
+In this work, we proposed semantic operators for bulk semantic processing and provide the first formalized model for general-purpose, language-based transformations over data.
+Our results across a diverse set of applications, including fact-checking, biomedical multi-label classification, search, and topic analysis, demonstrate the generality, expressiveness and robustness of the semantic operator model as well as our optimization approach.
+For each task, we find that LOTUS programs capture state-of-the-art AI applications with low development overhead, match or exceed result quality of recent LLM-powered analytics systems, and substantially reduce cost, while ensuring accuracy guarantees.
+Since open-sourcing LOTUS, we have seen a growing user-base, as well as adoption of semantic operators among large database vendors.
+We believe both represent exciting steps towards new data systems that integrate powerful AI reasoning capabilities over vast knowledge corpora.
 Acknowledgements.
@@
 - arx ([n.d.]) [n.d.]. arXiv.org ePrint archive. https://arxiv.org
-- cus ([n. d.]) [n. d.]. Custom Search JSON API | Programmable Search Engine. https://developers.google.com/custom-search/v1/overview
+- noa ([n.d.]a) [n.d.]a. Custom Search JSON API | Programmable Search Engine. https://developers.google.com/custom-search/v1/overview
 - dis ([n.d.]) [n.d.]. Discovery Insight Platform. https://www.findourview.com
-- gai ([n. d.]) [n. d.]. GAIR-NLP/factool: FacTool: Factuality Detection in Generative AI. https://github.com/GAIR-NLP/factool
-- int ([n. d.]) [n. d.]. Introducing Meta Llama 3: The most capable openly available LLM to date. https://ai.meta.com/blog/meta-llama-3/
+- noa ([n.d.]b) [n.d.]b. GAIR-NLP/factool: FacTool: Factuality Detection in Generative AI. https://github.com/GAIR-NLP/factool
+- noa ([n.d.]c) [n.d.]c. Introducing Meta Llama 3: The most capable openly available LLM to date. https://ai.meta.com/blog/meta-llama-3/
 - lan ([n.d.]) [n.d.]. LangChain. https://www.langchain.com/
+- noa ([n.d.]d) [n.d.]d. Large Language Model (LLM) Functions (Snowflake Cortex) | Snowflake Documentation. https://docs.snowflake.com/user-guide/snowflake-cortex/llm-functions
 - ver ([n.d.]) [n.d.]. LLM with Vertex AI only using SQL queries in BigQuery. https://cloud.google.com/blog/products/ai-machine-learning/llm-with-vertex-ai-only-using-sql-queries-in-bigquery
-- ope ([n. d.]) [n. d.]. OpenAI Platform. https://platform.openai.com
+- noa ([n.d.]e) [n.d.]e. OpenAI Platform. https://platform.openai.com
 - pan ([n.d.]) [n.d.]. pandas - Python Data Analysis Library. https://pandas.pydata.org/
-- pap ([n. d.]) [n. d.]. Papers with Code - Machine Learning Datasets. https://paperswithcode.com/datasets
+- noa ([n.d.]f) [n.d.]f.
+python-bigquery-dataframes/notebooks/experimental/semantic_operators.ipynb at main · googleapis/python-bigquery-dataframes.
+https://github.com/googleapis/python-bigquery-dataframes/blob/main/notebooks/experimental/semantic_operators.ipynb
 - lla ([n.d.]) [n.d.]. Querying - LlamaIndex 0.9.11.post1. https://docs.llamaindex.ai/en/stable/understanding/querying/querying.html
@@
 http://arxiv.org/abs/2309.04269 arXiv:2309.04269 [cs].
-- Anantha et al.
-(2021) Raviteja Anantha, Svitlana Vakulenko, Zhucheng Tu, Shayne Longpre, Stephen Pulman, and Srinivas Chappidi.
-2021.
-Open-Domain Question Answering Goes Conversational via Question Rewriting.
-In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.).
-Association for Computational Linguistics, Online, 520–534.
-https://doi.org/10.18653/v1/2021.naacl-main.44
+- Anderson et al.
+(2024) Eric Anderson, Jonathan Fritz, Austin Lee, Bohou Li, Mark Lindblad, Henry Lindeman, Alex Meyer, Parth Parmar, Tanvi Ranade, Mehul A.
+Shah, Benjamin Sowell, Dan Tecuci, Vinayak Thapliyal, and Matt Welsh.
+2024.
+The Design of an LLM-powered Unstructured Analytics System.
+https://doi.org/10.48550/arXiv.2409.00847 arXiv:2409.00847 [cs].
 - Arora et al.
@@
 http://arxiv.org/abs/2304.09433 arXiv:2304.09433 [cs].
+- Biswal et al.
+(2024) Asim Biswal, Liana Patel, Siddarth Jha, Amog Kamsetty, Shu Liu, Joseph E.
+Gonzalez, Carlos Guestrin, and Matei Zaharia.
+2024.
+Text2SQL is Not Enough: Unifying AI and Databases with TAG.
+https://doi.org/10.48550/arXiv.2408.14717 arXiv:2408.14717 [cs].
 - Braverman and Mossel ([n.d.]) Mark Braverman and Elchanan Mossel. [n.d.]. Noisy sorting without resampling. ([n. d.]).
@@
 https://doi.org/10.48550/arXiv.2307.13528 arXiv:2307.13528 [cs].
+- Codd (1970) E F Codd. 1970. A Relational Model of Data for Large Shared Data Banks. 13, 6 (1970).
+- Dai et al.
+(2024) Hanjun Dai, Bethany Yixin Wang, Xingchen Wan, Bo Dai, Sherry Yang, Azade Nova, Pengcheng Yin, Phitchaya Mangpo Phothilimthana, Charles Sutton, and Dale Schuurmans.
+2024.
+UQE: A Query Engine for Unstructured Databases.
+https://doi.org/10.48550/arXiv.2407.09522 arXiv:2407.09522 [cs].
 - Dasgupta ([n.d.]) Sanjoy Dasgupta. [n.d.]. The hardness of k-means clustering. ([n. d.]).
@@
 https://doi.org/10.48550/arXiv.2305.13395 arXiv:2305.13395 [cs].
+- Douze et al.
+(2024) Matthijs Douze, Alexandr Guzhva, Chengqi Deng, Jeff Johnson, Gergely Szilvasy, Pierre-Emmanuel Mazaré, Maria Lomeli, Lucas Hosseini, and Hervé Jégou.
+2024.
+The Faiss library.
+https://doi.org/10.48550/arXiv.2401.08281 arXiv:2401.08281 [cs].
 - Drozdov et al.
@@
 https://doi.org/10.48550/arXiv.2310.14408 arXiv:2310.14408 [cs].
-- Gao et al.
-(2023) Luyu Gao, Zhuyun Dai, Panupong Pasupat, Anthony Chen, Arun Tejasvi Chaganty, Yicheng Fan, Vincent Y.
-Zhao, Ni Lao, Hongrae Lee, Da-Cheng Juan, and Kelvin Guu.
-2023.
-RARR: Researching and Revising What Language Models Say, Using Language Models.
-https://doi.org/10.48550/arXiv.2210.08726 arXiv:2210.08726 [cs].
 - Hellerstein et al.
@@
 - Hoare (1961) C. A. R. Hoare. 1961. Algorithm 65: find. Commun. ACM 4, 7 (July 1961), 321–322. https://doi.org/10.1145/366622.366647
-- Izacard et al.
-(2022) Gautier Izacard, Patrick Lewis, Maria Lomeli, Lucas Hosseini, Fabio Petroni, Timo Schick, Jane Dwivedi-Yu, Armand Joulin, Sebastian Riedel, and Edouard Grave.
-2022.
-Atlas: Few-shot Learning with Retrieval Augmented Language Models.
-https://doi.org/10.48550/arXiv.2208.03299 arXiv:2208.03299 [cs].
 - Johnson et al. (2017) Jeff Johnson, Matthijs Douze, and Hervé Jégou. 2017. Billion-scale similarity search with GPUs. http://arxiv.org/abs/1702.08734 arXiv:1702.08734 [cs].
@@
 https://doi.org/10.14778/3476249.3476285
-- Kang et al.
-([n.
-d.]) Daniel Kang, John Guibas, Peter Bailis, Tatsunori Hashimoto, and Matei Zaharia.
-[n.
-d.].
-Task-agnostic Indexes for Deep Learning-based Queries over Unstructured Data.
-([n.
-d.]).
-- Khattab et al.
-(2021) Omar Khattab, Christopher Potts, and Matei Zaharia.
-2021.
-Relevance-guided Supervision for OpenQA with ColBERT.
-Transactions of the Association for Computational Linguistics 9 (2021), 929–944.
-https://doi.org/10.1162/tacl_a_00405 Place: Cambridge, MA Publisher: MIT Press.
+- Kang et al. ([n.d.]) Daniel Kang, John Guibas, Peter Bailis, Tatsunori Hashimoto, and Matei Zaharia. [n.d.]. Task-agnostic Indexes for Deep Learning-based Queries over Unstructured Data. ([n. d.]).
 - Khattab et al.
@@
 https://arxiv.org/abs/2309.06180v1
-- Lee et al.
-(2019) Kenton Lee, Ming-Wei Chang, and Kristina Toutanova.
-2019.
-Latent Retrieval for Weakly Supervised Open Domain Question Answering.
-In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Anna Korhonen, David Traum, and Lluís Màrquez (Eds.).
-Association for Computational Linguistics, Florence, Italy, 6086–6096.
-https://doi.org/10.18653/v1/P19-1612
+- Lewis et al.
+(2021) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela.
+2021.
+Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks.
+https://doi.org/10.48550/arXiv.2005.11401 arXiv:2005.11401 [cs].
 - Liang et al.
@@
 - Ma et al. (2023) Xueguang Ma, Xinyu Zhang, Ronak Pradeep, and Jimmy Lin. 2023. Zero-Shot Listwise Document Reranking with a Large Language Model. https://arxiv.org/abs/2305.02156v1
+- MotherDuck ([n.d.]) MotherDuck. [n.d.]. Introducing the prompt() Function: Use the Power of LLMs with SQL! - MotherDuck Blog. https://motherduck.com/blog/sql-llm-prompt-function-gpt-models/
 - Patel et al.
-(2024) Liana Patel, Peter Kraft, Carlos Guestrin, and Matei Zaharia.
-2024.
+(2024a) Liana Patel, Siddharth Jha, Carlos Guestrin, and Matei Zaharia.
+2024a.
+LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data.
+https://doi.org/10.48550/arXiv.2407.11418 arXiv:2407.11418 [cs] version: 1.
+- Patel et al.
+(2024b) Liana Patel, Peter Kraft, Carlos Guestrin, and Matei Zaharia.
+2024b.
 ACORN: Performant and Predicate-Agnostic Search Over Vector Embeddings and Structured Data.
@@
 http://arxiv.org/abs/1512.08949 arXiv:1512.08949 [cs, math, stat].
+- Shankar et al.
+(2024) Shreya Shankar, Tristan Chambers, Tarak Shah, Aditya G.
+Parameswaran, and Eugene Wu.
+2024.
+DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing.
+https://doi.org/10.48550/arXiv.2410.12189 arXiv:2410.12189 [cs].
 - Sun et al.
@@
 http://arxiv.org/abs/2212.03533 arXiv:2212.03533 [cs].
+- WilliamDAssafMSFT (2024) WilliamDAssafMSFT.
+2024.
+Intelligent Applications - Azure SQL Database.
+https://learn.microsoft.com/en-us/azure/azure-sql/database/ai-artificial-intelligence-intelligent-applications?view=azuresql
 - Wu et al.
@@
 http://arxiv.org/abs/2310.03094 arXiv:2310.03094 [cs].
-- Yuksekgonul et al.
-(2024) Mert Yuksekgonul, Federico Bianchi, Joseph Boen, Sheng Liu, Zhi Huang, Carlos Guestrin, and James Zou.
-2024.
-TextGrad: Automatic ”Differentiation” via Text.
-http://arxiv.org/abs/2406.07496 arXiv:2406.07496 [cs].
-- Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah D. Goodman. 2022. STaR: Bootstrapping Reasoning With Reasoning. https://doi.org/10.48550/arXiv.2203.14465 arXiv:2203.14465 [cs].
 - Zelle and Mooney (1996) John M Zelle and Raymond J Mooney. 1996. 1996-Learning to Parse Database Queries Using Inductive Logic Programming. (1996).
+- Zellers et al.
+(2019) Rowan Zellers, Ari Holtzman, Yonatan Bisk, Ali Farhadi, and Yejin Choi.
+2019.
+HellaSwag: Can a Machine Really Finish Your Sentence?
+https://doi.org/10.48550/arXiv.1905.07830 arXiv:1905.07830 [cs].
 - Zhang et al.
@@
 https://doi.org/10.48550/arXiv.2403.02951 arXiv:2403.02951 [cs].
-- Zhang et al.
-(2022) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola.
-2022.
-Automatic Chain of Thought Prompting in Large Language Models.
-https://doi.org/10.48550/arXiv.2210.03493 arXiv:2210.03493 [cs].
+- Zhuang et al.
+(2024) Honglei Zhuang, Zhen Qin, Kai Hui, Junru Wu, Le Yan, Xuanhui Wang, and Michael Bendersky.
+2024.
+Beyond Yes and No: Improving Zero-Shot LLM Rankers via Scoring Fine-Grained Relevance Labels.
+http://arxiv.org/abs/2310.14122 arXiv:2310.14122 [cs].